### v1

# Meritocracy and Its Discontents: Long-run Effects of Repeated School Admission Reforms

Abstract
What happens if selective colleges change their admission policies? We answer this question by analyzing the world’s first introduction of nationally centralized meritocratic admissions in the early twentieth century. We find a persistent meritocracy-equity tradeoff. Compared to the decentralized system, the centralized system admitted more high-achieving applicants, producing a greater number of top elite bureaucrats in the long run. However, this impact came at the distributional cost of equal regional access to elite higher education and career advancement. Several decades later, the meritocratic centralization increased the number of urban-born career elites (e.g., top income earners) relative to rural-born ones.
Keywords: Elite Education, Market Design, Strategic Behavior, Regional Mobility, Universal Access, Persistent Effects
1 Introduction
College and school admission processes vary across time and place. American college admissions use a decentralized system where each college makes its own admissions often based on opaque evaluation criteria. In contrast, China’s public college admissions illustrate a regionally-integrated, single-application, and single-offer system using a transparent admission criterion. How do different admissions institutions affect students’ education and future career? What are the impacts on the national production of high-skilled individuals and their distributions?
In this paper, we study the impacts of using centralized and meritocratic admissions. We do so by combining a series of natural experiments in history, newly assembled historical data, and theoretical insights. Our investigations reveal the pros and cons of centralized meritocratic admissions, especially a tradeoff between meritocracy and equal access to selective higher education and career achievements.
Our empirical setting is the first known transition from decentralized to nationally-centralized school admissions. At the end of the 19th century, to modernize its higher education system, the Japanese government set up National Higher Schools (roughly equivalent to today’s colleges) as an exclusive entry point to the most prestigious tertiary education. [^1] These schools were transformed into parts of today’s national universities after WWII. For example, the National Higher School in Tokyo became the first two years of the University of Tokyo. These schools later produced many of the most influential members of the society, including several Prime Ministers, Nobel Laureates, and founders of global companies like Toyota. Their selective and prestigious nature and social influences had similarities to Oxbridge in the UK, Grande Ecoles in France, and the Ivy League in the US. [^2] We study top elites on both substantive and pragmatic grounds. The individual holdings of prestigious economic, political, and cultural positions are skewed and concentrated on a tiny minority of the population who have big influences on the economy and society. For their social importance, we also have relatively rich historical data that allows us to trace back their careers from humble beginnings as students to wide societal recognition and acclaim.
Acceptance into these schools was merit-based, using annual entrance examinations. Initially, the government let each school run its own exam and admissions based on exam scores, similar to many of today’s decentralized K-12 and college admissions. Under this decentralized system, the schools typically held exams on the same day so that each applicant could apply for only one school. [^3] Similar restrictions on the number of applications exist today in the college admission systems of Italy, Nigeria, and the UK. As a result, many high achieving applicants applied to too competitive schools and failed to enter any school.
This problem motivated the government to introduce a centralized system in order to allocate the limited school seats to the highest-achieving students. In the new system introduced at the turn of the 20th century, applicants were allowed to rank multiple schools in the order of their preferences and take a single unified exam. [^4] As shown later, the defining feature of the centralized system was to allow applicants to list multiple schools, not the use of a single unified exam. Given their preferences and exam scores, each applicant is assigned to a school (or none if unsuccessful) based on a computational algorithm. The algorithm was a mix of the Deferred Acceptance and Immediate Acceptance (Boston) algorithms combined with a meritocracy principle that only assigns the highest-scoring applicants to any school. To our knowledge, this instance is the first nationwide use of any matching algorithm. [^5] The earliest known large-scale use of the Boston algorithm is the assignment of medical residents to hospitals in New York City in the 1920s (Roth, 1990). The oldest known national use of the Deferred Acceptance algorithm is the National Resident Matching Program (NRMP) in the 1950s (Roth, 1984). See Abdulkadiroğlu and Sönmez (2003) for the details of these algorithms in school admission contexts. Furthermore, for reasons detailed below, the government later re-decentralized and re-centralized the system several times, producing multiple natural experiments for studying the consequences of the different systems. [^6] During 1900–1930, there were three periods of centralization (1902–07, 1917–18, and 1926–27). We exploit these bidirectional institutional changes to identify the short- and long-run impacts of the meritocratic centralization.
A stylized theoretical model predicts the short-run effects of the reforms on application behavior and admission outcomes. Consistent with its stated goal, the model confirms that the centralized system produces more meritocratic school seat allocations than the decentralized system. It also predicts that centralization would cause applicants to apply to more selective schools and make more inter-regional applications.
To test these predictions and analyze the effects of meritocratic centralization, we newly collect and digitize data about application and enrollment. The data is based on administrative school records, Government Gazettes, and Ministry of Education Yearbooks. Our data records each student’s regional origin, which is a strong predictor of socioeconomic status both in our empirical setting and other settings like the contemporary US.
We find that meritocratic centralization had large and immediate effects on application and enrollment. First, consistent with the theoretical predictions, centralization caused stark strategic responses in application behavior. In particular, strategic incentives in the centralized system led applicants across the country to more often rank the most selective school first. This effect and the option to apply to multiple schools made students enroll at schools further away from their regional origin. The reform thus made the elite higher education market more meritocratic and regionally integrated.
The gains from centralization were not equally distributed, however. Because high-achieving students were disproportionately located in urban areas (mainly Tokyo), the centralized system caused a greater number of urban applicants to be admitted to schools in rural areas, typically after being rejected by their first choice schools. [^7] We use the nomenclature of “urban” and “rural” schools, but note that “rural” schools were located in regional cities rather than in the countryside. As a result, urban high-achievers crowded out rural applicants; the number of entrants to any national elite school coming from the urban area increased by about 10% during centralization. [^8] It is also empirically true that the centralized system made a greater number of rural applicants apply to and enter urban schools. The centralized system thus increased regional mobility across the country. But their net effects are such that urban high-achievers crowded out rural applicants.
Historical documents suggest that this distributional consequence upset rural schools and communities. Partly as a result of such rural discontents, the government went back and forth between decentralized and centralized systems, finally settling on a decentralized scheme. This series of bidirectional reforms enables us to identify the causal effects of meritocratic centralization more precisely than a usual, single policy change would. [^9] We provide evidence that the reform timings appear exogenous in that they are not correlated with other social and institutional outcomes.
Our main results are about the long-run effects of the centralized system. We start by looking at the persistence of the distributional impacts. The short-run analysis finds that centralization let urban areas disproportionately gain school access relative to rural areas. This result motivates a difference-in-differences analysis that compares long-term career outcomes of urban vs rural-born individuals by each cohort’s exposure to the centralized system. Our long-run career data comes from the Personnel Inquiry Records (PIR) published in 1939, more than thirty years after the first episode of meritocratic centralization in 1902–07. This data provides a list of distinguished individuals (e.g., high-income earners, national medal recipients, high-ranking politicians, corporate executives) along with their biographical information. We provide extensive investigations about the quality of the data. [^10] We find that the data covers a large fraction of the national population of elites (e.g., 53% of the top 0.01% income earners) by comparing our data with national administrative statistics. We also find no systematic variation in the sampling rates across prefectures, consistent with our assumption that sample selection bias is uncorrelated with the prefecture-cohort variation we use.
Distributional effects of meritocratic centralization turn out to be persistent. Almost four decades later, relative to the decentralized system, the centralized system produced a greater number of top income earners, prestigious medal recipients, and other elite professionals who came from urban areas compared to rural areas. The number of urban-born career elites increased by 10–20% for the cohorts exposed to meritocratic centralization. We also obtain suggestive evidence that the centralized system increased the number of career elites living in urban areas as adults in the long run.
We conclude the long-run analysis by examining the overall productive efficiency implication of meritocratic centralization for the whole country. We focus on a specific subgroup of career elites — top bureaucrats. Top bureaucrats were an exemplar category of career elites in our empirical context, as the best and brightest graduates from the top colleges competed to become bureaucrats. Using the comprehensive administrative data of all individuals who passed the national exams to become higher civil officials, we compare the national number of bureaucrats promoted to top ranks between cohorts exposed to the centralized admissions and cohorts exposed to the decentralized admissions. We find that the number of top-ranking bureaucrats in the centralized cohorts was 15% greater than that in the decentralized cohorts, suggesting that centralized cohorts won promotion competitions over close decentralized cohorts. Meritocratic centralization thus produced a larger number of high-quality bureaucratic elites for the whole country.
Overall, our findings reveal an equity-meritocracy tradeoff both in the short and long run. On the one hand, the meritocratic centralization of school admissions achieved the goal of rewarding applicants with higher academic performance in the short run. This meritocracy resulted in a greater number of top bureaucratic elites produced by the exposed cohorts in the long run. On the other hand, this meritocracy came at the equity cost of urban-born high-achieving applicants overwhelming rural-born ones. The design of admission rules therefore affects the production and distribution of highly educated and skilled individuals, which is an important determinant of economic growth and inequality (Glaeser, 2011; Moretti, 2012).
Literature. Our analysis contributes to the empirical literature that uses policy reforms to study the impacts of the design of admission systems (Abdulkadiroğlu et al., 2006; Abdulkadiroğlu, Agarwal and Pathak, 2017; Chen, Jiang and Kesten, 2020; Machado and Szerman, 2021; Grenet, He and Kübler, 2022; Kapor, Karnani and Neilson, 2023). [^11] Other studies measure the effects of selective schools conditional on a particular admission system (Dale and Krueger, 2002; Hastings, Neilson and Zimmerman, 2013; Pop-Eleches and Urquiola, 2013; Deming et al., 2014; Kirkeboen, Leuven and Mogstad, 2016; Clark and Del Bono, 2016; Abdulkadroǧlu et al., 2017; Zimmerman, 2019; Abdulkadiroglu et al., 2021; Michelman, Price and Zimmerman, 2021; Chetty, Deming and Friedman, 2023). While these prior studies focus on the short-run effects (on application, enrollment, academic achievement, and welfare), we estimate the long-run effects by taking advantage of bidirectional, repeated policy changes in history. [^12] With its interest in long-run effects, this paper also relates to studies of the long-term effects of educational reforms (e.g., Duflo, 2001; Meghir and Palme, 2005). These studies focus on the effects of expanding resources (such as school constructions and compulsory education extensions), while we investigate the effects of changing resource allocation mechanisms given a fixed amount of resources. This use of bidirectional policy changes echoes other studies with similar identification strategies (Niederle and Roth, 2003). Our analysis of the meritocracy vs equity tension has some similarities to ongoing policy discussions on affirmative actions, which are surveyed in Arcidiacono and Lovenheim (2016). Recent contributions in this area include Bleemer (2020, 2022), Kapor (2020) and Otero, Barahona and Dobbin (2021). [^13] Kamada and Kojima (2015) and Agarwal (2017) study regional equity in other matching markets.
From a broader historical perspective, this paper relates to the literature that investigates the long-term effects of institutions (Nunn, 2009), particularly resource allocation mechanisms (Bleakley and Ferrie, 2014, 2016). Our analysis is also related to Bai and Jia (2016), who examine political consequences of the abolition of a meritocratic elite recruitment system (civil service exam) in early twentieth-century China.
The next section provides the institutional background. Section 3 examines the short-term effects of meritocratic centralization, while Section 4 analyzes the long-term impacts. Finally, Section 5 summarizes our findings and outlines future directions.
2 Background
One major trend in modern college admissions is a growing degree of centralization with transparent admission criteria. Today, over 30 countries use regionally- or nationally-integrated, single-application and single-offer college admissions. Centralized college admissions are used in all continents except North America (Appendix Figure A.1). These systems have well-specified admission criteria, mixing meritocratic achievement elements (such as GPA and entrance exams), and affirmative action and other priority considerations.
Before the turn of the 20th century, however, no country used such a centralized system. Even today, many countries, including the US, continue to use decentralized systems. Such decentralized schemes tend to come with less transparent criteria for ranking applicants, as illustrated by recent court cases against American universities. Similar observations also apply to K-12 school admissions. How does the centralization of college and school admissions affect students’ application behavior, enrollment outcomes, and future careers? Understanding the costs and benefits of meritocratic centralization is the goal of our paper.
2.1 Bidirectional Admission Reforms
To evaluate the impacts of centralized admission systems, we take advantage of unique historical episodes in early twentieth-century Japan. Since the opening of the country in the mid 19th century, to catch up with Western knowledge, science, and technologies, education reforms became a central part of modernization efforts by the Japanese government. In 1894, the government set up a national higher education system consisting of one Imperial University and five National Higher Schools. By 1908, the system was expanded to four Imperial Universities (Tokyo, Kyoto, Tohoku, and Kyushu) and eight National Higher Schools located across different areas in Japan (First in Tokyo, Second in Sendai, Third in Kyoto, Fourth in Kanazawa, Fifth in Kumamoto, Sixth in Okayama, Seventh in Kagoshima, and Eighth in Nagoya, named after the order of establishment; see Appendix Figure A.2 for their locations). We refer to these eight National Higher Schools as Schools 1–8 for short. [^14] The number of higher education institutions increased after 1918, as the government permitted not only national but also local public and private higher schools and universities. See Appendix Section A.1 and Table A.1 for details. In our empirical analyses, we control for the number of National Higher Schools as well as other characteristics of higher education institutions. We focus on Schools 1–8.
Schools 1–8 served as an exclusive entry point to Imperial Universities, the most prestigious tertiary education. Virtually all graduates of Schools 1–8 were admitted to these universities without further selection well into the 1920s. Imperial University graduates were also partially or wholly exempted from the Higher Civil Service Examinations and other selective national qualification exams to become high-ranking administrators, diplomats, judges, and physicians. As a result, entering Schools 1–8 was considered as a passport into the elite class. Indeed, Schools 1–8 produced highly distinguished and influential individuals, including several Prime Ministers, Nobel Laureates, world-leading mathematicians, renowned novelists, and founders of global companies like Toyota. To apply to these schools, one must be male aged 17 or older and have completed a five-year middle school. [^15] The eligibility was changed in 1919 to males aged 16 or older who have completed the fourth year of middle school. Schools 1–8 admitted fewer than 2,300 students each year throughout 1900–1930, constituting less than 0.5% of the cohort of males aged 17. More descriptive statistics are presented in Appendix Table A.2.
Among Schools 1-8, School 1 in Tokyo was considered the most prestigious due to its location in the capital and its geographical proximity to Tokyo Imperial University (today’s University of Tokyo). The next most prestigious was School 3 in Kyoto, the ancient capital of Japan. By contrast, located in a remote southwest region, Schools 5 and 7 were considered the least prestigious among all schools. Consequently, the schools differed substantially in their popularity and selectiveness. [^16] Tuition was mostly uniform across Schools 1-8.
The admission to Schools 1–8 was merit-based and determined by annual entrance exams. Initially, the government let each school administer its own exam and admissions. Schools 1–8 typically held their exams on the same day so that each applicant could only apply to one school. Following the convention in the literature (Che and Koh, 2016; Hafalir et al., 2018), we call this system “decentralized admissions” or Dapp for short. The single choice aspect of Dapp captures an essential feature of decentralization, which incentivizes each applicant to self-select into an appropriate school by comparing the selectivity of schools with his own standing.
Under Dapp, however, many high-achieving students were rejected by popular schools, while lower-achieving students entered less popular schools. For the government, who wanted to send the most promising students to selective higher education, the decentralized system appeared inefficient. [^17] Applicants who failed to enter Schools 1-8 chose either (1) to retake the exam in the following year, (2) to enter local public or private higher educational institutions, or (3) to give up receiving higher education. The Education Minister criticized the decentralized system as follows:
[Under the decentralized system] many applicants are rejected by Schools 1 and 3 [in Tokyo and Kyoto], which attract a large number of high ability applicants, despite the fact that these applicants have superior academic performance to that of applicants admitted to other rural schools. (…) As a result, hundreds of applicants with sufficiently high academic ability to enter rural schools are idly wasting another year [to retake the exam]. This is not only a pity for them, but also a loss for the country.” (Educational Review No.1146, p.21, published in February 15, 1917)
To solve this problem, in 1901, the government first asked all schools to unify their entrance exams to a single one, while maintaining decentralized admission decisions. In 1902, the government launched a centralized system in which applicants apply for multiple schools, submit their preferences over schools, and take the unified exam. Based on their exam scores and preferences, applicants were then assigned to a school (or no school if unsuccessful) by an assignment algorithm announced ex ante. Each year, the government announced application procedures in April, three months before the exam in July, as a public notice in Government Gazette. We call this system “centralized admissions” or Capp for short.
The centralized assignment algorithm operates as follows. [^18] This is the assignment algorithm used in 1917. See Appendix Figure A.3 for a reprint of the original assignment algorithm published in April 1917. Strictly speaking, the centralized assignment algorithm was introduced first in 1902, slightly modified in 1903 and 1917, and revised further in 1926.
  1. (1)
    In the order of exam scores, select the same number of applicants as the sum of all schools’ capacities. In the case of a tie, decide by lottery.
  2. (2)
    For applicants selected in (1), in the order of exam scores, assign each applicant to the school of his first choice. In the case of a tie, decide by lottery.
  3. (3)
    For applicants selected in (1) and for whom the school of his first choice is already filled at the end of (2), in the order of exam scores, assign each applicant to the school of his second choice. In the case of a tie, decide by lottery.
  4. (4)
    For applicants selected in (1) and for whom the school of his second choice is already filled at the end of (3), then assign each applicant to the school of his third choice or below, repeating the same procedure as (3).
  5. (5)
    If all the schools in an applicant’s preference list are filled, then the applicant is not admitted to any school.
Written more than a century ago in natural language, the rule description is mathematically precise. The above algorithm imposes meritocracy up front in which only top-scoring applicants were considered for admission regardless of their preferences (Step (1)). Capp is meritocratic in the sense that the test score distribution among assigned students under Capp first order stochastically dominates that under any other mechanisms, including Dapp. This step selects only applicants who would be admitted by any school under the Deferred Acceptance algorithm (Abdulkadiroğlu and Sönmez, 2003), one of the most widely used algorithms in today’s college and selective K-12 admissions (Appendix B). These applicants are then assigned to one of Schools 1–8 using the Immediate Acceptance (Boston) algorithm (Steps (2) to (4)). This algorithm is therefore a variant of the Immediate Acceptance algorithm with a meritocracy constraint, making it closer to the Deferred Acceptance algorithm. To the best of our knowledge, this is the world’s first recorded nation-wide use of any assignment algorithm.
This institutional innovation was short-lived, however. Due to political and administrative reasons detailed below, the government switched back to Dapp (with a unified exam) in 1908. The government then continued to oscillate between decentralization and centralization, reintroducing Capp in 1917, moving back to Dapp (with a unified exam) in 1919, reinstituting Capp (with major modifications of allowing applicants to list at most two schools) in 1926, and finally settling down to Dapp (with separate exams) in 1928. In a space of thirty years, there were three periods of centralized admissions: first in 1902–1907, next in 1917–1918, and finally in 1926–1927.191919For details of the evolution of the admission system, see Appendix Table A.1.
We exploit the series of bidirectional policy changes to examine the impacts of centralization. Importantly, it was difficult for applicants to anticipate the exact timing of these policy changes, as the admission policy was decided and announced only a few months in advance. This suggests that the reform timing is close to exogenous, as we further explore in Section 3.5.
2.2 Predicted Impacts
To guide our empirical analysis, we first predict effects of centralization by using historical documents and theoretical insights. Historical documents suggest that the repeated policy changes resulted from intense bargaining between the Ministry of Education in the central government and the Council of School Principals representing local interests. [^20] Detailed accounts are provided in Moriguchi (2021), pp.195-199. While the Ministry of Education emphasized the importance of using centralized admissions to enroll the best and brightest to the national higher education, the Council of School Principals was opposed to centralization, pointing out several adverse consequences for rural schools and their local communities.
First, according to the Council of School Principals, the centralized system encourages applicants to be more risk-taking and rank the most competitive and prestigious urban school as the first choice. This is because the centralized system provides applicants with options to apply for lower choice schools as insurance.
Hypothesis 1. Under the centralized system, more applicants in all areas rank most prestigious School 1 as their first choice.
Second, Hypothesis 1 means that the centralized system made a greater number of rural applicants apply to and enter urban schools. In addition, the centralization allowed high-achieving applicants in urban areas to enter rural schools even if they failed at the most selective urban school. This, in turn, often resulted in the rejection of lower-achieving rural applicants who would have been able to enter their local schools under decentralized admissions. These effects would jointly increase the mobility of students across the country.
Hypothesis 2. Under the centralized system, a smaller share of each school’s seats are allocated to its local students.
The Council of School Principals also complained that the centralized system lets many reluctant and unmotivated students enter rural schools as fallback options, which lowered school morale.
Third, by the meritocracy constraint embedded in the centralized system, it allocated a higher fraction of seats at Schools 1–8 to high-achieving applicants. After reviewing the admission results of 1917 (the first year of the second centralization period), the Council of School Principals found that the number of rural middle-school graduates not admitted to any of Schools 1–8 increased considerably compared to the previous three years of decentralized admissions. This is presumably because high-achieving applicants were located disproportionately in urban areas.
Hypothesis 3. Under the centralized system, a larger share of all school seats in Schools 1–8 are allocated to urban-born applicants.
Online Appendix B formalizes these hypotheses as theoretical predictions in particular models. A noted historian summarizes these observations as follows (Takeuchi, 2011, p.121):
Urban applicants ‘overwhelm’ rural applicants by applying for rural schools as fallback options. Urban applicants rob rural applicants of opportunities that were once open to them. This ruins the meaning of building national higher schools across the nation.
This upset local communities of rural schools, especially because they contributed land and other resources to the construction of rural schools (Takeuchi, 2011, p.56 and pp.106-107). [^21] An additional concern against the centralized system was the administrative cost of implementing it in the absence of modern computers and photocopying technologies. Both the Ministry of Education officials and the school principals repeatedly pointed out the difficulty of grading thousands of exam sheets by hand in a short period of time and assigning applicants to schools according to the algorithm (Takeuchi, 2011, pp.118-119; Moriguchi, 2021, p.199). The school principals also deemed centralization as an intrusion on their power and autonomy.
These documents highlight a potential meritocracy-equity tradeoff of centralization. On the one hand, the centralized admissions made the school seat allocation more meritocratic, enabling high-achieving students to enter one of the elite schools even if they failed at the most selective one. On the other hand, this meritocracy might have come at the expense of equal regional access to higher education, whereby high-achieving urban applicants dominated rural applicants. Below we test the above hypotheses as well as their persistence in the long run.
3 Short-run Impacts
3.1 Data
To analyze short-run effects of the centralized admission system, we assemble and digitalize data on application and enrollment. First, we collect data on the number of applicants by school of their first choice from 1900 to 1930, using Ministry of Education Yearbook and other sources. For these two years (1916 and 1917), we also collect data on the number of applicants by school of their first-choice and by prefecture of their middle school. Appendix Section A.3.1 provides detailed explanations for data sources and definitions.
Second, we compile data on the number of entrants (first-year students) by school and by their birth prefecture from 1900 to 1930, using Student Registers of Schools 1–8. Finally, to control for the size of potential applicants and the number of competing schools, we collect data on the number of middle-school graduates by prefecture of their middle school, as well as the number of other Higher Schools (established in addition to Schools 1–8 starting in 1919) by prefecture from 1900 to 1930. See Appendix Section A.3.2 for detailed descriptions of the data and sample. Descriptive statistics of main variables are reported in Appendix Table A.2.
3.2 Strategic Responses by Applicants
As an immediate effect, switching back and forth between the centralized and the decentralized admission systems causes stark strategic responses in application behavior. Figure 1(a) shows that the three periods of Capp are associated with sharp increases in the share of applicants who chose the most selective School 1 in Tokyo as their first choice, as predicted.
To observe regional variations, we examine how the propensity of applicants to rank School 1 as their first choice changes between 1916 (under Dapp) and 1917 (under Capp). The results in Appendix Table A.3 show that the meritocratic centralization induced applicants in all regions to rank the most prestigious school first and to make more long-distance applications. As a result, the competition to enter School 1 became even more intense under Capp (Appendix Figure A.4). As a result, under Capp, only a small fraction of first-choice applicants were admitted to School 1, producing many high-scoring applicants who were rejected by School 1 and then assigned to their second or lower-choice school.
3.3 Regional Mobility in Enrollment
The changes in the mechanism and behavior influence enrollment outcomes. To measure the geographical mobility of entrants, we compute the “enrollment distance” defined as the direct distance between an entrant’s birth prefecture and the school he entered. As Figure 1(b) shows, the centralized system is associated with a sharp and discontinuous increase in enrollment distance. [^22] The effect is especially stark in the first two periods of Capp. The third period of Capp in 1926–27 was qualitatively different from that in the first and second periods. In the third period of Capp, schools were divided into two groups and applicants were allowed to choose and rank at most two schools (one school per group) in 1926–27. This increase in regional mobility is driven by a sharp reduction in the number of “local entrants” defined as entrants who entered a school in (or near) their birth prefecture. To show this, we estimate the following regression for each school $s$ separately:
$\begin{split}Y_{pt}^{s}=&\ \beta_{1}^{s}\times Centralized_{t}\times 1\{\text{school }s\text{ is located in prefecture }p\}\\ &+\beta_{2}^{s}\times Centralized_{t}\times 1\{\text{school }s\text{ is 1-100 km away from prefecture }p\}\\ &+\alpha^{s}X_{pt}+\gamma_{t}^{s}+\gamma_{p}^{s}+\epsilon_{pt}^{s},\end{split}$ (1)
[…]
As shown in Panel (a) of Table 1, Capp reduces the number of local entrants born in the school’s prefecture across the country. The coefficients $\beta_{1}^{s}$ are significantly negative for all schools. Column (1) shows that the number of School 1 entrants born in Tokyo prefecture declined by about 25%. Schools 4–7 experienced reductions in the number of entrants born not only from the school’s prefecture but also from surrounding prefectures. In other words, centralization weakened the local monopoly power of each school by creating a national market for higher education, consistent with our theoretical prediction in Appendix B (Proposition 3).
3.4 Meritocracy vs Equal Regional Access
Which regions gained more school seats under the centralized system? We now investigate how centralization altered the geographical allocation of school seats. Figure 2(a) plots the change in the total number of entrants to Schools 1–8 from Dapp to Capp by their birth prefecture, where the darker blue color indicates the greater decline and the darker red color indicates the greater increase. The figure shows that Tokyo prefecture (where School 1 was located) and its surrounding area gained school seats under Capp, while most of western and northern prefectures lost school seats.
To see this quantitatively, we regress the number of entrants to Schools 1–8 in year $t$ born in prefecture $p$ on the interaction terms between $Centralized_{t}$ and each of the following two indicator variables: the indicator that takes 1 if prefecture $p$ was Tokyo and the indicator that takes 1 if prefecture $p$ was 1-100 km away from Tokyo. To control for the decline of local entrants, we also add the interaction terms between $Centralized_{t}$ and each of the following two indicator variables: the indicator that takes 1 if any of Schools 1–8 was located in prefecture $p$ and the indicator that takes 1 if any of Schools 1–8 was located 1-100 km away from prefecture $p$. As in Equation (1), we also control for observable characteristics of prefecture $p$ and year $t$ and include year and prefecture fixed effects.
Column (1) in Table 1 Panel (b) shows the number of entrants born in Tokyo prefecture increased by 13% per year under Capp relative to its mean under Dapp. The increase in the number of entrants is even higher for its surrounding area; the number of entrants born in prefectures 1-100 km away from Tokyo increased by 45% per prefecture per year under Capp relative to its mean under Dapp. [^23] One possible reason for this concentration of the effect in surrounding prefectures is that applicants from this area might be marginal applicants who would not be admitted under Dapp but would be admitted under Capp. On the other hand, many high-scoring applicants from the Tokyo prefecture might be admitted even under Dapp. Consistent with this hypothesis, the number of entrants from Tokyo (201) is much larger than the number of entrants from surrounding prefectures (26 per prefecture) under Dapp.
Motivated by these observations, we define “Tokyo area” as the set of prefectures located within 100 km from Tokyo (see Appendix Table A.3 for its location). Figure 2(b) depicts the time evolution of the share of entrants to Schools 1–8 who were born in the Tokyo area. Consistent with the above results, the share of Tokyo-area born entrants rose during the periods of Capp.
To see which schools received more Tokyo-area born entrants under Capp, we regress the number of entrants to each school $s$ in year $t$ born in prefecture $p$ on the interaction terms between $Centralized_{t}$ and the variables indicating whether entrants’ birth prefecture $p$ is Tokyo or near Tokyo. We include the same variables used in Equation (1) to control for school locations and other prefecture and year characteristics. Under Capp, a greater number of Tokyo-area-born students entered less selective rural schools, as shown in Columns (2)–(9) in Table 1 Panel (b). [^24] The results remain almost the same when we additionally control for observable prefecture characteristics (see Appendix Table A.4).
In summary, the net effect of Capp is such that the increased inter-regional applications caused high-achieving students in the Tokyo area to crowd out lower-achieving, rural-born students from their local schools. This evidence is consistent with the historian’s observation (Takeuchi (2011) cited in Section 2.2) that, as a result of the centralization, urban applicants mainly from Tokyo dominated rural applicants in entering national elite schools.
The Tokyo area shares common characteristics with what is generally considered urban. Prefectures in the Tokyo area had a larger population, higher income per capita, and better educational infrastructure (indicated by a greater number of middle-school graduates eligible for higher education) compared to other prefectures (Appendix Table A.5). In Appendix Table A.6, we use these urban characteristics instead of the Tokyo area indicator in the same specification as Column (1) in Panel (b) of Table 1. The result shows that these urban characteristics, especially educational infrastructure, positively predict the areas which produced a greater number of entrants under Capp. Therefore, a greater number of high achieving students in the Tokyo area may explain why the Tokyo area gained more school seats under centralization.
Students’ heterogenous preferences across areas can be another reason why the Tokyo area gained more school seats under the centralized system. Students prefer their local school for geographical and cultural proximity. The local school for students born in the Tokyo area is the most competitive School 1. Holding academic achievement constant, during the decentralization, students born in the Tokyo area are more likely to apply to School 1 and therefore fail to enter any school (even though they might have been able to enter Schools 2–8 had they applied). This hypothesis is consistent with Appendix Table A.6, which shows that the share of applicants to School 1 during the decentralization is associated with a greater number of entrants to Schools 1–8 during the centralization.
In summary, our analysis suggests that a combination of better educational infrastructure and strong preference for School 1 explains why the Tokyo area gained more school seats under the centralized system.
3.5 Other Institutional Changes
We discuss potential threats to our empirical analysis. In particular, if changes in other institutional and policy factors were correlated with the admission reforms, it could influence application behavior and enrollment outcomes, explaining our finding that more applicants applied for School 1 during centralization. Such other institutional factors include simultaneous reforms in middle schools, the total number of applicants, capacities at Schools 1–8, and the capacity of School 1 relative to the capacities of other schools.
We investigate these concerns and confirm that centralization periods are not correlated with time-series changes in the following variables (Columns (1)–(6) in Appendix Table A.7): the number of middle school graduates, the national number of applicants as well as the level of competitiveness (measured by the number of entrants divided by the number of applicants), the total number of entrants to Schools 1–8, the share of entrants to School 1 in all entrants, the probability of unsuccessful applicants retaking the exam in subsequent years, and the average age of entrants.
A potential concern with the above time-series analysis is that the insignificant results in Appendix Table A.7 may be due to a small sample size (around 30). Yet, using the same empirical specification, we find that centralization is significantly correlated with our main outcome variables (the share of applicants to School 1, the enrollment distance, and the share of entrants who were born in the Tokyo area), as shown in Columns (7)–(9) of Appendix Table A.7. Taken together, these results suggest that our findings are unlikely to be driven by institutional changes other than the school admission reforms.
Finally, the centralization reform introduced not only the meritocratic assignment algorithm, but also the unified entrance exam that applicants could take at any school location (see Appendix Table A.1). As such, the estimated impacts of centralization may be confounded by the unification of entrance exams and more flexible exam location choices. To investigate this issue, we analyze how key outcomes change from 1900 to 1901, during which the government also introduced a single entrance exam that applicants were allowed to take anywhere while the assignment method remained unchanged (i.e., decentralized). Figures 1(a) and 1(b) show that this institutional change from 1900 to 1901 induced little changes in application and enrollment patterns. The estimated impacts of centralization are therefore likely due to the meritocratic assignment algorithm rather than the changes in exam contents and locations.
4 Long-run Impacts
To assess long-run effects of the meritocratic centralization, we provide two sets of empirical analysis. First, we investigate the persistence of the distributional effect of centralization that we document in the short-run analysis. We do so by a difference-in-differences strategy comparing labor-market outcomes of urban- and rural-born individuals across birth cohorts with differential exposure to the centralized admissions. Our analysis shows that the centralization increased both (a) the number of career elites born in urban areas and (b) the number of career elites living in urban areas as adults. The admission reforms thus influence the regional origins and destinations of career elites.
Second, we explore the national productive efficiency of the two admission systems. We find that cohorts more intensely exposed to the centralized admissions produced a larger national number of civil officials who were promoted to top ranks. This finding indicates that the centralization boosted the national production of career elites, at least for the specific important category of top-ranking government officials.
4.1 Regional Distribution of Career Elites
Personnel Inquiry Records Data
To analyze long-run effects of the centralization on students’ career outcomes, we use the Personnel Inquiry Records (PIR) published in 1939.252525Japan went into the military regime with the promulgation of the National General Mobility Act in 1938. Although the military government began massive social interventions, their effect was relatively small in 1939 (Moriguchi and Saez (2008)). In addition, most of the information in PIR (1939) is based on 1938 data. The PIR is an equivalent of Who’s Who, which compiles a selective list of socially distinguished individuals (including high-income earners, top business managers, elite professionals, and high-ranking public servants) and provides their biographical information. In total, the 1939 PIR lists 55,742 individuals or 0.15% of the adult Japanese population of that time. We detail the selection process in Appendix Section A.4.1.
To capture the effects of the first period of the centralized admission system in 1902–1907, we use the cohorts born in 1880–1894, who turned 17 years old (the age eligible for application) in 1897–1911. The cohorts born in 1880–1894 were 45 to 59 years old in 1939. The number of individuals listed in the PIR in each of these cohorts is about 1,800. We use the following information from the PIR data for each individual: full name, birth year, birth prefecture, prefecture of residence, final education, occupation titles and organization names, the medal for merit and the court rank awarded (if any), and the amount of national income tax and business tax paid (if any).
We define the following groups of elites as a subset of PIR-listed individuals: (1) the top 0.01% and 0.05% income earners according to the national income distribution, (2) prestigious medal recipients (civilians who received either the medal of the fifth order of merit or above, or the court rank of the junior fifth rank or above), (3) corporate executives (individuals who hold an executive position in a corporation and pay a positive amount of income or business tax), (4) top politicians and bureaucrats (individuals whose occupation is either Imperial Diet member, Cabinet member, or high-ranking central government official), and (5) Imperial University professors or associate professors. Appendix Section A.4.1 and Table A.8 provide detailed definitions and descriptions of each group. These categories encompass economic, social, cultural, political, and academic definitions of career elites.
We use this data to count the number of elites in each group defined by birth prefecture and birth cohort. These counts allow us to conduct a difference-in-differences analysis that compares long-term career outcomes of urban- and rural-born individuals by each cohort’s exposure to the centralized admission system. Descriptive statistics of main variables are summarized in Appendix Table A.2.
Assessing the Coverage and Bias of the PIR Data
Since our PIR data is not exhaustive administrative data, we are concerned about potential sample selection bias. For the top income earners and Imperial University professors, we can compute the exact sampling rates by comparing the number of individuals in our data against complete counts reported in government statistics. We find that the sampling rates are decent even by modern standards: 70% for Imperial University professors, and 53% and 39% for the top 0.01% and 0.05% income earners, respectively. Consistent with the nature of our data, which lists only distinguished individuals, the sampling rates increase with the income level (see Appendix Figure A.5).
Sample selection bias becomes a problem for our difference-in-differences analysis only if the difference in sampling rates between urban and rural areas changes with cohorts’ exposure to the centralized admission system. To assess this possibility, we examine the prefecture-level sampling rates for the top income earners. As Appendix Figure A.6 shows, the number of high income earners in our data and the complete count from tax statistics are highly correlated at the prefecture level, with similar sampling rates across prefectures. This result provides further support for the quality of our data. Even so, one potential concern is that Imperial University graduates might have a higher likelihood of being sampled by the PIR even after controlling for the income level. However, we find no positive correlation between the sampling rates of top income earners and the numbers of Imperial University graduates across prefectures (see Appendix Table A.9). This series of findings suggests that possible sample selection bias in the PIR data is unlikely to drive our empirical results.
Urban-Rural Disparity in Producing Career Elites
We estimate the long-run impacts of the centralized admission system by conducting a difference-in-differences analysis by birth cohorts and birth areas. The key idea behind our empirical strategy is that applicants born in the Tokyo area should experience a greater gain in entering Schools 1–8 under Capp relative to Dapp, since the centralized system is designed to be more meritocratic and high-achieving students are disproportionately located in urban areas. This expectation is confirmed in Figure 2 and Table 1. We exploit this differential gain in school access to compare the career outcomes of individuals born inside and outside the Tokyo area by the cohort’s exposure to Capp. If admission to Schools 1–8 increases one’s chance of becoming a career elite, we should observe a greater number of elites born inside the Tokyo area for the cohorts exposed to Capp. We estimate a difference-in-differences specification as follows:
$Y_{pt}=\beta\times Centralized_{t}\times Tokyo\_area_{p}+\gamma_{p}+\gamma_{t}+\epsilon_{pt},$
where $Y_{pt}$ is the number of elites born in cohort $t$ and prefecture $p$. $Centralized_{t}$ is a measure of cohort $t$’s exposure to Capp, which is, in the baseline specification, a binary indicator that cohort $t$ turned 17 during Capp (1902–1907). $Tokyo\_area_{p}$ is the indicator that takes 1 if prefecture $p$ is in the Tokyo area. $\gamma_{p}$ and $\gamma_{t}$ are prefecture and cohort fixed effects. To allow for serial correlation of $\epsilon_{pt}$ within prefecture over time, we cluster the standard errors at the prefecture level in our baseline specification. [^26] Bertrand, Duflo and Mullainathan (2004) evaluate approaches to deal with serial correlation within each cross-sectional unit in panel data. They suggest that clustering the standard errors on each cross-section unit performs well in settings with 50 or more cross-section units, as in our setting. In addition, we report the results of clustering at cohort level, which are estimated by wild cluster bootstrap (Cameron and Miller, 2015; Roodman et al., 2019) due to the small number of clusters (15 cohorts).
The above regression defines $Centralized_{t}$ to be a binary indicator, as the simplest proxy for the intensity of exposure to Capp. In reality, however, a nontrivial number of unsuccessful applicants retook the exam at age 18 and beyond. As a result, the cohorts who turned age 17 in 1899–1901 were partially and increasingly exposed to Capp (as some of them took the exam in 1902). The cohorts who turned age 17 in 1902–1904 were fully exposed to Capp. The cohorts who turned age 17 in 1905–1907 were partially and decreasingly exposed to Capp (as some of them took the exam in 1908). Finally, the intensity of exposure drops to zero for the cohorts who turned age 17 in 1908.272727The intensity of exposure to Capp is estimated to be 0.00, 0.02, 0.09, 0.36, 1.00, 0.98, 0.91, 0.64, and 0.00 for the cohort who turned age 17 in 1896–98, 1899, 1900, 1901, 1902–04, 1905, 1906, 1907, and 1908–11, respectively. See Appendix Section A.4.1 for data and methods. In the following visual analysis, we explicitly incorporate cohort’s intensity of exposure to Capp by coloring cohorts according to their intensity of exposure to Capp.
We first check whether the number of Imperial University graduates born inside the Tokyo area increased for the cohorts exposed to Capp. Since all Schools 1–8 graduates were automatically admitted to an Imperial University during this period, the areas that produced more Schools 1–8 entrants should produce more Imperial University graduates. Figure 3 confirms this expectation. Panel (a) compares the average number of Imperial University graduates who were born in prefectures inside and outside the Tokyo area by cohorts (represented by their birth year plus 17 on the horizontal axis). Figure 3 (b) shows that the urban-rural difference in the number of Imperial University graduates rises as the intensity of exposure to Capp increases and falls sharply when Capp ends in 1908. Column (1) in Table 2 shows that the estimate of $\beta$ in the above regression is positive and statistically significant.
Our main results are presented in Figure 3 (c)–(f) and Table 2 Columns (2)–(7). Figure 3 (c)–(f) show difference-in-differences plots that compare the number of elites (the top 0.05% income earners and medal recipients) who were born inside and outside the Tokyo area by the cohort’s exposure to Capp. For both elite categories, the plots show that the difference between the Tokyo area and the rest grows as the intensity of exposure to Capp increases, and then drops sharply when Capp ends. Appendix Figure A.7 shows qualitatively similar difference-in-differences plots for the top 0.01% income earners, corporate executives, top politicians and bureaucrats, and Imperial University professors.
Table 2 Columns (2)–(7) show that the long-run effects of Capp are economically and statistically significant. For the cohorts exposed to Capp, the number of career elites born inside the Tokyo area (compared to those born outside the Tokyo area) increases by 34% for the top 0.01% income earners, 24% for the top 0.05% income earners, 36% for medal recipients, 14% for corporate executives, 50% for top politicians and bureaucrats, and 47% for Imperial University professors (in Panel B). [^28] In Appendix Table A.10, we also examine other groups of elite professionals and obtain similar results for scholars, engineers, and physicians. In Panel A, we control only for cohort and prefecture fixed effects. In Panel B, we additionally control for time- and cohort-varying prefecture characteristics. [^29] Cohort birth population, the number of primary schools, the number of middle school graduates, and prefecture-level GDP. See Appendix Section A.4.1 for variable definitions and data sources. The coefficients fall slightly in magnitude after adding control variables, but remain sizable.
Table 2 Panel C shows that the effects are symmetric with respect to the direction of the admission reforms, i.e., the change from Dapp to Capp and the change from Capp to Dapp produce similar effects of the opposite sign. In Table 2 Panel D, we replace the centralization dummy by the cohort’s intensity of exposure to Capp. The results remain qualitatively the same as the baseline results.
These results suggest that almost four decades after its implementation, Capp had lasting effects on the career trajectories of students. The above results are robust to alternative specifications. First, the analysis in Table 2 Panel D assumes that the cohort’s intensity to exposure to Capp is exogenous and the same across years, which may be a strong assumption. However, even when we drop the cohorts who are exposed to both Capp and Dapp (i.e., cohorts who became age 17 in 1901 and 1907) from the sample, we still find qualitatively the same results (Appendix Table A.11). Second, we test if the assumption of parallel pre-event trends holds. Appendix Table A.12 verifies that the differences in pre-event trends between the areas of comparison are small and mostly insignificant.
Another potential threat to our identification strategy is that there may be some age-specific trends in the number of elites that covary with the cohort-region variation we use. Specifically, the number of elites selected in the 1939 long-term data peaks at around the cohort who were 51 years old in 1939 (corresponding to the cohort who turned age 17 in 1905) and gradually falls for younger and older cohorts, suggesting that there are certain ages at which individuals are more likely to be listed in the long-term data. Such age effects may generate different trends in the number of elites born in the Tokyo and other areas, due, for example, to differences in population size across these areas. To address this concern, we use an earlier edition of the PIR published in 1934, construct the prefecture-cohort level data for the same cohorts used in our main analysis (but observed 5 years earlier), and conduct similar regression analyses. The results in Appendix Table A.13 confirm that our key results remain qualitatively the same even when we use the 1934 PIR data.
Finally, we conduct placebo tests to examine if the results are driven by other factors such as the sample selection of the PIR or changes in cohort populations. The urban-rural difference in the cohort’s birth populations does not change significantly with the cohort’s exposure to Capp, as shown in Appendix Table A.14. As an additional placebo test, we also look at unrelated career outcomes. Among the distinguished individuals listed in the PIR, we expect that landlords (defined in Appendix Section A.4.1) are least likely to be affected by the introduction of Capp as receiving higher education was not a typical pathway to become a landlord. As shown in Appendix Table A.14, the estimated effect of Capp on the number of landlords is small and statistically insignificant. These findings further support our conclusion that the meritocratic centralization enlarged the urban-rural gap in the capacity to produce future career elites.
Geographical Destinations of Career Elites
Having established that Capp affects the geographic origins of elites, we now ask how it affects their geographic destinations. While the former is about regional inequality in educational opportunities, the latter is about regional inequality in the supply of highly skilled human capital, which potentially affects economic growth and inequality.
We first examine if the centralization-induced increase in inter-regional mobility in the short run boosted the geographical mobility of elites in the long run. Surprisingly, it did not: The urban-rural difference in the fraction of elites whose prefectures of residence differ from their birth prefectures did not increase under Capp, as shown in Appendix Table A.14. We find similar results when we use the distance between an elite’s birth prefecture and his prefecture of residence as an alternative measure of long-run mobility. This result suggests that, even though a greater number of students born in the Tokyo area entered rural schools under Capp, most of them returned to the Tokyo area when pursuing their careers
If the greater number of Tokyo-area born students admitted to rural schools under Capp returned to the Tokyo area eventually for their subsequent careers, we should observe a greater number of elites living in the Tokyo area for the cohorts exposed to Capp. To test this hypothesis, we redefine the outcome variables by changing the prefecture ($p$) from birth prefecture to prefecture of residence and estimate the same equation.
Table 3 shows large positive effects of Capp on the urban-rural gap in the number of elite residents. For the cohorts exposed to Capp relative to Dapp, the number of elites living in the Tokyo area in their middle age (compared to those living outside the Tokyo area) increases by 28% for the top 0.01% income earners, 23% for the top 0.05% income earners, 29% for medal recipients, 22% for corporate executives, 20% for top politicians and bureaucrats, and 50% for Imperial University professors (Panel B). These results suggest that the meritocratic centralization intensified the concentration of career elites in urban areas in the long run.
4.2 National Production of Career Elites
The above analysis examines the distributional consequence of the centralized admissions. A complementary problem is its overall impact for the whole country. To study it, we now turn to whether the meritocratic centralization improved the national production of career elites in the long run. We focus on a specific group of elites — higher civil officials — for whom we have complete-count data from administrative records. Top-ranking civil officials were also the most relevant group of career elites; in our empirical context, higher civil service was considered to be one of the most prestigious jobs due to their political and legislative influence. As a result, high-achieving students at the top universities competed to get jobs at the Ministry of Finance and other selective ministries (Shimizu, 2019). We investigate whether cohorts exposed to Capp produced a greater national number of top-ranking civil officials compared to cohorts exposed to Dapp.
Higher Civil Officials Data
Our main data source is the list of individuals who passed the Higher Civil Service Examinations (HCSE) and their biographical information (Hata, 1981). The HCSE were selective national qualification exams held annually from 1894 to 1947. The 1893 ordinance required all individuals to pass the HCSE for appointment in the administrative division of higher civil service with some exceptions for special appointments (Spaulding, 1967, Chapter 25; Shimizu, 2019, Chapter 5). We digitized the information of all individuals who passed the administrative division of the HCSE in 1894–1941, including their full name, education, year of university graduation, year of passing the exam, starting position, final position, year of retirement, and other notable positions held. Because education includes not only final but also the second to final education, we observe both university and higher school (if applicable).
In the following analysis, we define “top-ranking officials” as higher civil officials who were internally promoted to one of the top three ranks by the end of their career. In the prewar Japanese bureaucracy system, the higher civil service refers to the top ten ranks of national government offices in the administrative, judicial, and diplomatic divisions. Within the higher civil service, the top three ranks were distinctively called “imperial appointees.” The first rank consisted of minister level positions, while the second and third ranks consisted of vice minister level positions such as vice minister, director general, bureau chief, and prefectural governor. Appendix A.4.2 provides more details.
We count the number of top-ranking officials and the number of individuals who passed the administrative division of the HCSE (hereafter “exam passers”) produced by cohort. Here cohort is defined by “the year of entering a higher school or its equivalent.” Since we only observe the year of university graduation in the above data, we estimate “the year of entering a higher school or its equivalent” using the method described in Appendix A.4.2. We first find the exact year of entering a higher school for all top-ranking officials who graduated from Schools 1–8 by searching their names in Student Registers of Schools 1–8. Using this information, we then estimate “the year of entering a higher school or its equivalent” for the rest of higher civil officials.
Out of 6,255 exam passers in our dataset, 55.8% are Schools 1–8 graduates and 15.7% are top-ranking officials. Among 982 top-ranking officials, 71.4% are Schools 1–8 graduates. More descriptive statistics are shown in Appendix Table A.2.
Effect on Higher Civil Service Exams Passers
As an intermediate outcome, we first examine the impacts of the centralized admission system on the number of individuals who passed the selective HCSE. We divide exam passers into three mutually exclusive subgroups: (a) those who graduated from School 1, (b) those who graduated from Schools 2–8, and (c) those who are not Schools 1–8 graduates. For each subgroup, we count the number of exam passers by cohort. We estimate the following equation for the entire group and for each subgroup:
$Y_{t}=\theta Centralized_{t}+\xi_{1}X_{t}+\xi_{2}Trend_{t}+\xi_{2}Trend_{t}^{2}+\omega_{t},$
where $Y_{t}$ is the number of exam passers in a given group in cohort $t$ (defined by the year of entering a higher school or its equivalent), and $Centralized_{t}$ is the indicator that takes 1 if cohort $t$ entered a higher school or its equivalent during Capp. For a subgroup regression, we control for the total number of exam passers in cohort $t$ (denoted by $X_{t}$). We also control for a quadratic time trend where $Trend_{t}$ is the number of years since 1897.
Capp changed the composition of exam passers. In Table 4 Panel (a), Column (3) shows that Capp increased the number of exam passers who graduated from Schools 2–8 by 23%, while Column (4) indicates that Capp reduced the number of exam passers who are not Schools 1–8 graduates by 13%. By contrast, Capp did not have significant effect on the total number of exam passers or the number of exam passers who graduated from School 1 (Columns (1) and (2)). In other words, the centralization had little impact on the total number of exam passers, but had a major impact on the composition of exam passers, and its positive effect was concentrated on Schools 2–8 graduates. This finding is consistent with the result of our short-run analysis that the centralization led to a large increase in Tokyo-area-born high-achieving entrants to Schools 2–8 (Table 1 Columns (3)–(9)), which likely improved the academic standing of these schools.
Effect on Top-Ranking Higher Civil Officials
Finally, we analyze the impact of the centralization on the national production of top-ranking officials. Our main outcome variable is the number of higher civil officials who were internally promoted to the top three ranks by the end of their career. We count the number of top-ranking officials by cohort and run the same regression as above for the entire group and for each of the following subgroups: (a) those who graduated from School 1, (b) those who graduated from Schools 2–8, and (c) those who are not Schools 1–8 graduates. For all regressions, we control for the total number of exam passers from the cohort.
Meritocratic centralization produced a larger number of high-quality bureaucratic elites for the whole country. Table 4 Panel (b) Column (1) shows that the total number of top-ranking officials increased for Capp cohorts by 15%, with high statistical significance. This result is consistent with the idea that the centralized system (which assigned top-scoring students to Schools 1-8) increased the quality of Schools 1–8 entrants, which in turn improved the likelihood of Schools 1–8 graduates to win promotion competitions over close Dapp cohorts. Indeed, Column (3) indicates that Capp had a large, positive, and significant effect on the number of top-ranking officials who graduated from Schools 2–8, where the number of urban-born high-achieving entrants spiked during centralization. The number of top-ranking officials who graduated from Schools 2–8 increased by 47% for Capp cohorts compared to Dapp cohorts. By contrast, in Column (2), Capp only had a small and insignificant effect on the number of top-ranking officials who graduated from School 1, where the number of urban-born high-scoring entrants did not change much during centralization.
Schools 1-8’s impacts on the likelihood of promotion to top-ranking officials appear especially large for high-achieving students. Indeed, Capp’s positive effect on Schools 2–8 graduates was so large (Column (3)) that it dominated Capp’s small and negative effect on those who did not graduate from Schools 1–8 (Column (4)), resulting in a positive effect on the total number of top-ranking officials (Column (1)). This finding suggests that the meritocratic assignment of high achievers to Schools 1-8 produced a net positive effect on the national production of top-ranking officials. This result is inconsistent with the pure selection hypothesis which argues that the role of the centralized admissions is simply to select and send a fixed number of high-achieving students to national higher education and that national higher education itself does not give students any value added.
Schools 1-8’s particularly large value-added for high-achieving students can be due to multiple factors. First, human capital returns to education by these schools could be higher for high-achieving students than lower-achieving students (due to demanding curriculum, for example). Gathering high-achieving students in these schools may also produce higher peer effects. Second, high-achieving students may benefit more from their signaling value or from gaining connections with powerful alumni than low-achieving students. While our data and variation do not allow for a clear decomposition into these different factors, our analysis suggests that the effect of meritocratic centralization remains similar even after controlling for the size of same-cohort colleagues from the same schools or focusing on those with the highest educational qualification (Appendix Section A.4.4). Much of Schools 1-8’s effects therefore likely come from the first human capital factor.
One potential threat to our identification is that the number of available top-ranking positions might increase during the Capp periods. To investigate this concern, recall that we define our cohort by the year of entering a higher school or its equivalent (not by the year of becoming top-ranking officials). Within each cohort, the number of years taken from entering a higher school to the appointment for the first top-ranking position varied widely from 20 to 30 years (see Appendix Figure A.9 and Appendix A.4.3 for details). Individuals in a given cohort were therefore promoted to top-ranking positions in different years. This limits the concern that a potential correlation between the number of top-ranking positions and the lagged periods of centralized admissions may drive our results.
Another potential threat is that the number of top-ranking officials may be differentially impacted by wars between Dapp and Capp cohorts. In particular, Japan was occupied by the US after WWII. During the postwar US Occupation, a substantial number of top officials were purged from public service, and a few were sentenced to death for war crimes. However, we show that the number of top-ranking officials who died in wars, or were purged or executed after WWII, changes little between Dapp and Capp cohorts (Appendix Table A.16).
For career bureaucrats, the meritocratic centralization turns out to be more productively efficient than the decentralized system, as the government envisioned. Whether or not this was true for all career elites is an open question. For example, it is possible that the quality of career bureaucrats improved at the expense of other types of elites. To examine this possibility, we use the 1939 PIR data to compute the share of top bureaucrats among all “socially distinguished individuals” listed in the PIR. [^30] Top bureaucrats are defined as high-ranking central government officials listed in the PIR as in Section As Appendix Figure A.10 shows, we observe no positive association between the share of government officials and the cohort’s exposure to the centralized admissions (although the share of government officials is rising over time reflecting a growing public sector). Our result is unlikely to be driven by students’ brain drain from other elite occupations into top bureaucrats under the meritocratic centralization.
Finally, we obtain suggestive evidence that the meritocratic centralization might produce a greater number of elites in general. As Appendix Figure A.11 and Table A.17 show, higher exposure to the centralized system is associated with a greater total number of individuals listed in the PIR. A potential concern with this analysis is that this may be driven by the age effect. For example, people of certain ages may have had a higher chance of holding important positions and being selected into the PIR, while such ages may happen to be correlated with exposure to centralization. To distinguish the cohort effect from the age effect, we use both 1934 and 1939 editions of the PIR while controlling for the age of each cohort. The resulting estimates show that the number of individuals listed in the PIR is 4–5% higher among Capp cohorts. The centralized system may therefore be more productively efficient not only for producing top bureaucrats but also for producing other types of career elites.
5 Conclusion
The design of school admissions persistently impacts the production and geography of career elites. We reveal this fact by looking at the world’s first recorded use of nationally centralized admissions and its subsequent abolitions in early twentieth-century. While centralization was designed to make the school seat allocations more meritocratic, there turns out to be a tradeoff between meritocracy and equal regional access to elite higher education and career advancement. The meritocratic centralization led students to apply to more selective schools and make more inter-regional applications. As high-achieving students were located disproportionately in urban areas, however, centralization caused urban applicants to crowd out rural applicants from advancing to higher education.
Most importantly, these impacts persist: Several decades later, the centralized system produced more top-ranking government officials, indicating that the meritocratic centralization may be more productively efficient. The distributional effect also persists: The meritocratic centralization increased the number of high income earners, medal recipients, and other elite professionals born in urban areas relative to those born in rural areas.
Even though our study uses the admission reforms unique to our historical setting, the implications of our study may have broader relevance. For instance, distributional consequences of centralized meritocratic admissions may be a reason why many countries continue to use seemingly inefficient decentralized college admissions. While this paper focuses on the decentralized system with a single application, the takeaway from our results would be relevant for other types of decentralized admissions as long as costly applications make applicants self-select into a limited number of schools. For any such decentralized systems, the meritocracy-equity tradeoff would be an important potential concern.
Methodologically, the use of natural experiments in history may be also valuable for studying the long-run effects of market designs in other areas, such as housing, labor, and health markets. It is the multiple bidirectional policy changes in history that allow us to measure the long-run effects. The disadvantage of using historical events is the limited availability of data. The ideal way to alleviate the data concerns would be to use modern administrative data. For example, one may imagine linking administrative tax return data and school district data to measure the long-run effects of school choice reforms around the world in the past few decades. Such an effort would be a fruitful complement to our historical study.
A. Baseline Specification
Age 17 under centralization 3.18 0.61 1.68 2.82 1.84 0.84 0.41
  $\times$ Tokyo area ($<$100 km) (0.027)** (0.016)** (0.060)* (0.010)*** (0.088)* (0.002)*** (0.025)**
[0.000]*** [0.020]** [0.008]*** [0.001]*** [0.045]** [0.016]** [0.182]
B. With Control Variables
Age 17 under centralization 1.93 0.52 1.59 2.44 1.57 0.63 0.39
  $\times$ Tokyo area ($<$100 km) (0.002)*** (0.025)** (0.026)** (0.002)*** (0.071)* (0.001)*** (0.041)**
[0.014]** [0.017]** [0.007]*** [0.000]*** [0.040]** [0.041]** [0.143]
C. Bidirectional Specification with Control Variables
[…]
  $\times$ Tokyo area ($<$100 km) (0.165) (0.078)* (0.006)*** (0.004)*** (0.057)* (0.010)*** (0.149)
[0.034]** [0.049]** [0.006]*** [0.003]*** [0.023]** [0.170] [0.359]
[…]
  $\times$ Tokyo area ($<$100 km) (0.001)*** (0.022)** (0.183) (0.004)*** (0.338) (0.006)*** (0.025)**
[0.005]*** [0.131] [0.021]** [0.008]*** [0.488] [0.032]** [0.079]*
D. Centralization Exposure with Control Variables
Cohort’s exposure to centralization 2.03 0.64 1.73 2.51 1.73 0.60 0.43
  $\times$ Tokyo area ($<$100 km) (0.001)*** (0.008)*** (0.020)** (0.002)*** (0.024)** (0.006)*** (0.074)*
[0.005]*** [0.001]*** [0.007]*** [0.004]*** [0.055]* [0.017]** [0.188]
Observations 705 705 705 705 705 705 705
Birth cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes
Mean dep var 8.77 1.24 5.19 6.21 7.99 1.34 0.82
Mean dep var (Tokyo area 10.62 1.51 6.75 6.86 11.21 1.25 1.02
   under decentralization)
[…]
Table 3: Long-run Impacts of Centralization: Destinations of Career Elites
(1) (2) (3) (4) (5) (6) (7)
Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial
Univ. income income recipients executives politicians & Univ.
grads earners earners bureaucrats professors
A. Baseline Specification
Age 17 under centralization 4.20 0.48 2.03 3.11 3.47 0.61 0.71
  $\times$ Tokyo area ($<$100 km) (0.060)* (0.151) (0.320) (0.056)* (0.238) (0.000)*** (0.120)
[0.000]*** [0.154] [0.060]* [0.005]*** [0.003]*** [0.077]* [0.092]*
B. Adding Control Variables
Age 17 under centralization 3.16 0.74 2.44 2.90 3.64 0.41 0.54
  $\times$ Tokyo area ($<$100 km) (0.021)** (0.221) (0.263) (0.038)** (0.223) (0.017)** (0.043)**
[0.008]*** [0.010]*** [0.012]** [0.010]*** [0.002]*** [0.122] [0.134]
Observations 705 705 705 705 705 705 705
Birth cohort FE, Pref. FE Yes Yes Yes Yes Yes Yes Yes
Mean dep var 8.30 1.21 5.08 5.88 7.74 1.27 0.78
Mean dep var (Tokyo area 14.73 2.46 9.60 9.64 15.54 2.14 1.32
   under decentralization)
Notes: This table shows difference-in-differences estimates of the long-run effects of the centralized admission system on the geographical destinations of career elites. By birth cohort (born in 1880–1894), we count the number of elites who reside in each prefecture in 1939. Unlike the previous tables, all outcome variables are measured at the prefecture of residence. In Panel B, we control for time- and cohort-varying prefecture characteristics, i.e., the number of primary schools in the prefecture in the year when the cohort turned age 17, the number of middle-school graduates in the prefecture in the year when the cohort turned age 17, log GDP of the prefecture when the cohort turned age 20, and birth population of the cohort in the prefecture. Parentheses contain p-values based on standard errors clustered at the prefecture level. Square brackets contain wild cluster bootstrap p-values based on standard errors clustered at the cohort level. ***, **, and * mean significance at the 1%, 5%, and 10% levels, respectively. See Section 4.1 for discussions about this table.
Table 4: Long-run Impacts of Centralization: National Production of Top Government Officials
(1) (2) (3) (4)
Exam passers Exam passers Exam passers Exam passers
graduated graduated not
from from graduated
School 1 Schools 2–8 from
(Tokyo) Schools 1–8
Centralized 5.96 -2.81 14.93 -12.11
(0.802) (0.221) (0.014)** (0.044)**
Observations 33 33 33 33
Control time trend Yes Yes Yes Yes
Control exam passers No Yes Yes Yes
Mean dep var 189.55 37.29 68.47 83.80
Mean dep var (decentralization) 194.22 37.96 64.09 92.17
(1) (2) (3) (4)
Top-ranking Top-ranking Top-ranking Top-ranking
officials officials officials officials
graduated graduated not
from from graduated
School 1 Schools 2–8 from
(Tokyo) Schools 1–8
Centralized 4.19 -0.48 5.22 -0.55
(0.006)*** (0.648) (0.000)*** (0.596)
Observations 33 33 33 33
Control time trend & exam passers Yes Yes Yes Yes
Mean dep var 29.77 8.27 12.97 8.53
Mean dep var (decentralization) 28.66 8.30 11.22 9.14
Notes: Panel (a) shows OLS estimates of the effects of the centralized admissions on the number of individuals who passed the administrative division of the Higher Civil Service Exams (administrative HCSE). Panel (b) shows OLS estimates of the effects of the centralized admissions on the number of top-ranking higher civil officials. The estimates are based on the cohort level data (1898–1930), where cohort is defined by the year of entering a higher school or its equivalent. The data is compiled from the complete list of individuals who passed the administrative HCSE in 1894–1941 and their biographical information. “Exam passers” is the number of individuals in cohort $t$ who passed the administrative HCSE. “Top-ranking officials” is the number of top-ranking officials in cohort $t$ (i.e., the number of individuals who entered a higher school or its equivalent in year $t$, passed the administrative HCSE, and were internally promoted to the top three ranks of higher civil service in their lifetime). “Centralized” is the indicator variable that takes 1 if cohort $t$ entered a higher school or its equivalent under the centralized admissions in 1902–07, 1917–18, and 1926–27. “Mean dep var (decentralization)” is the mean of the dependent variable for the cohorts who entered a higher school or its equivalent under the decentralized admissions. In all regressions, we control for quadratic time trends. Parentheses contain P values based on Newey-West standard errors with the maximum lag order of 3. ***, **, and * mean significance at the 1%, 5%, and 10% levels, respectively. See Section 4.2 for discussions about this table.
Appendix A Online Empirical Appendix
A.1 The Evolution of the Admission System
Table A.1 shows changes in the number and the admission system of National Higher Schools 1-8 (Kanritsu Koutou Gakkou in Japanese) from 1900 to 1930.313131National Higher Schools were initially called the University Preparatory Division of National Higher School. See Moriguchi (2021) for historical details and the sources of information in Table A.1. Despite the population increase and the growing demand for higher education, the number of Schools 1-8 increased only slightly from 6 to 8 in 1900–1918 due to tight fiscal constraints. With the economic boom of WWI, the government expanded the higher education system and increased the number of Schools from 8 in 1918 to 25 in 1925.323232In the late 1920s, in addition to 25 national higher schools, 3 local public higher schools, 4 private higher schools, and one colonial higher school were established. Such expansions notwithstanding, Schools 1–8 remained the most distinguished among all higher schools throughout the pre-WWII period.
A.2 Relation to Historical Literature
[…]
A.3 Data for the Short-run Analysis
To analyze short-run effects of the centralized admission system (Capp), we collect data on application and enrollment outcomes for Schools 1–8 by newly digitizing several administrative and non-administrative sources.
A.3.1 Data on Application Behavior
First, we collect data on the number of applicants by school of their first choice (hereafter first-choice school) for 1900–1930 from the following sources: Ministry of Education Yearbook (Monbushou Nenpou in Japanese) for 1900, 1901, 1908–1916, 1919–1925, and 1928–1930; [^33] Digital images are available online at the National Diet Library Digital Collections. Correspondences from the Ministry of Education to the Tokyo Imperial University for 1903 and 1904; [^34] Digital images are available online at the University of Tokyo Digital Archives. the entrance exam preparation magazine called Middle School World (Chugaku Sekai in Japanese), vol.10, no.12, for 1907; and Higher School Entrance Examination Investigation Report (Koutou Gakkou Nyugaku Shiken ni kansuru Shochousa in Japanese) by the Ministry of Education for 1917, 1918 and 1927; [^35] Digital images are available online at the National Diet Library Digital Collections. no data is available for 1902, 1905, 1906, and 1926. In addition, we collect data on the number of entrants by school for 1900–1930 from Ministry of Education Yearbook.
First-choice school is defined as the school to which an applicant apply under the decentralized admission system and the school which an applicant rank as his first choice under the centralized admission system. There is one issue in this data that requires data cleaning. As Table A.1 indicates, as exceptions under the decentralized system, Schools 7 and 8 are allowed to conduct their entrance exams on different dates from other schools in 1901 (School 7), 1908 (Schools 7 and 8), and 1909–1910 (School 7). In these years, applicants can apply for School 7 or 8 in addition to one of the other schools. For this reason, the number of first-choice applicants to Schools 7 and 8 in these years are double-counted. To correct for double-counting, we estimate the number of first-choice applicants for Schools 7 and 8 in these years using 1) the number of entrants to Schools 7 and 8 in these years and 2) the average ratio of applicants to entrants in 1911–1916 for School 7 and 1909–1916 for School 8, respectively. Using this data, we compute the share of applicants who choose School 1 as their first choice (in Figure 1(a)) and the ratio of first-choice applicants to entrants for each school (in Appendix Figure A.4).
Taking advantage of more detailed data in a supplementary volume of Higher School Entrance Examination Investigation Report of 1917, we collect data on the number of applicants by their first-choice school and by their middle-school prefecture for 1916 and 1917. Middle-school prefecture is defined as the prefecture in which an applicant’s middle school is located. To measure the geographical mobility of applicants, we define application distance (i.e., the distance between an applicant’s middle school and his first-choice school) by the direct (straight-line) distance between the capital of the prefecture in which his middle school is located and the capital of the prefecture in which his first-choice school is located. The distance data is provided by the Geospatial Information Authority of Japan. These data is used to examine the impact of centralized admissions on (a) the share of applicants whose first choice is School 1 and (b) the application distance by regions (Appendix Table A.3).
A.3.2 Data on Enrollment Outcomes
To analyze enrollment outcomes, we use Higher School Student Registers (Gakkou Ichiran in Japanese) published annually by Schools 1–8.363636Digital images are available online at the National Diet Library Digital Collections. For each school, as a proxy for the number of entrants, we collect data on the number of first-year students in the university preparatory course (daigaku yoka in Japanese). Our data starts in 1896 for Schools 1, 2, 4, and 5, 1897 for School 3, 1900 for School 6, 1901 for School 7, and 1908 for School 8, reflecting the year of establishment for each school (School 3 was established in 1896, but the university preparatory course started in 1897). [^37] The data is missing in the following years: 1929 for School 3, 1904 and 1907 for School 5, 1920 for School 6, and 1909 for School 7. Birth prefecture is the prefecture of a student’s legal domicile (honsekichi in Japanese) recorded in the official family registry system (koseki in Japanese). We include students born in any prefectures, but exclude foreign-born students and students born in colonies.
To measure the geographical mobility of entrants in Figure 1(b), we define enrollment distance (i.e., the distance between an entrant’s birth prefecture and the school he entered) by the direct distance between the capital of the birth prefecture and the capital of the prefecture in which the school he entered is located. The distance data is provided by the Geospatial Information Authority of Japan. To produce Figure 1(b), we focus on the sample of schools that use single-school application under the decentralized system. As Schools 7 and 8 are allowed to conduct their entrance exams on different dates from other schools in 1901 (School 7), 1908 (Schools 7 and 8), and 1909–1910 (School 7), applicants can apply for these schools in addition to one of the rest of the schools in these years even under the decentralized system. We exclude these schools in these years from the sample, since the feature of multiple-school application contaminates the essential aspect of the decentralized system.
To analyze the impact of centralized admissions on the geographical composition of entrants (Figure 2(a), Figure 2(b), and Table 1), we control for the size of potential applicants and the number of competing schools in each prefecture, using the following variables. From Ministry of Education Yearbook, we collect data on the number of middle-school graduates (including both public and private schools) by prefecture (defined by school location)  [^38] The data on middle-school graduates is missing for 1920 and is linearly interpolated. as well as the number of higher schools (including national, public, and private schools) by prefecture (defined by school location) for 1900–1930. The total number of higher schools increased from 8 (consisting of 8 national schools) in 1918 to 32 (consisting of 25 national, 3 public, and 4 private schools) in 1930. The Tokyo area is defined by seven prefectures located within 100 km from Tokyo (i.e., Chiba, Gumma, Ibaraki, Kanagawa, Saitama, Tochigi, and Tokyo prefectures) measured by the distance between two prefectural capitals (Appendix Table A.3).
[…]
Finally, to test the exogeneity of the timing of admission reforms, in addition to the data collected above, we collect data on the mean age of entrants for 1905–1930 from Ministry of Education Yearbook. [^39] No data on the mean age of entrants is available for 1900–1904. We also collect data on government expenditures on national higher education (the sum of ordinary and extraordinary expenditures spent on National Higher Schools and Imperial Universities) for 1900–1930 from Ministry of Education Yearbook. Our analysis is presented in Appendix Table A.7.
A.4 Data for the Long-run Analysis
To assess long-run effects of the centralized admission system (Capp), we provide two sets of empirical analyses. In the first analysis, we compile prefecture-cohort level data of career elites and compare career outcomes of urban- and rural-born individuals across birth cohorts that differed in their exposure to Capp. In the second analysis, we construct cohort level data of central government officials using the complete list of individuals who have passed the Higher Civil Service Examinations and compare promotion outcomes of these individuals across cohorts that differed in their exposure to Capp.
A.4.1 Data on the Regional Distribution of Career Elites
Personnel Inquiry Records (PIR) data
To measure students’ career outcomes, the first long-run analysis uses Personnel Inquiry Records (Jinji Koushin-roku in Japanese), which compiles a highly selective list of “socially distinguished individuals” (shakai-teki meishi in Japanese) and provides their biographical information. We use the 1939 edition of the PIR in our main analysis and the 1934 edition for a supplementary analysis. [^40] Digital images of the PIR are available online at the National Diet Library Digital Collections. We thank Hidehiko Ichimura and Yasuyuki Sawada for sharing their digitized PIR data.
In selecting the socially distinguished individuals, the PIR refers to multiple sources such as the government personnel directory (Shokuin-roku in Japanese), the directories of banks and companies (Teikoku Ginkou Kaisha Youroku and Zenkoku Ginkou Kaisha Youroku in Japanese), the directory of the national chamber of commerce and industry members (Zenkoku Shoukou Kaigisho Giin Meibo in Japanese), and the directory of Japanese notables (Nihon Shinshi-roku in Japanese) (see PIR 1934, p.2, for details). [^41] Digital images of these directories are available online at the National Diet Library Digital Collections.
The government personnel directory provides a complete list of public servants (with their job titles and ranks) in national and local governments, including the national diet members, civil and military officials, and Imperial University professors. One of the directories of banks and companies provides a complete list of directors of banks and companies whose capital is 300,000 JPY or above (Teikoku Ginkou Kaisha Youroku 1938, p.2). The directory of Japanese notables includes “wealthy persons” (shisanka in Japanese, defined by individuals who paid income tax of 80 JPY or greater or business tax of 70 JPY or greater) living in urban areas in 24 prefectures (out of 47 prefectures) in Japan (Nihon Shinshi-roku, 1938, p.i).
We use the following information from the PIR for each individual: full name, birth year, birth prefecture, the prefecture of residence, final education, occupation titles and organization names, corporation type (if the organization is incorporated), the medal for merit and the court rank awarded (if any), and the amounts of national income tax and business tax paid in the previous year (if any). To check for accuracy, the PIR verifies the information of birth date and birth prefecture for each individual by obtaining a transcript of the official family register (PIR 1934, p.2).
Defining the Groups of Career Elites
As described in the main text, we define (mutually non-exclusive) groups of elites among the individuals listed in the PIR. We also define landlords as individuals whose occupations include landlord, but excluding the top 0.05% income earners, prestigious medal recipients, corporate executives, top politicians and bureaucrats, and Imperial University professors. To determine individual’s occupations, we check if one’s occupational titles and organization names contain specific Japanese characters that signify a given occupation. The precise definitions of corporate executives, top politicians and bureaucrats, Imperial University professors, landlords, scholars, engineers, physicians, and lawyers are provided in Appendix Table A.8.
[…]
Control Variables
In the regression analysis presented in Table 2 Panels B, C, and D, we control for time- and cohort-varying prefecture characteristics for the cohorts born in 1880–1894. For the robustness check to test the parallel trend assumption (Appendix Table A.12), we also use the cohorts born in 1874–1883.
[…]
Third, to control for local demographic changes, we estimate the prefecture-level male birth population of the cohorts born in 1874–1894 using the following data sources: (a) Japanese Population Census (Nihon Zenkoku Kokou-hyo in Japanese) in 1880–1892 that provides the number of male births by prefecture in each year and (b) Japanese Imperial Population Census (Nihon Teikoku Minseki Kokou-hyou in Japanese) in 1886 that provides the age-specific population by prefecture (from which we use the population of males who were born in 1874–1885 and were 1–12 years old in 1886). [^43] Digital images of population censuses are available online at the National Diet Library Digital Collections. For the cohorts born in 1874–1879, we estimate their birth population combining (a) and (b) as follows. For the cohorts born in 1880–1886, we define the survival rate of cohort $c$ up to 1886 in prefecture $j$ by the number of age-specific population of cohort $c$ in 1886 from (b) divided by the number of births of cohort $c$ from (a). We then estimate the following equation using OLS:
$Ln(Survival_{jc})=\phi_{j}Age_{c}+\theta Age_{c}^{2}+\epsilon_{jc},$ (2)
where $Age_{c}$ is the age of cohort $c$ in 1886. Using the estimated coefficients, we predict the log of survival rates up to 7–11 years old in each prefecture. The birth population of cohorts born in 1875–1879 is obtained by
$Ln(Birth\_Population_{jc})=Ln(Population_{jc}^{1886})-\widehat{Ln(Survival_{jc})}.$ (3)
where $Population_{jc}^{1886}$ is the population of cohort $c$ in 1886. There are several prefectures whose boundaries changed between 1874 and 1886, however. For these prefectures, we use birth population data in 1887–1892 to estimate the linear time trend of male birth population for each prefecture and impute the data prior to 1886. For the cohorts born in 1893 and 1894, we do not have data for their birth population by prefecture. Therefore, we impute them by the average birth population in the prefecture in 1891 and 1892.
Estimating the Intensity of Exposure to Capp
Before 1918, the qualifications for the higher school entrance exam were male middle-school graduates who were at least 17 years old. Accordingly, applicants typically took the exam for the first time at age 17. Given this, in the regression analyses in Table 2 Panels A–C, Table A.14, and Table 3, we simply assume that, among cohorts who were born in 1880–1894 and turned age 17 in 1897–1911, only the cohorts who turned age 17 during the first period of centralized admissions (Capp) in 1902–1907 were fully exposed to Capp, while the rest of the cohorts were fully exposed to Dapp.
Precisely speaking, however, a nontrivial number of unsuccessful applicants retook the exam multiple times in the subsequent years at age greater than 17. For this reason, we also estimate each cohort’s intensity of exposure to Capp that takes into account the presence of exam re-takers as follows. We use a special table in Government Gazette No.5838 (published in December 17, 1902) that reports the number of applicants and the number of admitted applicants in 1902 by the year of middle-school graduation. According to the table, 63.9% of the applicants in 1902 graduated from middle schools in 1902 (the same year), 27.3% in 1901 (the previous year), 7.2% in 1900 (two years before), 1.3% in 1899 (three years before), 0.2% in 1898 (four years before), and 0.1% in 1897 (five years before).
We use the 1902 data to compute the total number of exams taken and the number of exams taken under Capp for each cohort, where cohort is defined by the year of middle-school graduation. This exercise requires us to estimate two sets of parameters. The first set of parameters is the probability of passing the exam in year $t$ for applicants in cohort $c$ who took the exam $n$ times before year $t$, denoted by $a_{n}^{t}$ (where $n=0,...,5$ and $t=1896,...,1911$). Note that the subscript for cohort $c$ is suppressed because $c=t-n$. The second set of parameters is the probability of applicants (who took the exam $n$ times) retaking the exam in year $t+1$ conditional on failing the exam in year $t$, denoted by $r_{n}^{t}$. Because we only have the data for a single year, we assume that these parameters are constant across years (and therefore expressed as $a_{n}$ and $r_{n}$) and that the cohort size is constant across cohorts. Under these assumptions, the probability of passing the exam is estimated to be $a_{0}=0.352$, $a_{1}=0.377$, $a_{2}=0.313$, $a_{3}=0.286$, $a_{4}=0.429$, $a_{5}=0.200$. The probability of retaking the exam is estimated to be $r_{0}=0.660$, $r_{1}=0.423$, $r_{2}=0.255$, $r_{3}=0.175$, $r_{4}=0.011$, $r_{5}=0.000$.
Using these parameters, we simulate the total number of exams taken by each cohort and compute the share of the exams taken under Capp, which we define as the intensity of exposure to Capp. Here we assume that all cohorts graduated from middle-schools at age 17 (so that birth cohorts and cohorts defined by the year of middle-school graduation coincide). The estimated intensity of exposure to Capp is 0.00, 0.02, 0.09, 0.36, 1.00, 0.98, 0.91, 0.64, and 0.00 for the cohort who turned age 17 in 1896–98, 1899, 1900, 1901, 1902–04, 1905, 1906, 1907, and 1908–11, respectively. We color cohorts according to their intensity of exposure to Capp in Figure 3 and Appendix Figures A.7, A.10, and A.11. We also use the estimated intensity of exposure in the regression analyses in Table 2 Panel D and Appendix Table A.17.
A.4.2 Data on the National Production of Career Elites
In the second set of long-run analysis, we focus on a specific group of career elites, i.e., higher civil officials for whom we have complete count data.
Higher Civil Service Examinations (HCSE) data
Our main data source is Hata (1981), which provides not only the list of all individuals who passed the Higher Civil Service Examinations (HCSE, Bunkan Koutou Shiken in Japanese), but also their biographical information available as of 1981 (Hata, 1981, Section 3). The HCSE were highly selective national qualification exams held annually from 1894 to 1947. The administrative division of the HCSE consisted of two parts: the preliminary exams and the main exams. Until 1922, graduates of the law departments of imperial universities were exempted from the preliminary exams, but had to take the main exams (Hata, 1981, pp.663-666; Spaulding, 1967, Chapter 12).
We newly digitized the information of all individuals who passed the administrative division of the HCSE in 1894–1941, including their full name, education, year of university graduation, year of passing the HCSE, starting position, final position, year of retirement, and other notable positions held. Unlike the PIR data, we observe both university and higher school (or its equivalent) in the HCSE data. However, birth year and birth prefecture are missing in the HCSE data.
In the bureaucracy system in Japan, the higher civil service refers to the top ten ranks of national government offices in the administrative, judicial, and diplomatic divisions. Within the higher civil service, the top three ranks were distinctively called “imperial appointees” (chokunin-kan in Japanese) during the prewar period. The first rank consisted of minister level positions, and the second and third ranks consisted of vice minister level positions such as vice minister, director general, bureau chief, and prefectural governor (fuku-daijin, jikan, kyokuchou, and chiji in Japanese). The correspondence between civil service positions and their ranks was reported in the section of salary tables in the government personnel directory (Shokuin-roku in Japanese).
We define “top-ranking officials” as higher civil officials who were internally promoted to reach one of the top three ranks by the end of their career. More precisely, we define “top-ranking officials” as higher civil officials (a) whose final position was in the top three ranks excluding postwar governorship and (b) whose final and notable positions do not include positions in the first rank (i.e., minister-level positions). We exclude postwar governors from top-ranking officials, because starting in 1947 governors were no longer internally promoted but selected by direct election. We further exclude higher civil officials who were appointed to any minister-level positions, because these positions were filled by political appointments and not by internal promotion. See Spaulding (1967) and Shimizu (2019) for the establishment of the internal promotion system in prewar Japan.
Defining Cohorts (by the year of entering a higher school or its equivalent)
To identify each individual’s exposure to the centralized admission system, we must find out in which year each individual took the entrance exam and entered a higher school (or failed and entered an alternative school). However, since we only observe the year of university graduation in the HCSE data, we estimate “the year of entering a higher school or its equivalent” separately for (1) top-ranking officials who graduated from one of Schools 1–8 (and an imperial university), (2) officials who graduated from one of Schools 1–8 (and an imperial university) but are not top-ranking officials, and (3) the rest of officials (who did not graduate from Schools 1–8), as follows.
First, for top-ranking officials who graduated from Schools 1–8, we find the exact year by searching each individual’s full name in the list of first-year students in the Student Registers of Schools 1–8 in all editions. The exact match was found for 699 out of 733 individuals with a matching rate of 95.1%. For these individuals, in Appendix Figure A.8, we plot the number of years taken from entering Schools 1–8 to university graduation by cohort (defined by the year of university graduation) using a round marker. Because higher school and imperial university were both three-year programs, in principle one could complete both programs in 6 years. However, it was quite common for students to repeat the same year in both higher school and imperial university, especially in the earlier period when there was a strict system of holding back students who failed year-end exams. As a result, as the figure indicates, it took substantially longer than 6 years on average to complete both programs especially for the earlier cohorts.
Second, for non top-ranking officials who graduated from Schools 1–8, we assume that the distribution of the number of years taken from entering Schools 1–8 to university graduation is the same as that of the top-ranking officials (who graduated from Schools 1–8) of the same cohort (defined by the year of university graduate) computed above. To check the validity of this assumption, for two representative cohorts of 1914 and 1922, we randomly select 25% of the non top-ranking officials (who graduated from Schools 1–8) from each cohort and find the exact year of entering a higher school for each individual using the same method of searching a full name in the Student Registers. In Appendix Figure A.8, we plot the average number of years taken from entering Schools 1–8 to university graduation for each subsample in 1914 and 1922 using a diamond marker. The average number is not significantly different between top-ranking and non top-ranking officials (i.e., both numbers are mutually within the 95% confidence intervals) in both 1914 and 1922.
Third, for the officials who are not Schools 1–8 graduates, we assume that the year of entering a higher school or its equivalent is “the year of university graduation minus 6.” Because the number of National Higher Schools increased from 8 in 1918 to 25 by 1925, this group includes the officials who graduated from a higher school other than Schools 1–8 and an imperial university, as well as the officials who graduated from private higher education institutions (e.g., Waseda or Meiji University and its affiliated high school). In either case, most universities and higher schools or their equivalent were three-year programs. For a robustness check, we alternatively assume that the year of entering a higher school or its equivalent is “the year of university graduation minus 7” and show that the empirical results are qualitatively the same (see Appendix Table A.15).
To proceed to an empirical analysis, we redefine cohort by “the year of entering a higher school or its equivalent” estimated above (so that we can see which cohorts are exposed to Capp as opposed to Dapp). That is, to create cohort-level data, we count the number of exam passers (i.e., those who passed the administrative division of the HCSE) and the number of top-ranking officials by the year in which they entered a higher school or its equivalent. Our dataset consists of 6,255 individuals who entered a higher school or its equivalent in 1898–1930 and passed the HCSE in 1901–1941.
Out of 6,255 exam passers, 3,490 individuals or 55.8% are Schools 1–8 graduates, 4,767 individuals or 76.2% are Imperial University graduates, and 982 individuals or 15.7% are top-ranking officials. The share of Imperial University graduates is greater than that of Schools 1–8 graduates because the number of higher schools increased from 8 to 25 in 1919–1930. Among 982 top-ranking officials, 701 officials or 71.4% are Schools 1–8 graduates, and 891 officials or 90.7% are Imperial University graduates. The HCSE consisted of preliminary and main exams, and because Imperial University law graduates were the only group exempted from the administrative preliminary exams from 1894 to 1922, they had substantial advantages in passing the HCSE (Hata, 1981, pp.663-666; Spaulding, 1967, Chapter 12). This exemption, however, does not explain why Imperial University graduates had such a high share in the top-ranking officials.
A.4.3 Variation in Years Taken to be Promoted to Top-ranking Positions
One potential threat to our identification in Table 4 Panel (b) is a possibility that the number of available top-ranking positions happened to have increased during the periods of Capp. However, we argue that even if this was the case, it is not likely to affect our results, since our cohort is defined by the year of entering a higher school or its equivalent, not by the year of becoming top-ranking officials. Namely, as long as individuals in a given cohort were not promoted to a top-ranking position in the same year, a potential correlation between the number of top-ranking positions and the lagged periods of centralized admissions does not bias our results.
To examine this possibility, we first randomly selected two cohorts exposed to Capp (1903 and 1916) and two cohorts exposed to Dapp (1913 and 1922). For all top-ranking officials who graduated from Schools 1–8 in these cohorts, we searched the years in which these officials were appointed to their first top-ranking positions. Using online searches to find biographical information for each official, we obtained necessary information for 82% of these officials. As shown in Appendix Figure A.9, within each cohort, the number of years taken from entering a higher school to the appointment for the first top-ranking position varied widely from 20 to 30 years. This within-cohort variation is a sum of two variations, the first variation in the number of years taken from entering a higher school to passing the exams and the second variation in the number of years taken from passing the exams to becoming a top-ranking official. Therefore, it is unlikely that the result in Table 4 Panel (b) is driven by a greater number of available top-ranking positions that coincided with the periods of Capp.
A.4.4 Connections and Signaling Hypotheses
We investigate whether connections or signaling was an important mechanism through which the meritocratic centralization increased the total number of top-ranking officials. First, to explore the connection channel, we reexamine the result of Table 4 Panel (b) Column (3) by controlling for the number of exam passers from the same schools (Schools 2–8) in the same cohort, instead of the total number of exam passers. This variable takes into account the size of same-cohort colleagues from the same schools who are likely to have network values. The coefficient of Capp remains sizable (2.72) and statistically significant at the 1% level, implying 24% increase in the number of top-ranking officials from these schools relative to the mean during Dapp (11.22). Second, to investigate the signaling channel, we further restrict our sample to top-ranking officials who graduated from both Schools 2–8 and Tokyo Imperial University, the most prestigious institution with the strongest signaling value. We regress the number of this subgroup of top-ranking officials on Capp, controlling for the number of exam passers from the same schools in the same cohort. If signaling is a major channel, we expect the effect of Capp among this subgroup to be larger. The coefficient of Capp decreases to 1.76, implying 18% increase in the number of top-ranking officials from these schools relative to the mean during Dapp (10.00). This effect size is smaller than the effect size of 24% above.
Figure A.1: College Admissions around the World Today
Notes: This figure summarizes each country and territory’s college admission system today. Dark red color (e.g. Norway): Regionally- or nationally-centralized college admissions where a single-application, single-offer assignment algorithm (well-defined rule) is used to make admissions to both public and private universities. Medium orange color (e.g. Brazil): Semi-centralized college admissions defined as either (1) there is a centralized system, but not all universities (such as private universities) are included in the single-application, single-offer system or (2) students submit a single application and receive multiple offers. Light orange color (e.g. US): Decentralized college admissions where each college defines its own admission standards and rules. Yellow with diagonal lines (e.g. Chad): Not enough information available or if the country or territory does not have tertiary institutions. We summarize the information sources at https://www.scribd.com/document/437545135/Online-Appendix191018. See Section 2 for discussions about this figure.
Figure A.2: Map of Schools 1–8 and Definition of the Tokyo Area
Notes: This figure shows the locations of Schools 1–8 and the location of the Tokyo area (in dark gray color) defined as a set of 7 prefectures that are within 100 km from Tokyo (i.e., Tokyo, Chiba, Kanagawa, Saitama, Ibaraki, Tochigi, and Gunma prefectures). See Sections 2 and 3.4 for discussions about this figure.
Table A.1: The Evolution of the Admission System
Notes: This table shows changes in the number and the admission system of Schools 1-8 from 1900 to 1930. In 1901, all schools held their exams on the same date, with the exception of newly established School 7. In 1908, all schools held their exams on the same date, with the exception of Schools 7 and 8. See Section 2.1 and Appendix Section A.1 for discussions about this table.
Figure A.3: Centralized Assignment Rule
Notes: This figure is a reprint of the assignment algorithm of the centralized admission system stated in the Ordinance of the Ministry of Education No.4 published in Government Gazette No.1419, pp.580-581, on April 27, 1917. See Sections 2 and 3.3 for an English translation and discussions.
Table A.2: Summary Statistics
Variable Mean Std. Dev. Median N
Year level data on short-run outcomes, 1900–1930
No. of applicants to Schools 1–8 10777 4122 10187 27
Share of applicants choosing School 1 as their first choice 0.29 0.11 0.24 27
No. of entrants to Schools 1–8 2007 333 2147 31
Applicant level data on short-run outcomes, 1916–1917
Distance between middle-school prefecture and the first-choice school (km) 224.88 272.03 117 20913
Applying to School 1 as first choice 0.33 0.47 0 20913
Entrant level data on short-run outcomes, 1900–1930
Distance between birth prefecture and the school entered (km) 224 255 139 65251
Entering the nearest school from birth prefecture 0.49 0.5 0 66193
Born in Tokyo prefecture 0.09 0.29 0 66193
Born in the Tokyo area (7 prefectures within 100 km from Tokyo) 0.17 0.38 0 66193
Prefecture-year level data on short-run outcomes, 1900–1930
No. of entrants to Schools 1-8 45.06 37.45 34 1469
No. of entrants to School 1 7.88 14.11 5 1469
No. of entrants to School 2 5.50 10.40 2 1469
No. of entrants to School 3 6.19 10.34 3 1421
No. of entrants to School 4 5.64 9.91 3 1469
No. of entrants to School 5 6.27 14.20 1 1422
No. of entrants to School 6 5.19 11.80 2 1421
No. of entrants to School 7 5.03 12.99 2 1328
No. of entrants to School 8 5.67 12.45 2 1093
No. of public middle-school graduates 415.38 350.01 299 1410
No. of private middle-school graduates 118.49 435.12 0 1410
No. of national higher schools other than Schools 1–8 0.14 0.43 0 1469
Prefecture-cohort level data on long-run outcomes, PIR-listed individuals born in 1880–1894
No. of all Imperial University graduates 8.77 7.81 7 705
No. of individuals in the top 0.01% income group 1.24 2.05 1 705
No. of individuals in the top 0.05% income group 5.19 7.57 3 705
No. of civilians receiving medal of the Order of Fifth Class and above 6.21 4.94 5 705
No. of corporate executives with a positive amount of tax payment 7.99 8.91 5 705
No. of top politicians and high-ranking bureaucrats 1.34 1.38 1 705
No. of Imperial University professors 0.82 1.18 0 705
Cohort level data on long-run outcomes, government officials entering higher school or equivalent in 1898–1930
No. of passers of the Higher Civil Service Exams 189.55 84.27 159.7 33
No. of top-ranking officials (internally promoted to top three ranks) 29.77 8.36 29.1 33
No. of top-ranking official who are School 1 graduates 8.27 3.36 8 33
No. of top-ranking officials who are Schools 2–8 graduates 12.97 5.98 12 33
No. of top-ranking officials who are not Schools 1–8 graduates 8.53 5.73 6.8 33
Table A.3: Centralization Caused Applicants Across the Country to Apply More Aggressively
A. Selecting School 1 as First Choice
Centralized 0.159 0.192 0.151 0.146 0.128 0.168 0.180 0.166 0.114
(0.000)*** (0.000)*** (0.003)*** (0.001)*** (0.142) (0.001)*** (0.001)*** (0.007)*** (0.001)***
Constant 0.248 0.494 0.169 0.0892 0.178 0.107 0.184 0.0813 0.127
(0.001)*** (0.000)*** (0.002)*** (0.003)*** (0.018)** (0.002)*** (0.000)*** (0.015)** (0.088)*
Sample region All S1 Region S2 Region S3 Region S4 Region S5 Region S6 Region S7 Region S8 Region
Observations 20,913 6,505 2,555 3,248 1,266 2,730 2,276 615 1,718
B. Application Distance
Centralized -2.534 -92.88 10.95 2.080 -15.74 128.0 46.52 145.4 -25.57
(0.914) (0.000)*** (0.670) (0.720) (0.541) (0.003)*** (0.012)** (0.021)** (0.264)
Constant 226.2 231.7 289.7 158.8 166.7 252.6 294.1 218.0 154.2
(0.000)*** (0.000)*** (0.008)*** (0.002)*** (0.061)* (0.002)*** (0.001)*** (0.092)* (0.051)*
Sample region All S1 Region S2 Region S3 Region S4 Region S5 Region S6 Region S7 Region S8 Region
Observations 20,913 6,505 2,555 3,248 1,266 2,730 2,276 615 1,718
Notes: In Panel A, we estimate the effects of the centralized admissions on the propensity of an applicant to select the most prestigious and selective school (School 1 in Tokyo) as the first choice, using the applicant-level data in 1916 (under the decentralized system) and 1917 (under the centralized system). The prefecture-level application data is available only for these two years. We estimate the following regression: $Y_{it}=\alpha+\beta\times Centralized_{t}+\epsilon_{it},$ where $Y_{it}$ is the indicator that applicant $i$ in year $t$ selects School 1 as the first choice. $Centralized_{t}$ is the indicator that year $t$ is under the centralized system. To observe regional variation, we estimate the equation separately by region of the applicant’s middle school. More specifically, we group applicants into “school regions” based on which school (among Schools 1–8) is closest to the applicant’s middle school in 1916. The following map shows the locations of the eight school regions. In Panel B, we estimate the effects of centralization on the application distance defined by the distance between an applicant’s first-choice school and middle school. Parentheses p-values based on standard errors clustered at the prefecture level. ***, **, and * mean significance at the 1%, 5%, and 10% levels, respectively. See Section 3.2 for discussions about this table.
Figure A.4: Changes in the Competitiveness of Schools 1–8
Notes: This figure shows the time evolution of the competitiveness of Schools 1–8 (measured by the ratio of the number of applicants who rank the school first to the number of entrants to the school) from 1900 to 1930. No data is available for 1902, 1905, 1906, and 1926. Colored years (1902–07, 1917–18, 1926–27) indicate the periods of the centralized admission system. See Section 3.2 for discussions about this figure.
Table A.4: Centralization Increased Urban-born Entrants to Schools 2–8: Additional Control Variables
(1) (2) (3) (4) (5) (6) (7) (8) (9)
Dependent variable = No. of entrants to: All schools Sch. 1 Sch. 2 Sch. 3 Sch. 4 Sch. 5 Sch. 6 Sch. 7 Sch. 8
Centralized x Born in Tokyo prefecture 40.25 3.96 6.14 6.91 3.87 10.72 4.08 19.41
(0.000)*** (0.000)*** (0.000)*** (0.000)*** (0.000)*** (0.000)*** (0.016)** (0.000)***
[0.028]** [0.160] [0.036]** [0.010]** [0.031]** [0.000]*** [0.264] [0.000]***
Centralized x Born near Tokyo prefecture (1–100 km) 11.20 0.14 0.66 1.85 0.51 1.01 0.59 0.59
(0.000)*** (0.725) (0.072)* (0.000)*** (0.214) (0.040)** (0.240) (0.382)
[0.000]*** [0.769] [0.128] [0.000]*** [0.057]* [0.018]** [0.194] [0.044]**
Centralized x Born in school’s prefecture -17.77 -18.84 -18.47 -14.75 -23.00 -27.66 -21.99 -47.49 -13.33
(0.001)*** (0.000)*** (0.000)*** (0.000)*** (0.000)*** (0.000)*** (0.000)*** (0.000)*** (0.000)***
[0.025]** [0.006]*** [0.021]** [0.065]* [0.005]*** [0.001]*** [0.079]* [0.001]*** [0.386]
Centralized x Born near school’s prefecture (1–100 km) -3.79 -0.01 -2.86 -3.29 -8.96 -11.33 -1.65 -1.55 1.16
(0.116) (0.987) (0.271) (0.107) (0.007)*** (0.001)*** (0.177) (0.000)*** (0.231)
[0.047]** [0.980] [0.043]** [0.000]*** [0.001]*** [0.001]*** [0.062]* [0.510] [0.687]
No. middle-school grads in prefecture 6.02 -2.64 -0.77 0.78 -0.16 2.76 1.87 2.77 1.67
(0.200) (0.136) (0.336) (0.440) (0.800) (0.001)*** (0.434) (0.211) (0.406)
[0.010]*** [0.000]*** [0.021]** [0.077]* [0.645] [0.003]*** [0.003]*** [0.000]*** [0.013]**
No. middle-school grads in prefecture (1–100km) -1.62 -0.18 -0.27 1.72 -0.31 0.09 -0.43 -1.32 -0.17
(0.439) (0.580) (0.579) (0.191) (0.632) (0.898) (0.589) (0.140) (0.819)
[0.003]*** [0.303] [0.032]** [0.000]*** [0.274] [0.695] [0.067]* [0.000]*** [0.274]
No. other schools -24.06 -1.83 -2.74 -3.01 -1.94 -3.33 -2.40 -3.42 -4.87
(0.000)*** (0.031)** (0.011)** (0.054)* (0.005)*** (0.006)*** (0.080)* (0.027)** (0.066)*
[0.000]*** [0.001]*** [0.000]*** [0.000]*** [0.000]*** [0.000]*** [0.000]*** [0.000]*** [0.000]***
Population in prefecture 14.35 8.80 3.55 1.39 -0.33 -0.59 -2.32 -1.20 1.42
(0.043)** (0.066)* (0.014)** (0.066)* (0.729) (0.540) (0.380) (0.204) (0.674)
[0.000]*** [0.000]*** [0.000]*** [0.058]* [0.622] [0.465] [0.001]*** [0.118] [0.184]
GDP per capita in prefecture 18.36 2.84 0.25 4.65 4.60 2.86 5.34 -3.40 0.03
(0.074)* (0.209) (0.878) (0.087)* (0.034)** (0.033)** (0.031)** (0.530) (0.986)
[0.000]*** [0.000]*** [0.573] [0.001]*** [0.000]*** [0.012]** [0.000]*** [0.002]*** [0.978]
Observations 1,457 1,457 1,457 1,410 1,457 1,410 1,410 1,269 1,081
Year FE, Prefecture FE Yes Yes Yes Yes Yes Yes Yes Yes Yes
Mean dep var 45.43 7.95 5.55 6.23 5.69 6.32 5.23 5.02 5.74
Mean dep var (Tokyo pref under decentralization) 201.1 104.7 27.52 10.85 14.48 6.10 9.20 11.72 20.21
Mean dep var (Within 1–100km from Tokyo pref. 26.23 9.119 6.738 1.208 2.865 0.746 1.242 1.546 3.228
under decentralization)
Notes: This table uses the prefecture-year level data in 1900–1930. “Middle-school graduates in prefecture” is the number of students who graduated from middle schools in prefecture $p$ in year $t$. “Middle-school graduates in nearby prefectures” is the number of students who graduated from middle schools in the prefectures within 100 km from prefecture $p$ (excluding prefecture $p$) in year $t$. “Population in prefecture” is the population in prefecture $p$ in year $t$. “GDP per capita in prefecture” is gross value-added per capita in prefecture $p$ in year $t$. See Tables 1 for the definition of other variables. Parentheses contain p-values based on standard errors clustered at the prefecture level. Square brackets contain p-values based on standard errors clustered at the year level. ***, **, and * mean significance at the 1%, 5%, and 10% levels, respectively. See Section 3.4 for discussions about this table.
Table A.5: Characteristics of the Tokyo Area
Mean Std. Dev. Min Max N
Tokyo Area
   Population in prefecture (in million) 1.497 0.866 0.820 5.437 217
   GDP per capita in prefecture (in 1,000 yen) 0.209 0.095 0.108 0.469 217
   Middle-school graduates in prefecture 0.871 1.172 0.011 6.427 217
   Middle-school graduates in nearby prefectures (1-100 km) 4.604 2.150 0.395 12.212 217
   Share of applicants at School 1 under decentralization 0.379 0.090 0.270 0.545 217
Other Areas
   Population in prefecture (in million) 1.072 0.522 0.414 3.568 1240
   GDP per capita in prefecture (in 1,000 yen) 0.173 0.058 0.097 0.494 1240
   Middle-school graduates in prefecture 0.474 0.431 0.029 3.307 1240
   Middle-school graduates in nearby prefectures (1-100 km) 1.299 1.594 0.000 11.729 1240
   Share of applicants at School 1 under decentralization 0.156 0.078 0.044 0.421 1240
Notes: This table shows the characteristics of the Tokyo area (defined as 7 prefectures within 100 km of Tokyo as in the map in Appendix Figure A.2) compared to other areas. All numbers are the prefecture-level average in 1900–1930. GDP per capita is in real terms expressed in 1934–1936 prices.
Table A.6: Replacing the Tokyo Area Indicator by Urban Characteristics
(1) (2) (3) (4)
Entrants to Schools 1-8
Centralized $\times$ Population in prefecture 2.63
[0.183]
Centralized $\times$ GDP per capita in prefecture 3.16
[0.224]
Centralized $\times$ Middle-school graduates in prefecture 3.73
(0.003)***
[0.171]
Centralized $\times$ Middle-school graduates in nearby prefectures (1–100 km) 4.32
(0.001)***
[0.019]**
Centralized $\times$ Share of applicants at School 1 (under Decentralization) 5.48
(0.000)***
[0.002]***
Population in prefecture 11.74
(0.061)*
[0.000]***
GDP per capita in prefecture 17.30
(0.099)*
[0.001]***
Middle-school graduates in prefecture 10.70 13.64 15.65 16.59
(0.004)*** (0.001)*** (0.000)*** (0.000)***
[0.000]*** [0.000]*** [0.000]*** [0.000]***
Middle-school graduates in nearby prefectures (1–100 km) -0.73
[0.322]
Observations 1,457 1,457 1,457 1,457
Year FE, Prefecture FE Yes Yes Yes Yes
Mean dep var 45.43 45.43 45.43 45.43
Notes: This table uses the prefecture-year level data in 1900–1930. The dependent variable is the number of students from birth prefecture $p$ who entered one of Schools 1–8 in year $t$. “Population in prefecture” is population in prefecture $p$ in year $t$. “GDP per capita in prefecture” is real gross value-added per capita in prefecture $p$ in year $t$. “Middle-school graduates in prefecture” is the number of students who graduated from middle schools in prefecture $p$ in year $t$. “Middle-school graduates in nearby prefectures” is the number of students who graduated from middle schools in the prefectures within 100 km from prefecture $p$ (excluding prefecture $p$) in year $t$. “Share of applicants to School 1 (under decentralization)” is the share of applicants to School 1 among all applicants to Schools 1–8 in prefecture $p$ under the decentralized system in 1916 (the only year for which the data is available). All variables interacted with “Centralized” are standardized to be mean 0 and standard deviation 1. We control for year fixed effects, prefecture fixed effects, and the number of higher schools other than Schools 1–8 in prefecture $p$ in year $t$. We also control for “Born in school’s prefecture,” “Born near school’s prefecture (1-100 km),” and “Born near school’s prefecture (100-300 km)” as in Table 1. Parentheses contain p-values based on standard errors clustered at the prefecture level. Square brackets contain p-values based on standard errors clustered at the year level. ***, **, and * mean significance at the 1%, 5%, and 10% levels, respectively. See Section 3.4 for discussions about this table.
Table A.7: Testing Exogeneity of Centralization
Placebo outcomes Main outcomes
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10)
No. middle No. entrants Share of No. applicants Ratio of Mean Government Share of Enrollment Share of
school to Schools 1–8 entrants to Schools 1–8 entrants to age of expenditures applicants distance entrants
graduates to School 1 applicants entrants on higher to School 1 born in
education Tokyo area
Centralized 1.357 -0.101 0.00556 -0.460 0.0224 0.0523 0.760 0.164 57.02 0.0394
(0.418) (0.109) (0.104) (0.486) (0.110) (0.518) (0.676) (0.001)*** (0.000)*** (0.000)***
No. of middle-school graduates 0.00493 -0.000942 -0.150 0.00134 0.0656 -0.429 0.00194 1.345 0.00147
(0.291) (0.022)** (0.046)** (0.165) (0.000)*** (0.062)* (0.468) (0.129) (0.041)**
Observations 31 31 31 27 27 26 31 27 31 31
Mean dep var 23.41 2.007 0.180 10.78 0.217 19.03 13.56 0.291 228.9 0.174
Notes: Columns 1–7 test if important institutional variables are correlated with the timing of centralization using year-level data. Columns 8–10 examine the robustness of our main short-run outcomes using year-level data. All numbers are at the national-level from 1900 to 1930. The numbers of middle school graduates, entrants, and applicants are denominated by 1,000. In all regressions, quadratic time trends (i.e. trend and trend squared, where the trend is defined by “year - 1899”) are controlled. Parentheses contain p-values based on Newey-West standard errors with the maximum lag order of 3. See Section 3.5 for discussions about this table.
Table A.8: Definitions of Career Elites in the PIR
Notes: This table provides the precise definitions of career elites, defined as a subset of individuals listed in PIR (1939). See Section 4.1 for the analysis using these variables.
Figure A.5: Sampling Rates of High Income Earners in the PIR Data
Notes: This figure plots the sampling rate of the high income earners listed in PIR (1939) by the income level expressed as a top percentile of the national income distribution. The sampling rates and the top income percentiles are computed from the complete count data in National Tax Bureau Yearbook. The vertical lines indicate the top 0.05% and 0.01% thresholds used in our analysis. See Section 4.1 for discussions about this figure.
Figure A.6: High Income Earners in the PIR vs Income Tax Statistics across Prefectures
Notes: These figures compare the number of high income earners in each prefecture listed in PIR (1939) and the complete count of high income earners in each prefecture reported in National Tax Bureau Yearbook (1936). The vertical axis is the log of the number of individuals in the PIR who earned more than 50,000 JPY taxable income (corresponding to the top 0.01% income group) or 16,950 JPY taxable income (corresponding to the top 0.05% income group) in 1938. The horizontal axis is the log of the number of individuals in tax statistics who earned more than 30,000 JPY taxable income (corresponding to the top 0.013% income group) or 10,000 JPY taxable income (corresponding to the top 0.08% income group) in 1936 (the closest year to 1938 for which prefecture-level tax statistics are available). See Section 4.1 for discussions about these figures.
Table A.9: Correlations between Prefecture-level Sampling Rates and Outcome Variables
Top 0.01% Top 0.05% Top 0.01% Top 0.05%
income income income income
earners earners earners earners
Entrants to Schools 1–8 -0.000035 0.000056
(0.000059) (0.000043)
Imperial Univ. grads -0.000052 0.000030
(0.000032) (0.000022)
Observations 47 47 47 47
R-squared 0.003527 0.033977 0.006313 0.008153
Mean dep var 0.44 0.24 0.44 0.24
Notes: This table shows the results of regressing the sampling rates of PIR (1939) on our outcome variables using prefecture-level data. “Top 0.01% income earners” is the sampling rate of the top 0.01% income earners defined by the number of individuals with more than 50,000 JPY of taxable income in 1938 divided by the complete count of the number of individuals with more than 30,000 JPY of taxable income in 1936. “Top 0.05% income earners” is the sampling rate of the top 0.05% income earners defined by the number of individuals with more than 16,950 JPY of taxable income in 1938 divided by the complete count of the number of individuals with more than 10,000 JPY of taxable income in 1936. “Entrants to Schools 1–8 ” is the number of entrants to Schools 1–8 during 1900–1911 who were born in the prefecture (mean=590 and SD=383). “Imperial Univ. grads” is the number of individuals residing in the prefecture in 1938 who graduated from one of the Imperial Universities (mean=224 and SD=349). See Section 4.1 for discussions about this table.
Figure A.7: Long-run Impacts of Centralization: Geographical Origins of Other Career Elites
Notes: This figure shows additional difference-in-differences plots that compare the average number of elites born in prefectures inside and outside the Tokyo area by cohort (see Figure 3). The plots are based on the data from PIR (1939), which covers cohorts who were born in 1879–1894 and turned age 17 (main application age) in 1896–1911. The vertical axis shows the number of indicated elites who were born in the indicated area in the indicated cohort. The cohorts are colored according to their intensity of exposure to the centralized admissions in 1902–07, where the darker color indicates the higher intensity. See Section 4.1 for discussions about this figure.
Table A.10: Long-run Impacts: Other Elite Professionals
(1) (2) (3) (4)
Scholars Engineers Physicians Lawyers
Age 17 under centralization 0.87 0.66 0.46 0.18
$\times$ Tokyo area ($<$100 km) (0.026)** (0.048)** (0.038)** (0.420)
[0.094]* [0.125] [0.091]* [0.544]
Observations 705 705 705 705
Birth cohort FE, Birth pref. FE Yes Yes Yes Yes
Mean dep var 3.61 2.46 1.92 1.11
Mean dep var (Tokyo area 4.76 3.13 2.52 1.16
under decentralization)
Notes: This table shows difference-in-differences estimates of the long-run effects of the centralized admission system. The estimates are based on the birth-prefecture-cohort level data from PIR (1939) which covers cohorts who were born in 1880–1894 and turned age 17 (main application age) in 1897–1911. All regressions control for birth-prefecture fixed effects and cohort fixed effects. “Scholars,” “Engineers,” “Physicians,” and “Lawyers” are defined at the prefecture-cohort level as the number of individuals listed in PIR (1939) whose occupations include scholar, engineer, physician, and lawyer, respectively. “Age 17 under centralization” takes 1 if the cohort turned 17 years old during 1902–07 under the centralized admissions, and takes 0 otherwise. “Mean dep var” shows the mean of the dependent variable for all prefecture-cohort observations. “Mean dep var (Tokyo area under decentralization)” shows the mean of the dependent variable in the Tokyo area under the decentralized admissions. Parentheses contain p-values based on standard errors clustered at the prefecture level. Square brackets contain wild cluster bootstrap p-values based on standard errors clustered at the cohort level. See Section 4.1 for discussions about this table.
Table A.11: Long-run Impacts: Excluding Cohorts who Turned Age 17 in 1901 or 1907
Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial
Univ. income income recipients executives politicians & Univ.
grads earners earners bureaucrats professors
A. Baseline Specification
Age 17 under centralization 3.05 0.68 1.65 2.61 1.89 0.68 0.45
  $\times$ Tokyo area ($<$100 km) (0.003)*** (0.007)*** (0.048)** (0.007)*** (0.030)** (0.019)** (0.067)*
[0.000]*** [0.026]** [0.011]** [0.000]*** [0.066]* [0.023]** [0.189]
B. Adding Control Variables
Age 17 under centralization 1.70 0.58 1.53 2.18 1.60 0.45 0.41
   $\times$ Tokyo area ($<$100 km) (0.021)** (0.013)** (0.007)*** (0.000)*** (0.014)** (0.028)** (0.095)*
[0.017]** [0.008]*** [0.020]** [0.006]*** [0.069]* [0.086]* [0.186]
Observations 611 611 611 611 611 611 611
Birth cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes
Mean dep var 8.68 1.23 5.16 6.08 7.96 1.34 0.83
Mean dep var (Tokyo area 10.38 1.41 6.57 6.61 10.91 1.21 1.02
   under decentralization)
Notes: This table repeats the same difference-in-differences analysis in Table 2, but excluding the cohorts who turned age 17 (main application age) in 1901 or 1907 from the sample as these cohorts were exposed to both centralized and decentralized admission systems. All the variables are defined as in Table 2. Parentheses contain p-values based on standard errors clustered at the prefecture level. Square brackets contain wild cluster bootstrap p-values based on standard errors clustered at the cohort level. See Section 4.1 for discussions about this table.
Table A.12: Long-run Impacts: Pre-event Trends Are Parallel
Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial
Univ. income income recipients executives politicians & Univ.
grads earners earners bureaucrats professors
A. Baseline Specification
Tokyo area ($<$ 100 km) 0.19 -0.02 0.07 0.17 0.11 -0.01 0.02
  $\times$ Time trend (0.673) (0.802) (0.822) (0.476) (0.760) (0.879) (0.835)
B. Adding Control Variables
Tokyo area ($<$ 100 km) 0.21 -0.01 0.10 0.18 0.14 -0.01 0.02
  $\times$ Time trend (0.630) (0.836) (0.751) (0.448) (0.701) (0.885) (0.841)
C. Squared trend term
Tokyo area ($<$ 100km) $\times$ Time trend -0.93 -0.08 -0.48 0.02 0.09 -0.13 0.22
(0.415) (0.817) (0.275) (0.981) (0.903) (0.629) (0.412)
[…]
(0.465) (0.824) (0.288) (0.878) (0.927) (0.695) (0.333)
Observations 470 470 470 470 470 470 470
Birth cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes
Mean dep var 5.02 1.21 4.50 4.34 6.34 0.66 0.39
Mean dep var (Tokyo area 6.47 1.29 5.96 5.23 8.80 0.61 0.54
   under decentralization)
Notes: This table tests if there are differences in pre-event trends between urban and rural areas in the difference-in-differences analysis in Table 2. For cohorts who were born in 1874–1883 and thus turned age 17 in 1891–1900, we run the following regression in Panels A and B:
where $Trend_{t}$ is defined as the cohort’s birth year minus 1870 (the linear time trend). In Panel C, for the same cohorts, we run the following regression:
In Panels B and C, we control for the number of primary schools in the prefecture in the year when the cohort turned eligible age and the birth population of the cohort in the prefecture. All the other variables are defined in the same way as in Table 2. Parentheses contain p-values based on standard errors clustered at the prefecture level. See Section 4.1 for discussions about this table.
Table A.13: Long-run Impacts: Difference-in-Differences Estimates Using the PIR in 1934
(1) (2) (3) (4) (5) (6) (7)
Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial
Univ. income income recipients executives politicians & Univ.
grads earners earners bureaucrats professors
Age 17 under centralization 1.71 0.50 1.81 1.44 1.11 0.39 0.08
  $\times$ Tokyo area ($<$100 km) (0.016)** (0.044)** (0.029)** (0.070)* (0.013)** (0.018)** (0.717)
[0.004]*** [0.003]*** [0.004]*** [0.008]*** [0.008]*** [0.282] [0.496]
Observations 705 705 705 705 705 705 705
Birth cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes
Mean dep var 4.79 0.81 3.64 3.86 2.80 1.18 0.70
Mean dep var (Tokyo area 5.68 1.11 6.18 4.46 4.46 1.14 0.70
   under decentralization)
Notes: In this table, we repeat the same difference-in-differences analysis as in Table 2 Panel B, but using PIR (1934) instead of PIR (1939). In PIR (1934), we observe the cohorts born in 1880–1894 when they are 40 to 54 years old. Sampling rates in PIR (1934) for top income earners are similar to those in PIR (1939): 51% and 40% for the top 0.01% and 0.05% income earners, respectively. All the variables are defined as in Table 2. Parentheses contain p-values based on standard errors clustered at the prefecture level. Square brackets contain wild cluster bootstrap p-values based on standard errors clustered at the cohort level. See Section 4.1 for discussions about this table.
Table A.14: Long-run Impacts of Centralization: Placebo Tests and Pathways
(1) (2) (3) (4)
Placebo: Placebo: Pathway: Pathway:
Population Landlords Fraction Distance
moved moved
in the in the
long-run long-run
A. Baseline Specification
Age 17 under centralization 0.34 -0.10 -0.01 -18.41
   $\times$ Tokyo area ($<$100 km) (0.270) (0.588) (0.449) (0.173)
[0.245] [0.818] [0.422] [ 0.307]
B. Adding Control Variables
Age 17 under centralization -0.06 0.01 -0.02 -23.44
  $\times$ Tokyo area ($<$100 km) (0.872) (0.947) (0.360) (0.173)
[0.790] [0.969] [0.361] [0.259]
Observations 705 705 705 705
Birth cohort FE, Birth pref. FE Yes Yes Yes Yes
Mean dep var 11.67 0.94 0.29 297.92
Mean dep var (Tokyo area 13.18 2.89 0.37 463.52
   under decentralization)
Notes: This table provides placebo tests and explores pathways of the long-run effects. We construct the prefecture-cohort level data from PIR (1939) by counting the number of elites by birth prefecture and by birth cohort (born in 1880–1894). In (1), “Population” is the cohort’s birth population in the birth prefecture. In (2), “Landlords” is defined as individuals listed in the PIR whose occupations include landlord, but excluding the top 0.05% income earners, medal recipients, corporate executives, top politicians and bureaucrats, and Imperial University professors. In (3), “Fraction moved” is defined as the fraction of individuals listed in the PIR whose prefecture of residence is different from his birth prefecture. In (4), “Distance moved” is defined as the average distance between the birth prefecture and the prefecture of residence among individuals listed in the PIR. “Age 17 under centralization” is the indicator variable that takes 1 if the cohort became age 17 under the centralized admissions in 1902–07. “Mean dep var” shows the mean of the dependent variable for all prefecture-cohort observations. “Mean dep var (Tokyo area under decentralization)” shows the mean of the dependent variable in the Tokyo area under the decentralized admissions. In Panel B, we control for time- and cohort-varying prefecture characteristics, i.e., the number of primary schools in the prefecture in the year when the cohort turned eligible age, the number of middle-school graduates in the prefecture in the year when the cohort turned age 17, log GDP of the prefecture when the cohort turned age 20, and the birth population of the cohort in the prefecture (except for Column (1)). Parentheses contain p-values based on standard errors clustered at the prefecture level. Square brackets contain wild cluster bootstrap p-values based on standard errors clustered at the cohort level. ***, **, and * mean significance at the 1%, 5%, and 10% levels, respectively. See Section 4.1 for discussions about this table.
Figure A.8: Years Taken from Entering Higher School to University Graduation
Notes: This figure plots the average number of years taken from entering a higher school to graduating from an Imperial University by cohort (where cohort is defined by the year of university graduation) for two groups of individuals who passed the administrative division of the Higher Civil Service Exams. The first group is top-ranking officials who graduated from Schools 1–8 (shown in black round markers). The second group is non-top-ranking officials who graduated from Schools 1–8 (shown in red rectangular markers for two representative years, 1914 and 1922, only). For both groups, we find the exact year of entering Schools 1–8 for each individual using the Student Registers. The figure shows that the average number of years is not significantly different between the two groups. See Section 4.2 for discussions about this figure.
Figure A.9: Years Taken to Become Top-ranking Government Officials
Notes: This figure plots the distribution of the number of years taken from entering a higher school to becoming a top-ranking government official for representative cohorts. We focus on top-ranking officials who graduated from Schools 1–8 and four randomly selected cohorts (the cohorts entering a higher school in 1903, 1913, 1916, and 1922). For each top-ranking official in each cohort, we look for his biographical information by online searches to find the year of appointment to his first top-ranking position (with a success rate of 82%). The figure shows that, for all cohorts, there is a large within-cohort variation in the number of years taken to be promoted to a top-ranking position. See Section 4.2 for discussions about this figure.
Table A.15: Long-run Impacts on Top-Ranking Government Officials: Alternative Definition of the Year of Entering Higher School or Its Equivalent
(1) (2) (3) (4)
Top-ranking Top-ranking Top-ranking Top-ranking
officials officials officials officials
graduated graduated not
from from graduated
School 1 Schools 2-8 from
Schools 1-8
Centralized 4.15 -0.54 5.05 -0.36
(0.003)*** (0.600) (0.001)*** (0.702)
Observations 33 33 33 33
Control time trend & exam passers Yes Yes Yes Yes
Mean dep var 29.77 8.273 12.97 8.53
Mean dep var (decentralization) 28.57 8.30 11.22 9.05
Notes: This table shows OLS estimates of the long-run effects of the centralized admission system on the number of top-ranking government officials. The definitions of the variables and the specifications are the same as in Table 4, except that the computation of the year of entering a higher school or its equivalent for individuals who did not graduate from Schools 1–8 is changed from “year of graduating the final education - 6” to “year of graduating the final education - 7.” The results are qualitatively the same as the results in Table 4. In all regressions, we control for quadratic time trends and the number of individuals in the cohort who passed the administrative HCSE. Parentheses contain p-values based on Newey-West standard errors with the maximum lag order of 3. ***, **, and * mean significance at the 1%, 5%, and 10% levels, respectively. See Section 4.2 for discussions about this table.
Table A.16: Robustness Check: Government Officials Died in Wars or Purged after WWII
(1) (2) (3) (4)
Exam passers Exam passers Top-ranking Top-ranking
who were who died officials officials
purged in wars who were who died
purged in wars
Centralized 0.854 0.010 0.847 -0.037
(0.767) (0.952) (0.682) (0.214)
Observations 33 33 33 33
Control time trend & exam passers Yes Yes Yes Yes
Mean dep var 15.58 0.61 6.18 0.03
Mean dep var (decentralization) 15.03 0.69 5.88 0.0435
Notes: This table tests if the results in Table 4 are affected by war deaths and post-WWII purge of government officials. Column (1) shows an OLS estimate of the effect of the centralized admissions on the number of individuals who passed the administrative division of the Higher Civil Service Exams and who were purged from public service after WWII by the occupational authority. Column (2) shows an OLS estimate of the effect on the number of HCSE passers who died in wars or were executed after WWII for war crimes. Columns (3) and (4) show OLS estimates of the effects on the number of top-ranking officials who were purged after WWII (Column (3)) or died in wars or were executed after WWII (Columns (4)). The estimates are based on the cohort level data, 1898–1930, where cohort is defined by the year of entering a higher school or its equivalent. “Mean dep var (decentralization)” is the mean of the dependent variable for the cohorts who entered a higher school or its equivalent under the decentralized admissions. In all regressions, we control for quadratic time trends and the number of exam passers. Parentheses contain p-values based on Newey-West standard errors with the maximum lag order of 3. ***, **, and * mean significance at the 1%, 5%, and 10% levels, respectively. See Section 4.2 for discussions about this table.
Figure A.10: Share of Government Officials in PIR-listed Individuals
Notes: This figure plots the share of central government officials in all individuals listed in PIR (1939) by birth cohort. The cohorts are colored according to their exposure to the centralized admissions in 1902–1907, where darker color indicates higher intensity of exposure. We observe no significant increase in the share of central government officials for the cohorts who are exposed to the centralized admissions. See Section 4.2 for discussions about this table.
Figure A.11: Total Number of PIR-listed Individuals by Cohort
Notes: This figure plots the total number of individuals listed in PIR (1939) and PIR (1934) by birth cohort. The cohorts are colored according to their exposure to the centralized admissions in 1902–1907, where the darker color indicates the higher intensity of exposure. In both PIR (1939) and PIR (1934), cohorts exposed to the centralization tend to have a greater number of individuals selected into the PIR. See Appendix Table A.17 for a statistical analysis and Section 4.2 for discussions about this figure.
Table A.17: Long-run Impacts: Total Number of PIR-listed Individuals
(1) (2)
No. of Individuals in PIR
(No. of individuals aged 50 in PIR = 100)
Age 17 under centralization 3.77
(0.086)*
Cohort’s exposure to centralization 4.78
(0.060)*
Age 40.32 39.69
(0.000)*** (0.000)***
Age squared -0.39 -0.38
(0.000)*** (0.000)***
PIR (1939) dummy 0.37 0.37
(0.808) (0.805)
Observations 46 46
Notes: This table examines the long-run effects of the centralized admissions on the total number of career elites in the PIR. To distinguish the cohort effect from the age effect, we use two editions of the PIR, i.e., PIR (1934) and PIR (1939), count the number of individuals aged 40–62 by birth cohort in each edition (see Appendix Figure A.11), and pool these data. We standardize the dependent variable by setting the number of individuals at age 50 in each edition to be 100. “Age 17 under centralization” takes 1 if the cohort turned 17 years old under the centralized system in 1902–1907, and takes 0 otherwise. “Cohort’s exposure to centralization” is defined as in Table 2 Panel D. We control for quadratic age trends and the edition fixed effect. Parentheses contain p-values based on robust standard errors. ***, **, and * mean significance at the 1%, 5%, and 10% levels, respectively. See Section 4.2 for discussions about this table.
Appendix B Online Theoretical Appendix
To guide our empirical investigation, we develop a simple model to predict the impacts of centralization on application behavior and assignment. We first confirm that centralized admissions (Capp) was indeed designed to make the school seat allocation more meritocratic compared to decentralized admissions (Dapp). Our model also predicts changes in application behavior. First, a greater number of applicants apply to the most popular school under Capp than under Dapp (Proposition 1). Second, when high-achieving applicants are disproportionately located in a geographical area (the urban area for example), then a larger fraction of school seats are assigned to students coming from that advantageous area (Proposition 2). Finally, a smaller fraction of applicants are assigned to their local school under Capp relative to Dapp (Proposition 3).
A school admission problem is $(S,I,q,(t_{i})_{i\in I},\succ)$ where $S=\{s_{1},\dotso,s_{m}\}$ is the set of schools while $I=\{i_{1},\dotso,i_{n}\}$ is the set of students. A capacity vector is $q=(q_{s_{1}},\dotso,q_{s_{m}})$ where $q_{s}$ is the number of students school $s$ can accommodate. Motivated by our empirical setting, schools’ common priority order over students is based on test scores $(t_{i})_{i\in I}\in\mathbb{R}_{+}^{n}$ (the higher the better). Without loss of generality, sort students so that $t_{i_{j}}>t_{i_{k}}$ if $j<k$. We also assume that all students are acceptable for any school, which, in our institutional setting, is true conditional on the pool of eligible applicants. The profile of student (strict) reported preferences is $\succ=(\succ_{i_{1}},\dotso,\succ_{i_{n}}$) defined over $S\cup\{o\}$ where $o$ is the outside option. Let $P_{i}$ denote the set of all possible preference relations for student $i$. $P=\times_{i\in I}P_{i}$ is the set of all preference profiles. Let $\succ$, $\succ^{\prime}$ and so on denote students’ reported preference profiles. Motivated by the fact that Schools 1–8 are prestigious national schools with no significant competitors, we assume that every student prefers Schools 1–8 over the outside option: $s\succ_{i}o$ for all $i\in I$ and $s\in S$. We also assume $n>\Sigma_{s\in S}q_{s}$, which is always true in our data.
The outcome of a school admission problem is a matching $\nu:I\to S\cup I$ where $\nu(i)$ means the school that admits student $i$ (or no assignment if $\nu(i)=i$) with the following properties.
  • $\nu(i)\not\in S\implies\nu(i)=i$ for every $i\in I$, and
  • $\big{|}\nu^{-1}(s)\big{|}\leq q_{s}$ for every $s\in S$.
A mechanism is a systematic procedure that determines a matching for each reported preference profile. Formally, it is a function $\mu:P\to\mathcal{M}$ where $\mathcal{M}$ denotes the set of all matchings. Let $\mu_{s}(\succ)$ denote the set of students assigned to $s$ in mechanism $\mu$ for reported preference profile $\succ$. Let $\mu^{C}$ be the Capp mechanism introduced in Section 2.
We compare mechanisms with a thought experiment where the same set of applicants with the same true preferences and test scores participate in different mechanisms. Applicants may change their preference reports, depending on which mechanism they participate in. The set of schools and their capacities are also assumed to stay constant. Index each school seat by $j=1,\dotso,k\equiv\sum_{s\in S}q_{s}$. Let $t_{\mu(\succ)}(j)$ be the test score of the student assigned to seat $j$ under mechanism $\mu$ for preference profile $\succ$. $t_{\mu(\succ)}(j)=0$ if no student is assigned to seat $j$. Let $F_{\mu(\succ)}$ be the cumulative distribution of test scores among assigned students under any mechanism $\mu$ for preference profile $\succ$, defined as
${F}_{\mu(\succ)}(t)=\frac{\big{|}\{j\in\{1,\dotso,k\}\mid t_{\mu(\succ)}(j)\leq t\}\big{|}}{k}$
for all $t\in\mathbb{R}_{+}$.
As should be the case given the official goal of centralization, Capp is more meritocratic than any other mechanism, especially Dapp. Specifically, Capp induces a first-order-stochastic-dominance improvement of the test score distribution among admittees.
Fact 1.
For any school admission problem and any mechanism $\mu$, we have $F_{\mu^{C}(\succ)}(t)\leq F_{\mu(\succ^{\prime})}(t)$ for all $t\in\mathbb{R}_{+}$ and $\succ,\succ^{\prime}\in P$.
This fact implies that the worst test score among assigned students under Capp is weakly better than that under any other mechanism, including Dapp. Fact 1 holds regardless of how applicant behavior changes between mechanisms.
We also confirm that the centralized assignment rule is as meritocratic as the possibly most meritocratic mechanism, i.e., the Deferred Acceptance mechanism. Let $\mu^{I}$ be the mechanism that selects a matching based on the following Student-Proposing Deferred Acceptance algorithm.
  • Step 1. Each student $i$ proposes to her most-preferred school. Each school $s$ holds top $q_{s}$ students and rejects the rest. If less than $q_{s}$ students proposed, then it holds all the students that proposed to $s$.
  • Step $l$. Any student who was rejected at step $l-1$ makes a new proposal to his most-preferred school that has not yet rejected him. If no acceptable choices remain, she makes no proposal. Each school holds its most-preferred $q_{s}$ students to date and rejects the rest. If less than $q_{s}$ students proposed, then it holds all the students who proposed to $s$.
  • The algorithm terminates when there are no more rejections. Each student is assigned to the school that holds her in the last step.
The following property holds:
Fact 2.
$\cup_{s\in S}\mu_{s}^{C}(\succ)=\cup_{s\in S}\mu_{s}^{I}(\succ^{\prime})$ for all $\succ,\succ^{\prime}\in P$.
This result says that the same students are assigned to Schools 1–8 under Capp and the Deferred Acceptance algorithm, which is the most meritocratic mechanism we can design. Like Fact 1, Fact 2 holds regardless of applicant behavior.
To derive additional predictions about applicant behavior, we impose more structures on the model. Each applicant takes an action under each mechanism $\mu$. Under Capp, for example, each applicant submits a preference list $\succ_{i}$. Under Dapp, each applicant applies to a school. The mechanism then uses these actions to obtain a matching. This procedure induces a strategic form game, $\langle I,(A_{i})_{i\in I},\succ^{o}\rangle$. The set of players is the set of applicants $I$. The action space of each applicant is $A_{i}$. Under Capp, this is the set of all possible preference relations $P_{i}$ over schools. Under Dapp, this is the set of schools $S=\{s_{1},...,s_{m}\}$. The outcome is evaluated through the true preferences $\succ^{o}=(\succ^{o}_{i_{1}},\dotso,\succ^{o}_{i_{n}})$. A strategy vector $a=(a_{1},\dotso,a_{n})$ is an equilibrium if for each applicant $i\in I$ and each strategy $a^{\prime}_{i}\in A_{i}$, we have $\mu_{i}(a)\succeq_{i}\mu_{i}(a^{\prime}_{i},a_{-i}).$ We say that strategy $a_{i}$ is a (weakly) dominant strategy if we have $\mu_{i}(a)\succeq_{i}\mu_{i}(a^{\prime}_{i},a_{-i})$ for all $a^{\prime}_{i}\in A_{i}$ and $a_{-i}\in A_{-i}$.
Proposition 1.
Suppose that applicants share homogeneous preferences $s_{1}\succ^{o}_{i}s_{2}\succ^{o}_{i}...\succ^{o}_{i}s_{n}\succ^{o}_{i}\emptyset$. Also assume that every applicant submits the true preference whenever it is a dominant strategy. In any equilibrium, the number of applicants who apply to the most preferred school $s_{1}$ is weakly larger under centralized admissions than under decentralized admissions.
This assumption is motivated by our institutional setting, where applicants share a perceived vertical hierarchy over schools. Intuitively, Capp would incentivize applicants to give a shot at the most prestigious and selective school since Capp gives applicants a chance of acceptance by lower-choice schools even if they are rejected by the first-choice school.
Assume that each applicant lives in the local area of a school. Let $n_{j}$ be the number of students from school $s_{j}$’s area. Interpret $s_{1}$ as an advantageous urban school whose the test score distribution first order stochastically dominates that in any other school’s area.
Proposition 2.
In any equilibrium, a larger share of all school seats are allocated to urban-born applicants from $s_{1}$’s area under centralized admissions than under decentralized admissions.
[…]
As above, an admission mechanism induces a strategic form game $\langle I,(A_{i})_{i\in I},(U_{i})_{i\in I}\rangle$. The set of players and the action space remain the same. The outcome is now evaluated accordingly to cardinal utility. Define $U_{i}(.)$ as the expected payoff of player $i$ at the application stage. For Dapp, for example, $U_{i}(a_{i},a_{-i})=\min\{\dfrac{q_{a_{i}}}{\sum_{j\in I}1\{a_{j}=a\}},1\}\times U_{i{a_{i}}}$ if he plays action $a_{i}$, given action profile $a_{-i}$ of other applicants. An equilibrium $(a_{1},\dotso,a_{n})$ is called a symmetric equilibrium if $a_{i}=a_{j}$ for all $i$ and $j$ from the same area. We make the following assumptions for the rest of this section: Applicants play a symmetric equilibrium, which is assumed to exist.
For a given mechanism and an equilibrium play, $q_{jk}$ denotes the number of applicants assigned to school $s_{j}$ who come from school $s_{k}$’s area. Define the proportion of assigned applicants assigned to their local school as
$\dfrac{{q}_{11}+{q}_{22}}{{q}_{11}+{q}_{12}+{q}_{21}+{q}_{22}}.$
Proposition 3.
For sufficiently large $V$ or sufficiently large $|U_{1}-U_{2}|$, the proportion of assigned applicants assigned to their local school is higher under Dapp than under Capp.
Capp therefore reduces the number of local entrants born in the school’s prefecture.
B.1 Proofs
Proof of Fact 1. As required in step 1 of Capp, school seats are assigned to applicants $i_{1},\dotso,i_{k}$ under $\mu^{C}$, i.e., $\cup_{s\in S}\mu_{s}^{C}(\succ)=\{i_{1},\dotso,i_{k}\}$ and $\cup_{j\in\{1,...,k\}}\{t_{\mu^{C}(\succ)}(j)\}=\{t_{i_{1}},\dotso,t_{i_{k}}\}$. Here we use the assumption that $s\succ_{i}o$ for all $i\in I$ and $s\in S$. Let $\cup_{s\in S}\mu_{s}(\succ^{\prime})=\{i_{j_{1}},\dotso,i_{j_{l}}\}$ with $k\leq l$, $j_{1}<\dotso<j_{l}$ and $\{j_{1},\dotso,j_{l}\}\subseteq\{1,\dotso,n\}$. This gives $\cup_{j\in\{1,...,k\}}\{t_{\mu(\succ^{\prime})}(j)\}=\{t_{i_{j_{1}}},\dotso,t_{i_{j_{l}}},\cup_{i=1}^{k-l}\{0\}\}$. Since $t_{i_{1}}\geq t_{i_{j_{1}}}$,…, $t_{i_{l}}\geq t_{i_{j_{l}}}$, we have |$\big{|}\{j\in\{1,\dotso,k\}\mid t_{\mu^{C}(\succ)}(j)\leq t\}\big{|}\leq\big{|}\{j\in\{1,\dotso,k\}\mid t_{\mu(\succ^{\prime})}(j)\leq t\}\big{|}$ so that
${F}_{\mu^{C}(\succ)}(t)=\frac{\big{|}\{j\in\{1,\dotso,k\}\mid t_{\mu^{C}(\succ)}(j)\leq t\}\big{|}}{k}\leq\frac{\big{|}\{j\in\{1,\dotso,k\}\mid t_{\mu(\succ^{\prime})}(j)\leq t\}\big{|}}{k}={F}_{\mu(\succ^{\prime})}(t).$
Therefore, we have that $F_{\mu^{C}(\succ)}(t)\leq F_{\mu(\succ^{\prime})}(t)$ for all $t\in\mathbb{R}_{+}$ and $\succ,\succ^{\prime}\in P$.∎
Proof of Fact 2. As required in step 1 of Capp, school seats are assigned to applicants $i_{1},\dotso,i_{k}$ under $\mu^{C}$, i.e., $\cup_{s\in S}\mu_{s}^{C}(\succ)=\{i_{1},\dotso,i_{k}\}$ for all $\succ\in P$. Again, we use the assumption that $s\succ_{i}o$ for all $i\in I$ and $s\in S$. Under the assumption that any school is acceptable for any student, any student $i_{k^{\prime}}$ with $k^{\prime}>k$ will be rejected at some step of the student-proposing Deferred Acceptance algorithm. Therefore, the top $k$ students are assigned to some school under $\mu^{I}$, i.e., $\cup_{s\in S}\mu_{s}^{I}(\succ^{\prime})=\{i_{1},\dotso,i_{k}\}$ for all $\succ^{\prime}\in P$. Therefore, $\cup_{s\in S}\mu_{s}^{C}(\succ)=\cup_{s\in S}\mu_{s}^{I}(\succ^{\prime})=\{i_{1},\dotso,i_{k}\}$ for all $\succ,\succ^{\prime}\in P$.∎
Proof of Proposition 1. This result follows from the following intermediate facts.
Lemma 1.
Under Dapp, it is the only equilibrium that
  • the $q_{s_{1}}$ highest-scoring applicants apply to $s_{1}$,
  • the next $(q_{s_{1}}+1)$ to $(q_{s_{1}}+q_{s_{2}})$ highest-scoring applicants apply to $s_{2}$,
  • the next $(\sum^{s_{n-1}}_{s=s_{1}}q_{s}+1)$ to $(\sum^{s_{n}}_{s=s_{1}}q_{s})$ highest-scoring applicants apply to $s_{n}$.
Proof.
The proof is inductive. The $q_{1}$ highest-scoring applicants are assigned to the most preferred school $s_{1}$, so there is no incentive for them to deviate from this equilibrium. Given the behavior of the $q_{s_{1}}$ highest-scoring applicants, the next $(q_{s_{1}}+1)$ to $(q_{s_{1}}+q_{s_{2}})$ highest-scoring applicants cannot get the most preferred school $s_{1}$. They are assigned to the second most preferred school $s_{2}$, so there is no incentive for them to deviate from this equilibrium. Similar arguments show that the above behavior is an equilibrium. It is also the only equilibrium since any difference from the above for any applicant would give the applicant a profitable deviation with which the applicant could get the same result as in the above equilibrium. ∎
Lemma 2.
Under Capp, ranking $s_{1}$ is a dominant strategy at least for the $q_{s_{1}}$ highest-scoring applicants.
Proof.
It is because they would be assigned to the most preferred school $s_{1}$ whenever ranking $s_{1}$ first. ∎
The above lemmas jointly imply that in any equilibrium, the number of applicants who apply to the most popular school $s_{1}$ is weakly larger under centralized admissions than under decentralized admissions.∎
Proof of Proposition 2. Under Capp, by the meritocracy constraint and the assumption that $s\succ_{i}o$ for all $i\in I$ and $s\in S$, the top $k$ highest score students are assigned to any school. Count how many of the top $k(\equiv\Sigma^{s_{m}}_{s=s_{1}}q_{s})$ students are from $s_{1}$’s area. Under Dapp, in any equilibrium, any of these students from $s_{1}$’s area is also assigned to one of the schools. To show this, suppose not. There is student $i^{*}$ from another area whose test score is below the top $k$ scores. Let $s_{i^{*}}$ be the school that admits $i^{*}$. If any unassigned, top $k$ student $i_{1}$ from $s_{1}$’s area apply to $s_{i^{*}}$, $i_{1}$ would be assigned to $s_{i^{*}}$, a contradiction to the definition of an equilibrium. ∎
Proof of Proposition 3. We use the following lemma.
Lemma 3.
For sufficiently large $V$, all applicants apply to their local schools in any symmetric equilibrium under Dapp.
Proof of Lemma 3. First, we show that for sufficiently large $V$, none of the following symmetric equilibrium survives: (i) all applicants apply to $s_{i}$ (for $i=1,2$), or (ii) applicants from $s_{1}$’s area apply to $s_{2}$ while those from $s_{2}$’s area apply to $s_{1}$. Define $p(n,q)$ as the probability of being one of the top $q\in\mathbb{N}$ applicants among $n\in\mathbb{N}$ applicants as per i.i.d test scores, i.e., $p(n,q)=\min\{\dfrac{q}{n},1\}$. Let $\bar{n}_{a}=\sum_{j\in I}1\{a_{j}=a\}$.
Case (i). Applicants from school $j$’s area apply to $s_{i}$ if $p(n_{i}+n_{j},q_{s_{i}})\times U_{i}\geq U_{j}+V$. Since $V>p(n_{i}+n_{j},q_{s_{i}})\times U_{i}-U_{j}$, therefore, all applicants applying to $s_{i}$ (for $i=1,2$) cannot be a symmetric equilibrium.
Case (ii). Suppose applicants from $s_{1}$’s area apply to $s_{2}$ while those from $s_{2}$’s area apply to $s_{1}$. For applicants from $s_{1}$’s area, it must be the case that $p(n_{1},q_{s_{2}})\times U_{2}\geq p(n_{2}+1,q_{1})*(U_{1}+V)$. For applicants from $s_{2}$’s area, it must be the case that $p(n_{2},q_{s_{1}})\times U_{1}\geq p(n_{1}+1,q_{s_{2}})\times(U_{2}+V)$. For sufficiently large $V$, these inequalities cannot hold.
Now we show that, for large enough $V$, all students applying to their local schools is indeed a symmetric equilibrium. For applicants from $s_{1}$’s area to apply to $s_{1}$, it must be the case that $p(n_{1},q_{s_{1}})\times(U_{1}+V)\geq p(n_{2}+1,q_{s_{2}})\times U_{2}$. For applicants from $s_{2}$’s area to apply to $s_{2}$, it must be the case that $p(n_{2},q_{s_{2}})\times(U_{2}+V)\geq p(n_{1}+1,q_{s_{1}})\times U_{1}$. Since the left hand sides of both the inequalities are increasing in $V$, the equilibrium conditions hold for sufficiently large $V$. ∎
From Lemma 3, for sufficiently large $V$, applicants apply to their locals schools under Dapp. Therefore, the expected proportion of assigned applicants assigned to their local school under Dapp is 1 (the highest). With sufficiently large $|U_{1}-U_{2}|$, every applicant ranks $s_{1}$ first in any equilibrium under either mechanism, so the same appplicants are assigned to $s_{1}$ and $s_{2}$, meaning that the expected proportion of assigned applicants assigned to their local school stays the same. This completes the proof.
[…]

### v2

@@
 What happens if selective colleges change their admission policies?
-We answer this question by analyzing the world’s first introduction of nationally centralized meritocratic admissions in the early twentieth century.
+We study this question by analyzing the world’s first implementation of nationally centralized meritocratic admissions in the early twentieth century.
 We find a persistent meritocracy-equity tradeoff.
-Compared to the decentralized system, the centralized system admitted more high-achieving applicants, producing a greater number of top elite bureaucrats in the long run.
-However, this impact came at the distributional cost of equal regional access to elite higher education and career advancement.
-Several decades later, the meritocratic centralization increased the number of urban-born career elites (e.g., top income earners) relative to rural-born ones.
+Compared to the decentralized system, the centralized system admitted more high-achievers and produced more occupational elites (such as top income earners) decades later in the labor market.
+This gain came at a distributional cost, however.
+Meritocratic centralization also increased the number of urban-born elites relative to rural-born ones, undermining equal access to higher education and career advancement.
 Keywords: Elite Education, Market Design, Strategic Behavior, Regional Mobility, Universal Access, Persistent Effects
+JEL: D47, I23, I24, N35
 1 Introduction
-College and school admission processes vary across time and place.
-American college admissions use a decentralized system where each college makes its own admissions often based on opaque evaluation criteria.
-In contrast, China’s public college admissions illustrate a regionally-integrated, single-application, and single-offer system using a transparent admission criterion.
-How do different admissions institutions affect students’ education and future career?
-What are the impacts on the national production of high-skilled individuals and their distributions?
-In this paper, we study the impacts of using centralized and meritocratic admissions.
-We do so by combining a series of natural experiments in history, newly assembled historical data, and theoretical insights.
-Our investigations reveal the pros and cons of centralized meritocratic admissions, especially a tradeoff between meritocracy and equal access to selective higher education and career achievements.
-Our empirical setting is the first known transition from decentralized to nationally-centralized school admissions.
+One major trend in college admissions around the world is a growing degree of centralization.
+Today, over thirty countries use regionally- or nationally-integrated, single-application and single-offer college admissions (Appendix Figure A.1).
+These systems have well-specified admission criteria, mixing meritocratic achievement elements (such as GPA and entrance exams) and other priority considerations.
+Before the turn of the 20th century, however, no country used such a centralized system.
+How does the centralization of admission systems affect students’ life trajectories?
+What are their impacts on the national production of highly skilled individuals and their composition?
+A key challenge in studying these questions is the lack of clear policy changes and data about students’ long-run outcomes.
+In this paper, we study the impacts of centralized and meritocratic college admissions.
+Our investigations reveal their pros and cons, especially a tension between meritocracy and equal access to higher education and career achievements in the long run.
+We reach these findings by combining a series of natural experiments in history and newly assembled historical data that trace students over decades.
+Our empirical setting is the world’s first known transition from decentralized to nationally-centralized school admissions.
 At the end of the 19th century, to modernize its higher education system, the Japanese government set up National Higher Schools (roughly equivalent to today’s colleges) as an exclusive entry point to the most prestigious tertiary education.
-[^1] These schools were transformed into parts of today’s national universities after WWII.
-For example, the National Higher School in Tokyo became the first two years of the University of Tokyo.
-These schools later produced many of the most influential members of the society, including several Prime Ministers, Nobel Laureates, and founders of global companies like Toyota.
-Their selective and prestigious nature and social influences had similarities to Oxbridge in the UK, Grande Ecoles in France, and the Ivy League in the US.
-[^2] We study top elites on both substantive and pragmatic grounds.
-The individual holdings of prestigious economic, political, and cultural positions are skewed and concentrated on a tiny minority of the population who have big influences on the economy and society.
-For their social importance, we also have relatively rich historical data that allows us to trace back their careers from humble beginnings as students to wide societal recognition and acclaim.
-Acceptance into these schools was merit-based, using annual entrance examinations.
+[^1] After WWII, National Higher Schools were transformed into parts of today’s national universities.
+For example, the First National Higher School in Tokyo became the first two years of the University of Tokyo.
+We focus on these selective colleges, as they later produced many of the most influential leaders of the society, including several Prime Ministers, Nobel Laureates, and founders of global companies like Toyota.
+Their prestige and social influences were similar to Oxbridge in the UK, Grandes Écoles in France, and the Ivy League in the US.
+Acceptance into these schools was based on annual entrance exams.
 Initially, the government let each school run its own exam and admissions based on exam scores, similar to many of today’s decentralized K-12 and college admissions.
-Under this decentralized system, the schools typically held exams on the same day so that each applicant could apply for only one school.
+Under this decentralized system, the schools typically held exams on the same date so that each applicant could apply for only one school.
 [^2] Similar restrictions on the number of applications exist today in the college admission systems of Italy, Nigeria, and the UK.
-As a result, many high achieving applicants applied to too competitive schools and failed to enter any school.
-This problem motivated the government to introduce a centralized system in order to allocate the limited school seats to the highest-achieving students.
-In the new system introduced at the turn of the 20th century, applicants were allowed to rank multiple schools in the order of their preferences and take a single unified exam.
+As a result, many high achieving applicants applied to and were rejected by the most competitive school, failing to enter any school.
+This problem motivated the government to introduce a centralized system in 1902.
+The new system prioritized the admission of the highest-scoring students.
+Specifically, applicants were asked to submit preference rankings over multiple schools and to take a single unified exam.
 [^3] As shown later, the defining feature of the centralized system was to allow applicants to list multiple schools, not the use of a single unified exam.
-Given their preferences and exam scores, each applicant is assigned to a school (or none if unsuccessful) based on a computational algorithm.
-The algorithm was a mix of the Deferred Acceptance and Immediate Acceptance (Boston) algorithms combined with a meritocracy principle that only assigns the highest-scoring applicants to any school.
+Given their preference rankings and exam scores, each applicant was assigned to a school (or none if unsuccessful) based on a computational algorithm.
+The algorithm was a mix of the Deferred Acceptance and Immediate Acceptance (Boston) algorithms combined with a meritocracy principle that assigns only the highest-scoring applicants to any school.
 To our knowledge, this instance is the first nationwide use of any matching algorithm.
@@
 Furthermore, for reasons detailed below, the government later re-decentralized and re-centralized the system several times, producing multiple natural experiments for studying the consequences of the different systems.
-[^6] During 1900–1930, there were three periods of centralization (1902–07, 1917–18, and 1926–27).
-We exploit these bidirectional institutional changes to identify the short- and long-run impacts of the meritocratic centralization.
-A stylized theoretical model predicts the short-run effects of the reforms on application behavior and admission outcomes.
-Consistent with its stated goal, the model confirms that the centralized system produces more meritocratic school seat allocations than the decentralized system.
-It also predicts that centralization would cause applicants to apply to more selective schools and make more inter-regional applications.
-To test these predictions and analyze the effects of meritocratic centralization, we newly collect and digitize data about application and enrollment.
-The data is based on administrative school records, Government Gazettes, and Ministry of Education Yearbooks.
-Our data records each student’s regional origin, which is a strong predictor of socioeconomic status both in our empirical setting and other settings like the contemporary US.
-We find that meritocratic centralization had large and immediate effects on application and enrollment.
-First, consistent with the theoretical predictions, centralization caused stark strategic responses in application behavior.
-In particular, strategic incentives in the centralized system led applicants across the country to more often rank the most selective school first.
-This effect and the option to apply to multiple schools made students enroll at schools further away from their regional origin.
-The reform thus made the elite higher education market more meritocratic and regionally integrated.
-The gains from centralization were not equally distributed, however.
-Because high-achieving students were disproportionately located in urban areas (mainly Tokyo), the centralized system caused a greater number of urban applicants to be admitted to schools in rural areas, typically after being rejected by their first choice schools.
-[^7] We use the nomenclature of “urban” and “rural” schools, but note that “rural” schools were located in regional cities rather than in the countryside.
-As a result, urban high-achievers crowded out rural applicants; the number of entrants to any national elite school coming from the urban area increased by about 10% during centralization.
+[^5] During 1900–1930, there were three periods of centralization in 1902–1907, 1917–1918, and 1926–1927.
+We exploit these bidirectional institutional changes to identify the short- and long-run impacts of meritocratic centralization.
+For the short-run analysis, we newly collect and digitize application and enrollment data from administrative school records and government documents, including Government Gazettes and Ministry of Education Yearbooks.
+We first find that meritocratic centralization had large effects on application behavior and enrollment outcomes.
+In particular, centralization caused applicants in all areas to be more risk-taking and more often rank the most selective school first.
+Applicants also made longer-distance applications, leading students to be enrolled at schools further away from their regional origin.
+The reform thus made the selective higher education market more meritocratic, competitive, and regionally integrated.
+The effect of centralization was heterogeneous, however.
+Because high-achieving students were disproportionately located in urban areas (mainly Tokyo), the centralized system caused a greater number of urban applicants to be admitted to schools in rural areas, typically after being rejected by their first-choice schools.
+[^6] We use “urban areas” to refer to a region surrounding Tokyo (as defined in Section 3).
+As shown later, these areas were characterized by a greater population, higher income, and better educational infrastructure.
+As a result, urban high-achievers crowded out rural applicants: the number of urban-born entrants to any national higher school increased by about 10% during centralization.
 [^7] It is also empirically true that the centralized system made a greater number of rural applicants apply to and enter urban schools.
@@
 But their net effects are such that urban high-achievers crowded out rural applicants.
-Historical documents suggest that this distributional consequence upset rural schools and communities.
-Partly as a result of such rural discontents, the government went back and forth between decentralized and centralized systems, finally settling on a decentralized scheme.
-This series of bidirectional reforms enables us to identify the causal effects of meritocratic centralization more precisely than a usual, single policy change would.
-[^9] We provide evidence that the reform timings appear exogenous in that they are not correlated with other social and institutional outcomes.
-Our main results are about the long-run effects of the centralized system.
-We start by looking at the persistence of the distributional impacts.
-The short-run analysis finds that centralization let urban areas disproportionately gain school access relative to rural areas.
-This result motivates a difference-in-differences analysis that compares long-term career outcomes of urban vs rural-born individuals by each cohort’s exposure to the centralized system.
-Our long-run career data comes from the Personnel Inquiry Records (PIR) published in 1939, more than thirty years after the first episode of meritocratic centralization in 1902–07.
-This data provides a list of distinguished individuals (e.g., high-income earners, national medal recipients, high-ranking politicians, corporate executives) along with their biographical information.
-We provide extensive investigations about the quality of the data.
-[^10] We find that the data covers a large fraction of the national population of elites (e.g., 53% of the top 0.01% income earners) by comparing our data with national administrative statistics.
-We also find no systematic variation in the sampling rates across prefectures, consistent with our assumption that sample selection bias is uncorrelated with the prefecture-cohort variation we use.
+Historical documents suggest that this distributional consequence upset rural schools and their local communities.
+Such rural discontents were a reason why the government oscillated between the centralized and decentralized systems.
+Our main results are the long-run effects of the centralized system on the national production and regional distribution of occupational elites.
+To study the distributional consequence, we compare urban vs rural-born individuals’ long-term career outcomes by each cohort’s exposure to the centralized admission system.
+The career outcome data is based on the Personnel Inquiry Records (PIR) published in 1939 (more than thirty years after the first period of meritocratic centralization).
+The data provides a list of socially distinguished individuals—encompassing economic, political, and cultural elites—and their biographical information.
+[^8] We provide extensive investigations about the quality of the data (such as the coverage and sampling bias) in Section 4.1.
 Distributional effects of meritocratic centralization turn out to be persistent.
-Almost four decades later, relative to the decentralized system, the centralized system produced a greater number of top income earners, prestigious medal recipients, and other elite professionals who came from urban areas compared to rural areas.
-The number of urban-born career elites increased by 10–20% for the cohorts exposed to meritocratic centralization.
-We also obtain suggestive evidence that the centralized system increased the number of career elites living in urban areas as adults in the long run.
-We conclude the long-run analysis by examining the overall productive efficiency implication of meritocratic centralization for the whole country.
-We focus on a specific subgroup of career elites — top bureaucrats.
-Top bureaucrats were an exemplar category of career elites in our empirical context, as the best and brightest graduates from the top colleges competed to become bureaucrats.
-Using the comprehensive administrative data of all individuals who passed the national exams to become higher civil officials, we compare the national number of bureaucrats promoted to top ranks between cohorts exposed to the centralized admissions and cohorts exposed to the decentralized admissions.
-We find that the number of top-ranking bureaucrats in the centralized cohorts was 15% greater than that in the decentralized cohorts, suggesting that centralized cohorts won promotion competitions over close decentralized cohorts.
-Meritocratic centralization thus produced a larger number of high-quality bureaucratic elites for the whole country.
-Overall, our findings reveal an equity-meritocracy tradeoff both in the short and long run.
-On the one hand, the meritocratic centralization of school admissions achieved the goal of rewarding applicants with higher academic performance in the short run.
-This meritocracy resulted in a greater number of top bureaucratic elites produced by the exposed cohorts in the long run.
-On the other hand, this meritocracy came at the equity cost of urban-born high-achieving applicants overwhelming rural-born ones.
-The design of admission rules therefore affects the production and distribution of highly educated and skilled individuals, which is an important determinant of economic growth and inequality (Glaeser, 2011; Moretti, 2012).
+Almost four decades after the reform, relative to the decentralized system, the centralized system produced a greater number of occupational elites (such as top income earners and top politicians and bureaucrats) who came from urban areas compared to rural areas.
+The number of urban-born elites increased by 10–20% for the cohorts exposed to meritocratic centralization.
+We also provide suggestive evidence that the centralized system increased the number of elites living in urban areas in their late adulthood.
+Meritocratic centralization thus impacted both the origins and destinations of highly skilled individuals.
+We conclude our analysis by examining the impact of meritocratic centralization on the national production of occupational elites.
+We first focus on a specific subgroup, i.e., top bureaucrats.
+Top bureaucrats were an important category of elites in our empirical context, where higher civil service was one of the most prestigious and coveted jobs due to their political and legislative influence in building the state (Shimizu, 2019).
+Using comprehensive administrative data of all individuals who have passed the national exams to become higher civil officials, we compare the total number of bureaucrats promoted to top ranks between cohorts exposed to the centralized admissions and cohorts exposed to the decentralized admissions.
+We find that the number of top-ranking bureaucrats was 15% greater in the centralized cohorts than in the decentralized cohorts, suggesting that centralized cohorts won promotion competitions over decentralized cohorts.
+Moreover, we find that meritocratic centralization produced a greater total number of all occupational elites listed in the PIR data, including top-income earners.
+Meritocratic centralization thus produced a greater number of occupational elites for the whole country.
+Overall, our findings show that the design of admission rules affects the production and distribution of highly educated and skilled individuals, which is an important determinant of economic growth and inequality (Glaeser, 2011; Moretti, 2012).
+On one hand, meritocratic centralization achieved the goal of recruiting applicants with higher academic performance.
+Moreover, this meritocracy produced a greater number of occupational elites in the long run.
+The policymaker may therefore embrace meritocratic centralization on the ground that it produces more high-quality leaders for the society.
+On the other hand, this gain came at the equity cost of urban-born high-achievers overwhelming rural-born ones both in the short and long run.
+Meritocratic centralization expanded the urban-rural gap in the ability to produce highly skilled individuals.
+Meritocracy may thus dampen the society’s capability to diversify the backgrounds of individuals in leadership positions and represent a broad range of socioeconomic groups.
 Literature.
-Our analysis contributes to the empirical literature that uses policy reforms to study the impacts of the design of admission systems (Abdulkadiroğlu et al., 2006; Abdulkadiroğlu, Agarwal and Pathak, 2017; Chen, Jiang and Kesten, 2020; Machado and Szerman, 2021; Grenet, He and Kübler, 2022; Kapor, Karnani and Neilson, 2023).
-[^11] Other studies measure the effects of selective schools conditional on a particular admission system (Dale and Krueger, 2002; Hastings, Neilson and Zimmerman, 2013; Pop-Eleches and Urquiola, 2013; Deming et al., 2014; Kirkeboen, Leuven and Mogstad, 2016; Clark and Del Bono, 2016; Abdulkadroǧlu et al., 2017; Zimmerman, 2019; Abdulkadiroglu et al., 2021; Michelman, Price and Zimmerman, 2021; Chetty, Deming and Friedman, 2023).
+Our analysis contributes to the empirical literature that uses policy reforms to study the impacts of the design of admission systems (Abdulkadiroğlu et al., 2006; Abdulkadiroğlu, Agarwal and Pathak, 2017; Chen, Jiang and Kesten, 2020; Machado and Szerman, 2021; Terrier, Pathak and Ren, 2021; Kapor, Karnani and Neilson, 2023; Campos and Kearns, 2024).
+[^9] Other studies measure the effects of selective schools conditional on particular admission systems (Dale and Krueger, 2002; Hastings, Neilson and Zimmerman, 2013; Pop-Eleches and Urquiola, 2013; Deming et al., 2014; Kirkeboen, Leuven and Mogstad, 2016; Clark and Del Bono, 2016; Abdulkadroǧlu et al., 2017; Zimmerman, 2019; Abdulkadiroglu et al., 2021; Narita, 2021; Michelman, Price and Zimmerman, 2021; Chetty, Deming and Friedman, 2023).
 While these prior studies focus on the short-run effects (on application, enrollment, academic achievement, and welfare), we estimate the long-run effects by taking advantage of bidirectional, repeated policy changes in history.
-[^12] With its interest in long-run effects, this paper also relates to studies of the long-term effects of educational reforms (e.g., Duflo, 2001; Meghir and Palme, 2005).
+[^10] With its interest in long-run effects, this paper also relates to studies of the long-term effects of educational reforms (e.g., Duflo, 2001; Meghir and Palme, 2005; Ichimura et al., 2024).
 These studies focus on the effects of expanding resources (such as school constructions and compulsory education extensions), while we investigate the effects of changing resource allocation mechanisms given a fixed amount of resources.
@@
 Our analysis of the meritocracy vs equity tension has some similarities to ongoing policy discussions on affirmative actions, which are surveyed in Arcidiacono and Lovenheim (2016).
-Recent contributions in this area include Bleemer (2020, 2022), Kapor (2020) and Otero, Barahona and Dobbin (2021).
+Recent important contributions in this area include Bleemer (2020, 2022), Kapor (2020) and Otero, Barahona and Dobbin (2021).
 [^11] Kamada and Kojima (2015) and Agarwal (2017) study regional equity in other matching markets.
-From a broader historical perspective, this paper relates to the literature that investigates the long-term effects of institutions (Nunn, 2009), particularly resource allocation mechanisms (Bleakley and Ferrie, 2014, 2016).
+From a broader historical perspective, this study relates to the literature that investigates the long-term effects of institutions (Nunn, 2009), particularly the effects of resource allocation mechanisms (Bleakley and Ferrie, 2014).
 Our analysis is also related to Bai and Jia (2016), who examine political consequences of the abolition of a meritocratic elite recruitment system (civil service exam) in early twentieth-century China.
-The next section provides the institutional background.
+The next section provides the historical and institutional background.
 Section 3 examines the short-term effects of meritocratic centralization, while Section 4 analyzes the long-term impacts.
-Finally, Section 5 summarizes our findings and outlines future directions.
+Section 5 summarizes our findings and outlines future directions.
 2 Background
-One major trend in modern college admissions is a growing degree of centralization with transparent admission criteria.
-Today, over 30 countries use regionally- or nationally-integrated, single-application and single-offer college admissions.
-Centralized college admissions are used in all continents except North America (Appendix Figure A.1).
-These systems have well-specified admission criteria, mixing meritocratic achievement elements (such as GPA and entrance exams), and affirmative action and other priority considerations.
-Before the turn of the 20th century, however, no country used such a centralized system.
-Even today, many countries, including the US, continue to use decentralized systems.
-Such decentralized schemes tend to come with less transparent criteria for ranking applicants, as illustrated by recent court cases against American universities.
-Similar observations also apply to K-12 school admissions.
-How does the centralization of college and school admissions affect students’ application behavior, enrollment outcomes, and future careers?
-Understanding the costs and benefits of meritocratic centralization is the goal of our paper.
 2.1 Bidirectional Admission Reforms
-To evaluate the impacts of centralized admission systems, we take advantage of unique historical episodes in early twentieth-century Japan.
+To analyze the impact of centralized admissions, we take advantage of unique historical episodes in early twentieth-century Japan.
 Since the opening of the country in the mid 19th century, to catch up with Western knowledge, science, and technologies, education reforms became a central part of modernization efforts by the Japanese government.
-In 1894, the government set up a national higher education system consisting of one Imperial University and five National Higher Schools.
-By 1908, the system was expanded to four Imperial Universities (Tokyo, Kyoto, Tohoku, and Kyushu) and eight National Higher Schools located across different areas in Japan (First in Tokyo, Second in Sendai, Third in Kyoto, Fourth in Kanazawa, Fifth in Kumamoto, Sixth in Okayama, Seventh in Kagoshima, and Eighth in Nagoya, named after the order of establishment; see Appendix Figure A.2 for their locations).
+In 1894, the government set up a national higher education system consisting of one Imperial University (four-year program) and five National Higher Schools (five-year program).
+By 1908, the system was expanded to four Imperial Universities (Tokyo, Kyoto, Tohoku, and Kyushu) and eight National Higher Schools located across the country.
+[^12] First in Tokyo, Second in Sendai, Third in Kyoto, Fourth in Kanazawa, Fifth in Kumamoto, Sixth in Okayama, Seventh in Kagoshima, and Eighth in Nagoya, named after the order of establishment; see Appendix Figure A.2 for their locations.
 We refer to these eight National Higher Schools as Schools 1–8 for short.
@@
 In our empirical analyses, we control for the number of National Higher Schools as well as other characteristics of higher education institutions.
-We focus on Schools 1–8.
+Our analysis focuses on Schools 1–8.
 Schools 1–8 served as an exclusive entry point to Imperial Universities, the most prestigious tertiary education.
@@
 As a result, entering Schools 1–8 was considered as a passport into the elite class.
-Indeed, Schools 1–8 produced highly distinguished and influential individuals, including several Prime Ministers, Nobel Laureates, world-leading mathematicians, renowned novelists, and founders of global companies like Toyota.
+Indeed, Schools 1–8 produced highly distinguished and influential individuals, including Prime Ministers, Nobel Laureates, world-leading mathematicians, renowned novelists, and founders of global companies like Toyota.
 To apply to these schools, one must be male aged 17 or older and have completed a five-year middle school.
@@
 Schools 1–8 admitted fewer than 2,300 students each year throughout 1900–1930, constituting less than 0.5% of the cohort of males aged 17.
-More descriptive statistics are presented in Appendix Table A.2.
-Among Schools 1-8, School 1 in Tokyo was considered the most prestigious due to its location in the capital and its geographical proximity to Tokyo Imperial University (today’s University of Tokyo).
+Descriptive statistics are presented in Appendix Table A.2.
+Among Schools 1–8, School 1 in Tokyo was considered the most prestigious due to its location in the capital and its geographical proximity to Tokyo Imperial University (today’s University of Tokyo).
 The next most prestigious was School 3 in Kyoto, the ancient capital of Japan.
@@
 Consequently, the schools differed substantially in their popularity and selectiveness.
-[^16] Tuition was mostly uniform across Schools 1-8.
-The admission to Schools 1–8 was merit-based and determined by annual entrance exams.
+[^15] Tuition was mostly uniform across Schools 1–8.
+The admission to Schools 1–8 was based on annual entrance exams.
 Initially, the government let each school administer its own exam and admissions.
-Schools 1–8 typically held their exams on the same day so that each applicant could only apply to one school.
-Following the convention in the literature (Che and Koh, 2016; Hafalir et al., 2018), we call this system “decentralized admissions” or Dapp for short.
-The single choice aspect of Dapp captures an essential feature of decentralization, which incentivizes each applicant to self-select into an appropriate school by comparing the selectivity of schools with his own standing.
-Under Dapp, however, many high-achieving students were rejected by popular schools, while lower-achieving students entered less popular schools.
+Typically, Schools 1–8 coordinated to hold their exams on the same date so that each applicant could only apply to one school.
+We call this system “decentralized admissions” (or “decentralization” for short).
+The single choice aspect captures an essential feature of decentralization, which incentivizes each applicant to self-select into an appropriate school by comparing the selectivity of schools with his own standing (Che and Koh, 2016).
+Under decentralization, however, many high-achieving students were rejected by popular schools, while lower-achieving students entered less popular schools.
 For the government, who wanted to send the most promising students to selective higher education, the decentralized system appeared inefficient.
-[^17] Applicants who failed to enter Schools 1-8 chose either (1) to retake the exam in the following year, (2) to enter local public or private higher educational institutions, or (3) to give up receiving higher education.
+[^16] Applicants who failed to enter Schools 1–8 chose either (1) to retake the exam in the following year, (2) to enter local public or private higher educational institutions, or (3) to give up receiving higher education.
 The Education Minister criticized the decentralized system as follows:
-[Under the decentralized system] many applicants are rejected by Schools 1 and 3 [in Tokyo and Kyoto], which attract a large number of high ability applicants, despite the fact that these applicants have superior academic performance to that of applicants admitted to other rural schools.
+[Under the decentralized system] many applicants are rejected by Schools 1 and 3 [in Tokyo and Kyoto], which attract a large number of high ability applicants, despite the fact that these applicants have superior academic performance to that of applicants admitted to other more rural schools.
 (…) As a result, hundreds of applicants with sufficiently high academic ability to enter rural schools are idly wasting another year [to retake the exam].
-This is not only a pity for them, but also a loss for the country.” (Educational Review No.1146, p.21, published in February 15, 1917)
+This is not only a pity for them, but also a loss for the country.
+[^17] A quote from Educational Review No.1146, p.21, published in February 15, 1917, right before the second introduction of the centralized system.
 To solve this problem, in 1901, the government first asked all schools to unify their entrance exams to a single one, while maintaining decentralized admission decisions.
-In 1902, the government launched a centralized system in which applicants apply for multiple schools, submit their preferences over schools, and take the unified exam.
-Based on their exam scores and preferences, applicants were then assigned to a school (or no school if unsuccessful) by an assignment algorithm announced ex ante.
-Each year, the government announced application procedures in April, three months before the exam in July, as a public notice in Government Gazette.
-We call this system “centralized admissions” or Capp for short.
-The centralized assignment algorithm operates as follows.
-[^18] This is the assignment algorithm used in 1917.
-See Appendix Figure A.3 for a reprint of the original assignment algorithm published in April 1917.
-Strictly speaking, the centralized assignment algorithm was introduced first in 1902, slightly modified in 1903 and 1917, and revised further in 1926.
+In 1902, the government launched a centralized system.
+Under the new system, applicants were allowed to apply for multiple schools by submitting their preference rankings over schools before taking the unified exam.
+Based on their exam scores and school preferences, applicants were then assigned to one school (or no school if unsuccessful) by an assignment algorithm announced ex ante.
+Each year, the government announced application procedures in April, three months before the exam in July, as a public notice in the Government Gazette.
+We call this system centralized admissions (or centralization for short).
+The centralized assignment algorithm was specified as follows.
+[^18] This is the assignment algorithm in 1917 (see Appendix Figure A.3 for a reprint of the original assignment algorithm published in the Government Gazette).
+The centralized assignment algorithm was first introduced in 1902 and reintroduced in 1917 with a slight modification.
   1. (1)
@@
   4. (4)
-    For applicants selected in (1) and for whom the school of his second choice is already filled at the end of (3), then assign each applicant to the school of his third choice or below, repeating the same procedure as (3).
+    For applicants selected in (1) and for whom the school of his second choice is already filled at the end of (3), assign each applicant to the school of his third choice or below, repeating the same procedure as (3).
   5. (5)
@@
 The above algorithm imposes meritocracy up front in which only top-scoring applicants were considered for admission regardless of their preferences (Step (1)).
-Capp is meritocratic in the sense that the test score distribution among assigned students under Capp first order stochastically dominates that under any other mechanisms, including Dapp.
-This step selects only applicants who would be admitted by any school under the Deferred Acceptance algorithm (Abdulkadiroğlu and Sönmez, 2003), one of the most widely used algorithms in today’s college and selective K-12 admissions (Appendix B).
+The centralized system is meritocratic in the sense that the test score distribution among the assigned students first order stochastically dominates that under any other mechanisms, including the original decentralized system.
+Step (1) also selects the same applicants as those who would be admitted by any school under the Deferred Acceptance algorithm, one of the most widely used algorithms in today’s college and selective K-12 admissions (see Appendix B for details).
 These applicants are then assigned to one of Schools 1–8 using the Immediate Acceptance (Boston) algorithm (Steps (2) to (4)).
@@
 To the best of our knowledge, this is the world’s first recorded nation-wide use of any assignment algorithm.
+2.2 Political Economy of the Admission Reforms
 This institutional innovation was short-lived, however.
-Due to political and administrative reasons detailed below, the government switched back to Dapp (with a unified exam) in 1908.
-The government then continued to oscillate between decentralization and centralization, reintroducing Capp in 1917, moving back to Dapp (with a unified exam) in 1919, reinstituting Capp (with major modifications of allowing applicants to list at most two schools) in 1926, and finally settling down to Dapp (with separate exams) in 1928.
+Due to the opposition detailed below, the government switched back to decentralization (with a unified exam) in 1908.
+The government then continued to oscillate between the two regimes, reintroducing centralization in 1917, moving back to decentralization (with a unified exam) in 1919, reinstituting centralization (with modifications of allowing applicants to list at most two schools) in 1926, and finally settling down to decentralization (with separate exams) in 1928.
 In a space of thirty years, there were three periods of centralized admissions: first in 1902–1907, next in 1917–1918, and finally in 1926–1927.191919For details of the evolution of the admission system, see Appendix Table A.1.
 We exploit the series of bidirectional policy changes to examine the impacts of centralization.
-Importantly, it was difficult for applicants to anticipate the exact timing of these policy changes, as the admission policy was decided and announced only a few months in advance.
-This suggests that the reform timing is close to exogenous, as we further explore in Section 3.5.
+Historical documents suggest that the repeated policy changes were a result of annual bargaining between the Ministry of Education (representing the central government) and the Council of School Principals (representing local interests of Schools 1–8).
+[^20] Detailed accounts are provided in Moriguchi (2021), pp.195-199.
+The admission policy, as a result of bargaining, was decided only a few months before the exam, making it difficult for applicants to anticipate the exact timing of the reforms.
+Indeed, we confirm that the timing of the reforms is not associated with the size and composition of applicants, school capacities, and other potential confounding factors (see Section 3.6).
+Throughout the bargaining process, the Ministry of Education insisted on centralized admissions to select the best and brightest, while the Council of School Principals was opposed to centralization, pointing out its adverse impacts on rural schools located in remote regions.
+[^21] Other reasons against centralization included a loss of school autonomy and independence as well as the high administrative cost of implementing centralized admissions.
+They claimed that these rural schools lost the most talented students in their local areas (who would have entered local schools under decentralization) to urban schools, especially School 1.
+They also complained that rural schools instead received many reluctant and unmotivated students who only came to these schools as fallback options.
+Moreover, a review of admission results revealed that the number of applicants from rural areas (areas where rural schools were located) not admitted to any of Schools 1–8 increased considerably under centralization.
+[^22] Ministry of Education (1917), Report on National Higher School Entrance Examination of 1917: Extra Issue.
+This was upsetting to rural schools and their local communities, as rural communities contributed land and other resources to construct rural schools.
+A noted historian summarizes the situation as follows (Takeuchi, 2011, p.121):
+Urban applicants ‘overwhelmed’ rural applicants by applying for rural schools as fallback options.
+Urban applicants robbed rural applicants of opportunities that were once open to them.
+This ruined the meaning of building national higher schools across the nation.
+These observations highlight a potential meritocracy-equity tradeoff of centralization.
+On one hand, the centralized admissions made the school seat allocation more meritocratic, enabling high-achieving students to enter one of Schools 1–8 even if they failed at the most prestigious one.
+On the other hand, this meritocracy might have come at the expense of equal regional access to national higher education, whereby high-achieving urban applicants dominated rural applicants.
+3 Short-run Impacts
 3.1 Predicted Impacts
-To guide our empirical analysis, we first predict effects of centralization by using historical documents and theoretical insights.
-Historical documents suggest that the repeated policy changes resulted from intense bargaining between the Ministry of Education in the central government and the Council of School Principals representing local interests.
-[^20] Detailed accounts are provided in Moriguchi (2021), pp.195-199.
-While the Ministry of Education emphasized the importance of using centralized admissions to enroll the best and brightest to the national higher education, the Council of School Principals was opposed to centralization, pointing out several adverse consequences for rural schools and their local communities.
-First, according to the Council of School Principals, the centralized system encourages applicants to be more risk-taking and rank the most competitive and prestigious urban school as the first choice.
-This is because the centralized system provides applicants with options to apply for lower choice schools as insurance.
-Hypothesis 1. Under the centralized system, more applicants in all areas rank most prestigious School 1 as their first choice.
-Second, Hypothesis 1 means that the centralized system made a greater number of rural applicants apply to and enter urban schools.
-In addition, the centralization allowed high-achieving applicants in urban areas to enter rural schools even if they failed at the most selective urban school.
-This, in turn, often resulted in the rejection of lower-achieving rural applicants who would have been able to enter their local schools under decentralized admissions.
-These effects would jointly increase the mobility of students across the country.
-Hypothesis 2. Under the centralized system, a smaller share of each school’s seats are allocated to its local students.
-The Council of School Principals also complained that the centralized system lets many reluctant and unmotivated students enter rural schools as fallback options, which lowered school morale.
-Third, by the meritocracy constraint embedded in the centralized system, it allocated a higher fraction of seats at Schools 1–8 to high-achieving applicants.
-After reviewing the admission results of 1917 (the first year of the second centralization period), the Council of School Principals found that the number of rural middle-school graduates not admitted to any of Schools 1–8 increased considerably compared to the previous three years of decentralized admissions.
-This is presumably because high-achieving applicants were located disproportionately in urban areas.
-Hypothesis 3. Under the centralized system, a larger share of all school seats in Schools 1–8 are allocated to urban-born applicants.
-Online Appendix B formalizes these hypotheses as theoretical predictions in particular models. A noted historian summarizes these observations as follows (Takeuchi, 2011, p.121):
-Urban applicants ‘overwhelm’ rural applicants by applying for rural schools as fallback options.
-Urban applicants rob rural applicants of opportunities that were once open to them.
-This ruins the meaning of building national higher schools across the nation.
-This upset local communities of rural schools, especially because they contributed land and other resources to the construction of rural schools (Takeuchi, 2011, p.56 and pp.106-107).
-[^21] An additional concern against the centralized system was the administrative cost of implementing it in the absence of modern computers and photocopying technologies.
-Both the Ministry of Education officials and the school principals repeatedly pointed out the difficulty of grading thousands of exam sheets by hand in a short period of time and assigning applicants to schools according to the algorithm (Takeuchi, 2011, pp.118-119; Moriguchi, 2021, p.199).
-The school principals also deemed centralization as an intrusion on their power and autonomy.
-These documents highlight a potential meritocracy-equity tradeoff of centralization.
-On the one hand, the centralized admissions made the school seat allocation more meritocratic, enabling high-achieving students to enter one of the elite schools even if they failed at the most selective one.
-On the other hand, this meritocracy might have come at the expense of equal regional access to higher education, whereby high-achieving urban applicants dominated rural applicants.
-Below we test the above hypotheses as well as their persistence in the long run.
-3 Short-run Impacts
+Motivated by the above historical observations, we first make predictions on the short-run effects of centralization.
+Online Appendix B formalizes the predictions using a simple model (Propositions 1, 2, and 3).
+First, we predict changes in application behavior.
+Specifically, the centralized system would encourage applicants all over the country to rank the most preferred and selective urban school as their first choice.
+This is because the centralized system provides applicants with an option to apply for second and lower choice schools as insurance.
+Second, if high-achieving applicants are disproportionately located in a specific geographical area, such as an urban area, then a larger fraction of all school seats in Schools 1–8 would be assigned to students coming from that area under the centralized system.
+This is because of the meritocratic constraint (Step 1) in the centralized assignment algorithm, which assigns only the top-scoring applicants to any school.
+Finally, a smaller fraction of applicants would be assigned to their local school under centralized admissions.
+Intuitively, meritocratic centralization enables high-scoring applicants from urban areas to enter rural schools even if they fail to enter the most competitive urban school.
+At the same time, the centralized system would make a greater number of high-scoring rural applicants apply to and enter urban schools.
+Under the centralized system, therefore, rural students would more often move to urban areas, while urban students would more often move to rural areas.
+These effects would jointly increase the spatial mobility of students.
 3.2 Data
-To analyze short-run effects of the centralized admission system, we assemble and digitalize data on application and enrollment.
-First, we collect data on the number of applicants by school of their first choice from 1900 to 1930, using Ministry of Education Yearbook and other sources.
-For these two years (1916 and 1917), we also collect data on the number of applicants by school of their first-choice and by prefecture of their middle school.
+To empirically analyze short-run effects of the centralized admission system, we newly assemble data on application and enrollment.
+First, we collect data on the number of applicants by school of their first choice from 1900 to 1930, using the Ministry of Education Yearbook and other sources.
+For two specific years (1916 and 1917), we have additional data on the number of applicants by school of their first-choice and by prefecture of their middle school.
+Japan was divided into 47 prefectures.
 Appendix Section A.3.1 provides detailed explanations for data sources and definitions.
@@
 As an immediate effect, switching back and forth between the centralized and the decentralized admission systems causes stark strategic responses in application behavior.
-Figure 1(a) shows that the three periods of Capp are associated with sharp increases in the share of applicants who chose the most selective School 1 in Tokyo as their first choice, as predicted.
-To observe regional variations, we examine how the propensity of applicants to rank School 1 as their first choice changes between 1916 (under Dapp) and 1917 (under Capp).
-The results in Appendix Table A.3 show that the meritocratic centralization induced applicants in all regions to rank the most prestigious school first and to make more long-distance applications.
-As a result, the competition to enter School 1 became even more intense under Capp (Appendix Figure A.4).
-As a result, under Capp, only a small fraction of first-choice applicants were admitted to School 1, producing many high-scoring applicants who were rejected by School 1 and then assigned to their second or lower-choice school.
+Figure 1(a) shows that the three periods of centralization are associated with sharp increases in the share of applicants who chose the most selective School 1 in Tokyo as their first choice, as predicted.
+To observe regional variations, we examine how the propensity of applicants to rank School 1 as their first choice changes between 1916 (under decentralization) and 1917 (under centralization).
+The results in Appendix Table A.3 show that meritocratic centralization induced applicants in all regions to rank the most prestigious school first and to make more long-distance applications.
+The competition to enter School 1 thus became even more intense under centralization (Appendix Figure A.4).
+As a result, under centralization, only a small fraction of first-choice applicants were admitted to School 1, producing many high-scoring applicants who were rejected by School 1 and then assigned to their second or lower-choice school.
 3.4 Regional Mobility in Enrollment
-The changes in the mechanism and behavior influence enrollment outcomes.
+The changes in the assignment algorithm and application behavior influence enrollment outcomes.
 To measure the geographical mobility of entrants, we compute the “enrollment distance” defined as the direct distance between an entrant’s birth prefecture and the school he entered.
 As Figure 1(b) shows, the centralized system is associated with a sharp and discontinuous increase in enrollment distance.
-[^22] The effect is especially stark in the first two periods of Capp.
-The third period of Capp in 1926–27 was qualitatively different from that in the first and second periods.
-In the third period of Capp, schools were divided into two groups and applicants were allowed to choose and rank at most two schools (one school per group) in 1926–27.
+[^23] The effect is especially stark in the first two periods of centralization.
+The third period of centralization in 1926–27 was qualitatively different from that in the first and second periods.
+In the third period of centralization, schools were divided into two groups and applicants were allowed to choose and rank at most two schools (one school per group) in 1926–27.
 This increase in regional mobility is driven by a sharp reduction in the number of “local entrants” defined as entrants who entered a school in (or near) their birth prefecture.
 To show this, we estimate the following regression for each school $s$ separately:
-$\begin{split}Y_{pt}^{s}=&\ \beta_{1}^{s}\times Centralized_{t}\times 1\{\text{school }s\text{ is located in prefecture }p\}\\ &+\beta_{2}^{s}\times Centralized_{t}\times 1\{\text{school }s\text{ is 1-100 km away from prefecture }p\}\\ &+\alpha^{s}X_{pt}+\gamma_{t}^{s}+\gamma_{p}^{s}+\epsilon_{pt}^{s},\end{split}$ (1)
+$\begin{split}Y_{pt}^{s}=&\ \beta_{1}^{s}\times Centralized_{t}\times 1\{\text{school }s\text{ is located in prefecture }p\}\\ &+\beta_{2}^{s}\times Centralized_{t}\times 1\{\text{school }s\text{ is 1--100km away from prefecture }p\}\\ &+\alpha^{s}X_{pt}+\gamma_{t}^{s}+\gamma_{p}^{s}+\epsilon_{pt}^{s},\end{split}$ (1)
 […]
-As shown in Panel (a) of Table 1, Capp reduces the number of local entrants born in the school’s prefecture across the country.
+As shown in Panel (a) of Table 1, centralization reduces the number of local entrants born in the school’s prefecture across the country.
 The coefficients $\beta_{1}^{s}$ are significantly negative for all schools.
-Column (1) shows that the number of School 1 entrants born in Tokyo prefecture declined by about 25%.
-Schools 4–7 experienced reductions in the number of entrants born not only from the school’s prefecture but also from surrounding prefectures.
-In other words, centralization weakened the local monopoly power of each school by creating a national market for higher education, consistent with our theoretical prediction in Appendix B (Proposition 3).
-3.4 Meritocracy vs Equal Regional Access
-Which regions gained more school seats under the centralized system?
-We now investigate how centralization altered the geographical allocation of school seats.
-Figure 2(a) plots the change in the total number of entrants to Schools 1–8 from Dapp to Capp by their birth prefecture, where the darker blue color indicates the greater decline and the darker red color indicates the greater increase.
-The figure shows that Tokyo prefecture (where School 1 was located) and its surrounding area gained school seats under Capp, while most of western and northern prefectures lost school seats.
-To see this quantitatively, we regress the number of entrants to Schools 1–8 in year $t$ born in prefecture $p$ on the interaction terms between $Centralized_{t}$ and each of the following two indicator variables: the indicator that takes 1 if prefecture $p$ was Tokyo and the indicator that takes 1 if prefecture $p$ was 1-100 km away from Tokyo.
-To control for the decline of local entrants, we also add the interaction terms between $Centralized_{t}$ and each of the following two indicator variables: the indicator that takes 1 if any of Schools 1–8 was located in prefecture $p$ and the indicator that takes 1 if any of Schools 1–8 was located 1-100 km away from prefecture $p$.
+Column (1) shows that the number of School 1 entrants born in Tokyo prefecture declined by about 26% relative to its mean under decentralization.
+Schools 2–7 experienced reductions in the number of entrants born not only from the school’s prefecture but also from surrounding prefectures.
+In other words, centralization weakened the local monopoly power of each school by integrating local markets into a national market, consistent with our prediction.
+3.5 Meritocracy vs Equal Access
+We now investigate the distributional effect of the centralized system on the allocation of school seats at Schools 1–8.
+Figure 2(a) plots the change in the number of entrants to Schools 1–8 from decentralization to centralization by their birth prefecture, where the darker blue color indicates the greater decline and the darker red color indicates the greater increase.
+The figure shows that Tokyo prefecture (where School 1 was located) and its surrounding area gained school seats under centralization, while most of western and northern prefectures lost school seats.
+To see this quantitatively, we regress the number of entrants to Schools 1–8 in year $t$ born in prefecture $p$ on the interaction terms between $Centralized_{t}$ and each of the following two indicator variables: the indicator that takes 1 if prefecture $p$ was Tokyo and the indicator that takes 1 if prefecture $p$ was 1–100 km away from Tokyo.
+To control for the decline of local entrants, we also add the interaction terms between $Centralized_{t}$ and each of the following two indicator variables: the indicator that takes 1 if any of Schools 1–8 was located in prefecture $p$ and the indicator that takes 1 if any of Schools 1–8 was located 1–100 km away from prefecture $p$.
 As in Equation (1), we also control for observable characteristics of prefecture $p$ and year $t$ and include year and prefecture fixed effects.
-Column (1) in Table 1 Panel (b) shows the number of entrants born in Tokyo prefecture increased by 13% per year under Capp relative to its mean under Dapp.
-The increase in the number of entrants is even higher for its surrounding area; the number of entrants born in prefectures 1-100 km away from Tokyo increased by 45% per prefecture per year under Capp relative to its mean under Dapp.
-[^23] One possible reason for this concentration of the effect in surrounding prefectures is that applicants from this area might be marginal applicants who would not be admitted under Dapp but would be admitted under Capp.
-On the other hand, many high-scoring applicants from the Tokyo prefecture might be admitted even under Dapp.
-Consistent with this hypothesis, the number of entrants from Tokyo (201) is much larger than the number of entrants from surrounding prefectures (26 per prefecture) under Dapp.
-Motivated by these observations, we define “Tokyo area” as the set of prefectures located within 100 km from Tokyo (see Appendix Table A.3 for its location).
+Column (1) in Table 1 Panel (b) shows the number of entrants born in Tokyo prefecture increased by 12% per year under centralization relative to its mean under decentralization.
+The increase in the number of entrants is even higher for its surrounding area; the number of entrants born in prefectures 1–100 km away from Tokyo increased by 45% per year under centralization relative to its mean under decentralization.
+[^24] One possible reason for this concentration of the effect in surrounding prefectures is that applicants from this area might be marginal applicants who would not be admitted under decentralization but would be admitted under centralization.
+By contrast, many high-scoring applicants from the Tokyo prefecture might be admitted even under decentralization.
+Consistent with this hypothesis, the number of entrants from Tokyo (201) is much larger than the number of entrants from surrounding prefectures (26 per prefecture) under decentralization.
+Motivated by these observations, we define the “Tokyo area” as the set of prefectures located within 100 km from Tokyo (see Appendix Table A.3 for its location).
 Figure 2(b) depicts the time evolution of the share of entrants to Schools 1–8 who were born in the Tokyo area.
-Consistent with the above results, the share of Tokyo-area born entrants rose during the periods of Capp.
-To see which schools received more Tokyo-area born entrants under Capp, we regress the number of entrants to each school $s$ in year $t$ born in prefecture $p$ on the interaction terms between $Centralized_{t}$ and the variables indicating whether entrants’ birth prefecture $p$ is Tokyo or near Tokyo.
+Consistent with the above results, the share of Tokyo-area-born entrants rose during the periods of centralization.
+To see which schools received more Tokyo-area-born entrants under centralization, we regress the number of entrants to each school $s$ in year $t$ born in prefecture $p$ on the interaction terms between $Centralized_{t}$ and the variables indicating whether entrants’ birth prefecture $p$ is Tokyo or near Tokyo.
 We include the same variables used in Equation (1) to control for school locations and other prefecture and year characteristics.
-Under Capp, a greater number of Tokyo-area-born students entered less selective rural schools, as shown in Columns (2)–(9) in Table 1 Panel (b).
+Under centralization, more Tokyo-area-born students entered less selective rural schools, as shown in Columns (2)–(8) in Table 1 Panel (b).
 [^25] The results remain almost the same when we additionally control for observable prefecture characteristics (see Appendix Table A.4).
-In summary, the net effect of Capp is such that the increased inter-regional applications caused high-achieving students in the Tokyo area to crowd out lower-achieving, rural-born students from their local schools.
-This evidence is consistent with the historian’s observation (Takeuchi (2011) cited in Section 2.2) that, as a result of the centralization, urban applicants mainly from Tokyo dominated rural applicants in entering national elite schools.
-The Tokyo area shares common characteristics with what is generally considered urban.
+In summary, the net effect of centralization is such that the increased inter-regional applications caused high-achieving, urban-born students to crowd out lower-achieving, rural-born students from their local schools.
+The Tokyo area shares common characteristics with what is generally considered urban and with high socioeconomic status.
 Prefectures in the Tokyo area had a larger population, higher income per capita, and better educational infrastructure (indicated by a greater number of middle-school graduates eligible for higher education) compared to other prefectures (Appendix Table A.5).
 In Appendix Table A.6, we use these urban characteristics instead of the Tokyo area indicator in the same specification as Column (1) in Panel (b) of Table 1.
-The result shows that these urban characteristics, especially educational infrastructure, positively predict the areas which produced a greater number of entrants under Capp.
-Therefore, a greater number of high achieving students in the Tokyo area may explain why the Tokyo area gained more school seats under centralization.
-Students’ heterogenous preferences across areas can be another reason why the Tokyo area gained more school seats under the centralized system.
+The result shows that these urban characteristics, especially educational infrastructure, positively predict the areas which produced a greater number of entrants under centralization.
+Therefore, richer educational infrastructures and a greater number of high achieving students in the Tokyo area may explain why the Tokyo area gained more school seats under centralization.
+Students’ heterogeneous preferences across school locations can be another reason why the Tokyo area gained more school seats under the centralized system.
 Students prefer their local school for geographical and cultural proximity.
@@
 Holding academic achievement constant, during the decentralization, students born in the Tokyo area are more likely to apply to School 1 and therefore fail to enter any school (even though they might have been able to enter Schools 2–8 had they applied).
-This hypothesis is consistent with Appendix Table A.6, which shows that the share of applicants to School 1 during the decentralization is associated with a greater number of entrants to Schools 1–8 during the centralization.
-In summary, our analysis suggests that a combination of better educational infrastructure and strong preference for School 1 explains why the Tokyo area gained more school seats under the centralized system.
+This hypothesis is consistent with Appendix Table A.6, which shows that the share of applicants to School 1 during decentralization is associated with a larger increase in the number of entrants to Schools 1–8 during centralization.
+In summary, the net effect of centralization is such that the increased regional mobility caused urban-born high achievers with high socioeconomic statuses to crowd out rural students from elite higher education.
+Key mechanisms behind this distributional effect are better educational infrastructures and strong preference for School 1.
 3.6 Other Institutional Changes
-We discuss potential threats to our empirical analysis.
+We finish the short-run analysis by discussing potential threats.
 In particular, if changes in other institutional and policy factors were correlated with the admission reforms, it could influence application behavior and enrollment outcomes, explaining our finding that more applicants applied for School 1 during centralization.
 Such other institutional factors include simultaneous reforms in middle schools, the total number of applicants, capacities at Schools 1–8, and the capacity of School 1 relative to the capacities of other schools.
-We investigate these concerns and confirm that centralization periods are not correlated with time-series changes in the following variables (Columns (1)–(6) in Appendix Table A.7): the number of middle school graduates, the national number of applicants as well as the level of competitiveness (measured by the number of entrants divided by the number of applicants), the total number of entrants to Schools 1–8, the share of entrants to School 1 in all entrants, the probability of unsuccessful applicants retaking the exam in subsequent years, and the average age of entrants.
+We investigate these concerns and confirm that centralization periods are not correlated with time-series changes in the following variables (Columns (1)–(6) in Appendix Table A.7): the number of middle school graduates, the national number of applicants as well as the level of competitiveness (measured by the number of entrants divided by the number of applicants), the total number of entrants to Schools 1–8, the share of entrants to School 1 in all entrants, the probability of unsuccessful applicants retaking the exam in subsequent years, the average age of entrants, and government expenditure on higher education.
 A potential concern with the above time-series analysis is that the insignificant results in Appendix Table A.7 may be due to a small sample size (around 30).
@@
 As such, the estimated impacts of centralization may be confounded by the unification of entrance exams and more flexible exam location choices.
-To investigate this issue, we analyze how key outcomes change from 1900 to 1901, during which the government also introduced a single entrance exam that applicants were allowed to take anywhere while the assignment method remained unchanged (i.e., decentralized).
+To investigate this issue, we analyze how key outcomes change from 1900 to 1901, during which the government also introduced a single entrance exam that applicants were allowed to take at any school while the assignment method remained decentralized.
 Figures 1(a) and 1(b) show that this institutional change from 1900 to 1901 induced little changes in application and enrollment patterns.
@@
 4 Long-run Impacts
-To assess long-run effects of the meritocratic centralization, we provide two sets of empirical analysis.
+To assess long-run effects of meritocratic centralization, we provide two sets of empirical analysis.
 First, we investigate the persistence of the distributional effect of centralization that we document in the short-run analysis.
 We do so by a difference-in-differences strategy comparing labor-market outcomes of urban- and rural-born individuals across birth cohorts with differential exposure to the centralized admissions.
-Our analysis shows that the centralization increased both (a) the number of career elites born in urban areas and (b) the number of career elites living in urban areas as adults.
-The admission reforms thus influence the regional origins and destinations of career elites.
-Second, we explore the national productive efficiency of the two admission systems.
-We find that cohorts more intensely exposed to the centralized admissions produced a larger national number of civil officials who were promoted to top ranks.
-This finding indicates that the centralization boosted the national production of career elites, at least for the specific important category of top-ranking government officials.
-4.1 Regional Distribution of Career Elites
+Our analysis shows that centralization increased both (a) the number of occupational elites born in urban areas and (b) the number of occupational elites living in urban areas as adults.
+The admission reforms thus influence the regional origins and destinations of elites.
+Second, we compare the national production of elites between the two admission systems.
+We find that cohorts more intensely exposed to the centralized admissions produced a larger national number of civil officials who were promoted to top ranks, as well as other types of occupational elites (such as top income earners).
+This finding indicates that centralization boosted the national production of individuals reaching leadership positions.
+4.1 Regional Distribution of Elites
 Personnel Inquiry Records Data
-To analyze long-run effects of the centralization on students’ career outcomes, we use the Personnel Inquiry Records (PIR) published in 1939.252525Japan went into the military regime with the promulgation of the National General Mobility Act in 1938.
-Although the military government began massive social interventions, their effect was relatively small in 1939 (Moriguchi and Saez (2008)).
-In addition, most of the information in PIR (1939) is based on 1938 data.
-The PIR is an equivalent of Who’s Who, which compiles a selective list of socially distinguished individuals (including high-income earners, top business managers, elite professionals, and high-ranking public servants) and provides their biographical information.
-In total, the 1939 PIR lists 55,742 individuals or 0.15% of the adult Japanese population of that time.
-We detail the selection process in Appendix Section A.4.1.
-To capture the effects of the first period of the centralized admission system in 1902–1907, we use the cohorts born in 1880–1894, who turned 17 years old (the age eligible for application) in 1897–1911.
-The cohorts born in 1880–1894 were 45 to 59 years old in 1939.
-The number of individuals listed in the PIR in each of these cohorts is about 1,800.
-We use the following information from the PIR data for each individual: full name, birth year, birth prefecture, prefecture of residence, final education, occupation titles and organization names, the medal for merit and the court rank awarded (if any), and the amount of national income tax and business tax paid (if any).
-We define the following groups of elites as a subset of PIR-listed individuals: (1) the top 0.01% and 0.05% income earners according to the national income distribution, (2) prestigious medal recipients (civilians who received either the medal of the fifth order of merit or above, or the court rank of the junior fifth rank or above), (3) corporate executives (individuals who hold an executive position in a corporation and pay a positive amount of income or business tax), (4) top politicians and bureaucrats (individuals whose occupation is either Imperial Diet member, Cabinet member, or high-ranking central government official), and (5) Imperial University professors or associate professors.
+To analyze the long-run effects of centralization on students’ career outcomes, we use the Personnel Inquiry Records (PIR) published in 1939.262626Japan went into the military regime after the National General Mobility Act of 1938.
+Although the military government began massive economic interventions, their effect was relatively small in 1939 (Moriguchi and Saez, 2008).
+Moreover, most of the information in the 1939 PIR was based on 1938 data.
+The PIR is an equivalent of Who’s Who, which compiles a highly selective list of socially distinguished individuals—encompassing economic, political, cultural, and aristocratic elites—and provides their biographical information.
+The 1939 PIR lists 55,742 individuals or 0.15% of the adult Japanese population of that time.
+[^27] We discuss the PIR’s selection criteria in Appendix Section A.4.1.
+According to their occupational distribution in Appendix Table A.8 Column (1), 66% hold managerial positions, 7.5% are politicians or bureaucrats, and the rest consists mainly of professionals such as scholars and engineers.
+Throughout this section, we refer to the individuals listed in the PIR as “elites.”
+To capture the effects of the first period of the centralized admissions in 1902–1907, we use the cohorts born in 1880–1894, who turned 17 years old (the age eligible for application) in 1897–1911.
+These cohorts were 45 to 59 years old in 1939.
+The number of elites in each of these cohorts is about 1,800.
+We use the following information from the PIR for each individual: full name, birth year, birth prefecture, prefecture of residence, final education, occupation titles and organization names, the medal for merit and the court rank awarded (if any), and the amount of national income tax and business tax paid (if any).
+Among all elites in our sample, 23% graduated from Imperial Universities.
+We define the following subgroups of elites (as a subset of PIR-listed individuals): (1) the top 0.01% and 0.05% income earners according to the national income distribution, (2) prestigious medal recipients (civilians who received high-ranking imperial medals for their exceptional service or merit in various fields), [^28] These prestigious medal recipients held a broad range of occupations, including scholars, managers, engineers, and bureaucrats as shown in Appendix Table A.8 Column (4).
+(3) corporate executives (individuals who hold an executive position in a corporation and pay a positive amount of income or business tax), (4) top politicians and bureaucrats (individuals whose occupation is either Imperial Diet member, Cabinet member, or high-ranking central government official), (5) Imperial University professors or associate professors, (6) all occupational elites (all individuals listed in the PIR except for hereditary elites defined as individuals whose sole occupation is either peerage or landlord).
 Appendix Section A.4.1 and Table A.9 provide detailed definitions and descriptions of each group.
-These categories encompass economic, social, cultural, political, and academic definitions of career elites.
-We use this data to count the number of elites in each group defined by birth prefecture and birth cohort.
+These categories encompass economic, political, and cultural definitions of occupational elites.
+We use this data to count the number of elites in each subgroup defined by birth prefecture and birth cohort.
 These counts allow us to conduct a difference-in-differences analysis that compares long-term career outcomes of urban- and rural-born individuals by each cohort’s exposure to the centralized admission system.
@@
 This series of findings suggests that possible sample selection bias in the PIR data is unlikely to drive our empirical results.
-Urban-Rural Disparity in Producing Career Elites
+Urban-Rural Disparity in Producing Elites
 We estimate the long-run impacts of the centralized admission system by conducting a difference-in-differences analysis by birth cohorts and birth areas.
-The key idea behind our empirical strategy is that applicants born in the Tokyo area should experience a greater gain in entering Schools 1–8 under Capp relative to Dapp, since the centralized system is designed to be more meritocratic and high-achieving students are disproportionately located in urban areas.
-This expectation is confirmed in Figure 2 and Table 1.
-We exploit this differential gain in school access to compare the career outcomes of individuals born inside and outside the Tokyo area by the cohort’s exposure to Capp.
-If admission to Schools 1–8 increases one’s chance of becoming a career elite, we should observe a greater number of elites born inside the Tokyo area for the cohorts exposed to Capp.
+The key idea behind our empirical strategy is that urban-born applicants experienced a greater gain in entering Schools 1–8 under centralization relative to decentralization, as shown in the short-run analysis in Figure 2 and Table 1.
+We exploit this differential gain in school access to compare the career outcomes of individuals born inside and outside the Tokyo area by the cohort’s exposure to centralization.
+If admission to Schools 1–8 increases one’s chance of becoming an elite, we should observe more elites born inside the Tokyo area for the cohorts exposed to centralization.
 We estimate a difference-in-differences specification as follows:
@@
 where $Y_{pt}$ is the number of elites born in cohort $t$ and prefecture $p$.
-$Centralized_{t}$ is a measure of cohort $t$’s exposure to Capp, which is, in the baseline specification, a binary indicator that cohort $t$ turned 17 during Capp (1902–1907).
+$Centralized_{t}$ is a measure of cohort $t$’s exposure to centralization, which is, in the baseline specification, a binary indicator that cohort $t$ turned 17 during centralization (1902–1907).
 $Tokyo\_area_{p}$ is the indicator that takes 1 if prefecture $p$ is in the Tokyo area.
@@
 In addition, we report the results of clustering at cohort level, which are estimated by wild cluster bootstrap (Cameron and Miller, 2015; Roodman et al., 2019) due to the small number of clusters (15 cohorts).
-The above regression defines $Centralized_{t}$ to be a binary indicator, as the simplest proxy for the intensity of exposure to Capp.
+The above regression defines $Centralized_{t}$ to be a binary indicator as the simplest proxy for the intensity of exposure to centralization.
 In reality, however, a nontrivial number of unsuccessful applicants retook the exam at age 18 and beyond.
-As a result, the cohorts who turned age 17 in 1899–1901 were partially and increasingly exposed to Capp (as some of them took the exam in 1902).
-The cohorts who turned age 17 in 1902–1904 were fully exposed to Capp.
-The cohorts who turned age 17 in 1905–1907 were partially and decreasingly exposed to Capp (as some of them took the exam in 1908).
-Finally, the intensity of exposure drops to zero for the cohorts who turned age 17 in 1908.272727The intensity of exposure to Capp is estimated to be 0.00, 0.02, 0.09, 0.36, 1.00, 0.98, 0.91, 0.64, and 0.00 for the cohort who turned age 17 in 1896–98, 1899, 1900, 1901, 1902–04, 1905, 1906, 1907, and 1908–11, respectively.
+As a result, the cohorts who turned age 17 in 1899–1901 were partially and increasingly exposed to centralization (as some of them took the exam in 1902).
+The cohorts who turned age 17 in 1902–1904 were fully exposed to centralization.
+The cohorts who turned age 17 in 1905–1907 were partially and decreasingly exposed to centralization (as some of them took the exam in 1908).
+Finally, the intensity of exposure drops to zero for the cohorts who turned age 17 in 1908.303030The intensity of exposure to centralization is estimated to be 0.00, 0.02, 0.09, 0.36, 1.00, 0.98, 0.91, 0.64, and 0.00 for the cohort who turned age 17 in 1896–98, 1899, 1900, 1901, 1902–04, 1905, 1906, 1907, and 1908–11, respectively.
 See Appendix Section A.4.1 for data and methods.
-In the following visual analysis, we explicitly incorporate cohort’s intensity of exposure to Capp by coloring cohorts according to their intensity of exposure to Capp.
-We first check whether the number of Imperial University graduates born inside the Tokyo area increased for the cohorts exposed to Capp.
+In the following visual analysis, we explicitly incorporate cohort’s intensity of exposure to centralization by coloring cohorts according to their intensity of exposure to centralization.
+We first check whether the number of Imperial University graduates born inside the Tokyo area increased for the cohorts exposed to centralization.
 Since all Schools 1–8 graduates were automatically admitted to an Imperial University during this period, the areas that produced more Schools 1–8 entrants should produce more Imperial University graduates.
 Figure 3 confirms this expectation.
-Panel (a) compares the average number of Imperial University graduates who were born in prefectures inside and outside the Tokyo area by cohorts (represented by their birth year plus 17 on the horizontal axis).
-Figure 3 (b) shows that the urban-rural difference in the number of Imperial University graduates rises as the intensity of exposure to Capp increases and falls sharply when Capp ends in 1908.
+Panel (a1) compares the average number of Imperial University graduates who were born inside and outside the Tokyo area by cohorts (represented by their birth year plus 17 on the horizontal axis).
+Panel (a2) shows that the urban-rural difference in the number of Imperial University graduates rises as the intensity of exposure to centralization increases and falls sharply when centralization ends in 1908.
 Column (1) in Table 2 shows that the estimate of $\beta$ in the above regression is positive and statistically significant.
-Our main results are presented in Figure 3 (c)–(f) and Table 2 Columns (2)–(7).
-Figure 3 (c)–(f) show difference-in-differences plots that compare the number of elites (the top 0.05% income earners and medal recipients) who were born inside and outside the Tokyo area by the cohort’s exposure to Capp.
-For both elite categories, the plots show that the difference between the Tokyo area and the rest grows as the intensity of exposure to Capp increases, and then drops sharply when Capp ends.
+Our main results are presented in Figure 3 Panels (b)–(d) and Table 2 Columns (2)–(8).
+Figure 3 Panels (b1)–(d1) show difference-in-differences plots that compare the number of elites (the top 0.05% income earners, medal recipients, and all occupational elites) who were born inside and outside the Tokyo area by the cohort’s exposure to centralization.
+For these elite categories, Panels (b2)–(d2) show that the difference between the Tokyo area and the rest grows as the intensity of exposure to centralization increased, and then drops sharply when centralization ended.
 Appendix Figure A.7 shows qualitatively similar difference-in-differences plots for the top 0.01% income earners, corporate executives, top politicians and bureaucrats, and Imperial University professors.
-Table 2 Columns (2)–(7) show that the long-run effects of Capp are economically and statistically significant.
-For the cohorts exposed to Capp, the number of career elites born inside the Tokyo area (compared to those born outside the Tokyo area) increases by 34% for the top 0.01% income earners, 24% for the top 0.05% income earners, 36% for medal recipients, 14% for corporate executives, 50% for top politicians and bureaucrats, and 47% for Imperial University professors (in Panel B).
-[^28] In Appendix Table A.10, we also examine other groups of elite professionals and obtain similar results for scholars, engineers, and physicians.
+Table 2 Columns (2)–(7) show that the long-run effects of centralization are economically and statistically significant.
+For the cohorts exposed to centralization, the number of elites born inside the Tokyo area (compared to those born outside the Tokyo area) increases by 34% for the top 0.01% income earners, 24% for the top 0.05% income earners, 36% for medal recipients, 13% for corporate executives, 50% for top politicians and bureaucrats, 38% for Imperial University professors, and 14% for all occupational elites (in Panel B).
+[^31] In Appendix Table A.11, we also examine other types of occupational elites and obtain similar results for scholars, engineers, and physicians.
 In Panel A, we control only for cohort and prefecture fixed effects.
@@
 The coefficients fall slightly in magnitude after adding control variables, but remain sizable.
-Table 2 Panel C shows that the effects are symmetric with respect to the direction of the admission reforms, i.e., the change from Dapp to Capp and the change from Capp to Dapp produce similar effects of the opposite sign.
-In Table 2 Panel D, we replace the centralization dummy by the cohort’s intensity of exposure to Capp.
+Table 2 Panel C shows that the effects are symmetric with respect to the direction of the admission reforms, i.e., the change from decentralization to centralization and the change from centralization to decentralization produce similar effects of the opposite sign.
+In Table 2 Panel D, we replace the centralization dummy by the cohort’s intensity of exposure to centralization.
 The results remain qualitatively the same as the baseline results.
-These results suggest that almost four decades after its implementation, Capp had lasting effects on the career trajectories of students.
+These results suggest that almost four decades after its implementation, centralization had lasting effects on the career trajectories of students.
 The above results are robust to alternative specifications.
-First, the analysis in Table 2 Panel D assumes that the cohort’s intensity to exposure to Capp is exogenous and the same across years, which may be a strong assumption.
-However, even when we drop the cohorts who are exposed to both Capp and Dapp (i.e., cohorts who became age 17 in 1901 and 1907) from the sample, we still find qualitatively the same results (Appendix Table A.11).
+First, the analysis in Table 2 Panel D assumes that the cohort’s intensity to exposure to centralization is exogenous and the same across years, which may be a strong assumption.
+However, even when we drop the cohorts who are exposed to both centralization and decentralization (i.e., cohorts who became age 17 in 1901 and 1907) from the sample, we still find qualitatively the same results (Appendix Table A.12).
 Second, we test if the assumption of parallel pre-event trends holds.
@@
 Another potential threat to our identification strategy is that there may be some age-specific trends in the number of elites that covary with the cohort-region variation we use.
-Specifically, the number of elites selected in the 1939 long-term data peaks at around the cohort who were 51 years old in 1939 (corresponding to the cohort who turned age 17 in 1905) and gradually falls for younger and older cohorts, suggesting that there are certain ages at which individuals are more likely to be listed in the long-term data.
+Specifically, the number of elites listed in the 1939 PIR data peaks at around the cohort who were 51 years old in 1939 (corresponding to the cohort who turned age 17 in 1905) and gradually falls for younger and older cohorts.
+This suggests that there are certain ages at which individuals are more likely to be listed in the long-term data.
 Such age effects may generate different trends in the number of elites born in the Tokyo and other areas, due, for example, to differences in population size across these areas.
 To address this concern, we use an earlier edition of the PIR published in 1934, construct the prefecture-cohort level data for the same cohorts used in our main analysis (but observed 5 years earlier), and conduct similar regression analyses.
-The results in Appendix Table A.13 confirm that our key results remain qualitatively the same even when we use the 1934 PIR data.
+The results in Appendix Table A.14 confirm that our key results remain qualitatively the same even when we use 1934 PIR data.
 Finally, we conduct placebo tests to examine if the results are driven by other factors such as the sample selection of the PIR or changes in cohort populations.
-The urban-rural difference in the cohort’s birth populations does not change significantly with the cohort’s exposure to Capp, as shown in Appendix Table A.14.
+The urban-rural difference in the cohort’s birth populations does not change significantly with the cohort’s exposure to centralization, as shown in Appendix Table A.15.
 As an additional placebo test, we also look at unrelated career outcomes.
-Among the distinguished individuals listed in the PIR, we expect that landlords (defined in Appendix Section A.4.1) are least likely to be affected by the introduction of Capp as receiving higher education was not a typical pathway to become a landlord.
-As shown in Appendix Table A.14, the estimated effect of Capp on the number of landlords is small and statistically insignificant.
-These findings further support our conclusion that the meritocratic centralization enlarged the urban-rural gap in the capacity to produce future career elites.
-Geographical Destinations of Career Elites
-Having established that Capp affects the geographic origins of elites, we now ask how it affects their geographic destinations.
-While the former is about regional inequality in educational opportunities, the latter is about regional inequality in the supply of highly skilled human capital, which potentially affects economic growth and inequality.
+Among the distinguished individuals listed in the PIR, we expect that landlords (defined in Appendix Section A.4.1) are least likely to be affected by the introduction of centralization as receiving higher education was not a typical pathway to become a landlord.
+As shown in Appendix Table A.15, the estimated effect of centralization on the number of landlords is small and statistically insignificant.
+These findings further support our conclusion that meritocratic centralization enlarged the urban-rural gap in the capacity to produce future individuals in leadership positions.
+Geographical Destinations of Elites
+Having established that centralization affects the geographic origins of occupational elites, we now ask how it affects their geographic destinations.
+While the former is about regional inequality in educational opportunities, the latter is about regional inequality in the supply of highly skilled human capital.
 We first examine if the centralization-induced increase in inter-regional mobility in the short run boosted the geographical mobility of elites in the long run.
-Surprisingly, it did not: The urban-rural difference in the fraction of elites whose prefectures of residence differ from their birth prefectures did not increase under Capp, as shown in Appendix Table A.14.
+It did not: The urban-rural difference in the fraction of elites whose prefectures of residence differ from their birth prefectures did not increase under centralization, as shown in Appendix Table A.15.
 We find similar results when we use the distance between an elite’s birth prefecture and his prefecture of residence as an alternative measure of long-run mobility.
-This result suggests that, even though a greater number of students born in the Tokyo area entered rural schools under Capp, most of them returned to the Tokyo area when pursuing their careers
-If the greater number of Tokyo-area born students admitted to rural schools under Capp returned to the Tokyo area eventually for their subsequent careers, we should observe a greater number of elites living in the Tokyo area for the cohorts exposed to Capp.
+This result suggests that, even though a greater number of students born in the Tokyo area entered rural schools under centralization, most of them returned to the Tokyo area when pursuing their careers.
+If the increased number of Tokyo-area-born students admitted to rural schools under centralization returned to the Tokyo area eventually for their careers, we should observe a greater number of elites living in the Tokyo area for the cohorts exposed to centralization.
 To test this hypothesis, we redefine the outcome variables by changing the prefecture ($p$) from birth prefecture to prefecture of residence and estimate the same equation.
-Table 3 shows large positive effects of Capp on the urban-rural gap in the number of elite residents.
-For the cohorts exposed to Capp relative to Dapp, the number of elites living in the Tokyo area in their middle age (compared to those living outside the Tokyo area) increases by 28% for the top 0.01% income earners, 23% for the top 0.05% income earners, 29% for medal recipients, 22% for corporate executives, 20% for top politicians and bureaucrats, and 50% for Imperial University professors (Panel B).
-These results suggest that the meritocratic centralization intensified the concentration of career elites in urban areas in the long run.
-4.2 National Production of Career Elites
+Table 3 shows large positive effects of centralization on the urban-rural gap in the number of elite residents.
+For the cohorts exposed to centralization relative to decentralization, the number of elites living in the Tokyo area in their middle age (compared to those living outside the Tokyo area) increases by 30% for the top 0.01% income earners, 25% for the top 0.05% income earners, 30% for medal recipients, 23% for corporate executives, 19% for top politicians and bureaucrats, 41% for Imperial University professors, and 20% for all occupational elites (Panel B).
+These results suggest that meritocratic centralization intensified the concentration of elites in urban areas in the long run.
+4.2 National Production of Elites
 The above analysis examines the distributional consequence of the centralized admissions.
 A complementary problem is its overall impact for the whole country.
-To study it, we now turn to whether the meritocratic centralization improved the national production of career elites in the long run.
-We focus on a specific group of elites — higher civil officials — for whom we have complete-count data from administrative records.
-Top-ranking civil officials were also the most relevant group of career elites; in our empirical context, higher civil service was considered to be one of the most prestigious jobs due to their political and legislative influence.
-As a result, high-achieving students at the top universities competed to get jobs at the Ministry of Finance and other selective ministries (Shimizu, 2019).
-We investigate whether cohorts exposed to Capp produced a greater national number of top-ranking civil officials compared to cohorts exposed to Dapp.
+To study it, we now turn to whether meritocratic centralization increased the national production of occupational elites in the long run.
+We first focus on a specific subgroup of elites—higher civil officials—for whom we have complete-count data from administrative records.
+Top-ranking civil officials are also a key group of occupational elites; in our empirical context, higher civil service was considered to be one of the most prestigious jobs due to their political and legislative influence.
+As a result, high-achieving students at the top universities competed to enter the Ministry of Finance and other selective ministries (Shimizu, 2019).
+We investigate whether cohorts exposed to the centralized admissions produced a greater number of top-ranking civil officials compared to cohorts exposed to the decentralized admissions.
+After that, we also analyze the impact of centralization on the national production of all occupational elites listed in the PIR.
 Higher Civil Officials Data
@@
 The 1893 ordinance required all individuals to pass the HCSE for appointment in the administrative division of higher civil service with some exceptions for special appointments (Spaulding, 1967, Chapter 25; Shimizu, 2019, Chapter 5).
-We digitized the information of all individuals who passed the administrative division of the HCSE in 1894–1941, including their full name, education, year of university graduation, year of passing the exam, starting position, final position, year of retirement, and other notable positions held.
-Because education includes not only final but also the second to final education, we observe both university and higher school (if applicable).
-In the following analysis, we define “top-ranking officials” as higher civil officials who were internally promoted to one of the top three ranks by the end of their career.
+We digitized the information of all individuals who passed the administrative division of the HCSE in 1894–1941, including their full name, education (both university and higher school), year of university graduation, year of passing the exam, starting position, final position, year of retirement, and other notable positions held.
+We define “top-ranking officials” as higher civil officials who were internally promoted to one of the top three ranks by the end of their career.
 In the prewar Japanese bureaucracy system, the higher civil service refers to the top ten ranks of national government offices in the administrative, judicial, and diplomatic divisions.
-Within the higher civil service, the top three ranks were distinctively called “imperial appointees.” The first rank consisted of minister level positions, while the second and third ranks consisted of vice minister level positions such as vice minister, director general, bureau chief, and prefectural governor.
+Within the higher civil service, the top three ranks were distinctively called “imperial appointees.” The first rank consisted of minister level positions, while the second and third ranks consisted of vice minister level positions (such as vice minister, director general, bureau chief, and prefectural governor).
 Appendix A.4.2 provides more details.
-We count the number of top-ranking officials and the number of individuals who passed the administrative division of the HCSE (hereafter “exam passers”) produced by cohort.
-Here cohort is defined by “the year of entering a higher school or its equivalent.” Since we only observe the year of university graduation in the above data, we estimate “the year of entering a higher school or its equivalent” using the method described in Appendix A.4.2.
-We first find the exact year of entering a higher school for all top-ranking officials who graduated from Schools 1–8 by searching their names in Student Registers of Schools 1–8.
-Using this information, we then estimate “the year of entering a higher school or its equivalent” for the rest of higher civil officials.
+We count the number of top-ranking officials and the number of individuals who passed the administrative division of the HCSE (hereafter “exam passers”) by cohort.
+Here cohort is defined by “the year of entering a higher school or its equivalent.” Since we only observe the year of university graduation in the data, we estimate “the year of entering a higher school or its equivalent.” We first find the exact year of entering a higher school for all top-ranking officials who graduated from Schools 1–8 by searching their names in Student Registers of Schools 1–8.
+Using this information, we then estimate “the year of entering a higher school or its equivalent” for the rest of higher civil officials using the method described in Appendix A.4.2.
 Out of 6,255 exam passers in our dataset, 55.8% are Schools 1–8 graduates and 15.7% are top-ranking officials.
@@
 More descriptive statistics are shown in Appendix Table A.2.
-Effect on Higher Civil Service Exams Passers
+Effect on Higher Civil Service Exam Passers
 As an intermediate outcome, we first examine the impacts of the centralized admission system on the number of individuals who passed the selective HCSE.
-We divide exam passers into three mutually exclusive subgroups: (a) those who graduated from School 1, (b) those who graduated from Schools 2–8, and (c) those who are not Schools 1–8 graduates.
+We divide exam passers into three mutually exclusive subgroups: (a) those who graduated from most selective urban School 1, (b) those who graduated from Schools 2–8, and (c) those who are not Schools 1–8 graduates.
 For each subgroup, we count the number of exam passers by cohort.
@@
 $Y_{t}=\theta Centralized_{t}+\xi_{1}X_{t}+\xi_{2}Trend_{t}+\xi_{2}Trend_{t}^{2}+\omega_{t},$
-where $Y_{t}$ is the number of exam passers in a given group in cohort $t$ (defined by the year of entering a higher school or its equivalent), and $Centralized_{t}$ is the indicator that takes 1 if cohort $t$ entered a higher school or its equivalent during Capp.
+where $Y_{t}$ is the number of exam passers in a given group in cohort $t$ (defined by the year of entering a higher school or its equivalent), and $Centralized_{t}$ is the indicator that takes 1 if cohort $t$ entered a higher school or its equivalent during centralization.
 For a subgroup regression, we control for the total number of exam passers in cohort $t$ (denoted by $X_{t}$).
 We also control for a quadratic time trend where $Trend_{t}$ is the number of years since 1897.
-Capp changed the composition of exam passers.
-In Table 4 Panel (a), Column (3) shows that Capp increased the number of exam passers who graduated from Schools 2–8 by 23%, while Column (4) indicates that Capp reduced the number of exam passers who are not Schools 1–8 graduates by 13%.
-By contrast, Capp did not have significant effect on the total number of exam passers or the number of exam passers who graduated from School 1 (Columns (1) and (2)).
-In other words, the centralization had little impact on the total number of exam passers, but had a major impact on the composition of exam passers, and its positive effect was concentrated on Schools 2–8 graduates.
-This finding is consistent with the result of our short-run analysis that the centralization led to a large increase in Tokyo-area-born high-achieving entrants to Schools 2–8 (Table 1 Columns (3)–(9)), which likely improved the academic standing of these schools.
+Centralization changed the composition of exam passers.
+In Table 4 Panel (a), Column (3) shows that centralization increased the number of exam passers who graduated from Schools 2–8 by 23%, while Column (4) indicates that centralization reduced the number of exam passers who are not Schools 1–8 graduates by 13%.
+By contrast, centralization did not have a significant effect on the total number of exam passers or the number of exam passers who graduated from School 1 (Columns (1) and (2)).
+This reflects the fact that the total number of exam passers was determined mainly by demand-side factors (e.g., ministries’ needs) rather than supply-side factors (e.g., the quality of exam takers).
+Thus, the main impact of centralization was not on the number but on the composition of exam passers, and its positive effect was concentrated on Schools 2–8 graduates.
+This finding is consistent with the result of our short-run analysis that centralization led to a large increase in Tokyo-area-born high-achieving entrants to Schools 2–8 (Table 1 Panel (b) Columns (2)–(8)), which likely improved the academic standing of graduates of these schools.
 Effect on Top-Ranking Higher Civil Officials
-Finally, we analyze the impact of the centralization on the national production of top-ranking officials.
+We now analyze the impact of centralization on the national production of top-ranking officials.
 Our main outcome variable is the number of higher civil officials who were internally promoted to the top three ranks by the end of their career.
-We count the number of top-ranking officials by cohort and run the same regression as above for the entire group and for each of the following subgroups: (a) those who graduated from School 1, (b) those who graduated from Schools 2–8, and (c) those who are not Schools 1–8 graduates.
+We count the number of top-ranking officials by cohort and run the same regression as above for each group.
 For all regressions, we control for the total number of exam passers from the cohort.
 Meritocratic centralization produced a larger number of high-quality bureaucratic elites for the whole country.
-Table 4 Panel (b) Column (1) shows that the total number of top-ranking officials increased for Capp cohorts by 15%, with high statistical significance.
-This result is consistent with the idea that the centralized system (which assigned top-scoring students to Schools 1-8) increased the quality of Schools 1–8 entrants, which in turn improved the likelihood of Schools 1–8 graduates to win promotion competitions over close Dapp cohorts.
-Indeed, Column (3) indicates that Capp had a large, positive, and significant effect on the number of top-ranking officials who graduated from Schools 2–8, where the number of urban-born high-achieving entrants spiked during centralization.
-The number of top-ranking officials who graduated from Schools 2–8 increased by 47% for Capp cohorts compared to Dapp cohorts.
-By contrast, in Column (2), Capp only had a small and insignificant effect on the number of top-ranking officials who graduated from School 1, where the number of urban-born high-scoring entrants did not change much during centralization.
-Schools 1-8’s impacts on the likelihood of promotion to top-ranking officials appear especially large for high-achieving students.
-Indeed, Capp’s positive effect on Schools 2–8 graduates was so large (Column (3)) that it dominated Capp’s small and negative effect on those who did not graduate from Schools 1–8 (Column (4)), resulting in a positive effect on the total number of top-ranking officials (Column (1)).
-This finding suggests that the meritocratic assignment of high achievers to Schools 1-8 produced a net positive effect on the national production of top-ranking officials.
-This result is inconsistent with the pure selection hypothesis which argues that the role of the centralized admissions is simply to select and send a fixed number of high-achieving students to national higher education and that national higher education itself does not give students any value added.
-Schools 1-8’s particularly large value-added for high-achieving students can be due to multiple factors.
-First, human capital returns to education by these schools could be higher for high-achieving students than lower-achieving students (due to demanding curriculum, for example).
+Table 4 Panel (b) Column (1) shows that the total number of top-ranking officials increased for centralized cohorts by 15%, with high statistical significance.
+This result is consistent with the idea that the centralized system (which assigned top-scoring students to Schools 1–8) increased the quality of Schools 1–8 entrants, which in turn improved the likelihood of Schools 1–8 graduates to win promotion competitions over close decentralized cohorts.
+Indeed, Column (3) indicates that centralization had a large, positive, and significant effect on the number of top-ranking officials who graduated from Schools 2–8, where the number of urban-born high-achieving entrants spiked during centralization.
+The number of top-ranking officials who graduated from Schools 2–8 increased by 47% for centralized cohorts compared to decentralized cohorts.
+By contrast, in Column (2), centralization only had a small and insignificant effect on the number of top-ranking officials who graduated from School 1, where the number of urban-born high-scoring entrants did not change much during centralization.
+Schools 1–8’s impacts on the likelihood of promotion to top-ranking officials appear especially large for high-achieving students.
+Indeed, centralization’s positive effect on Schools 2–8 graduates was so large (Column (3)) that it dominated centralization’s small and negative effect on those who did not graduate from Schools 1–8 (Column (4)), resulting in a positive effect on the total number of top-ranking officials (Column (1)).
+This finding suggests that the meritocratic assignment of high achievers to Schools 1–8 produced a net positive effect on the national production of top-ranking officials.
+This result is inconsistent with the pure selection hypothesis that the role of the centralized admissions is simply to select and send a fixed number of high-achieving students to national higher education and that the value added of national higher education is homogeneous across all students.
+Schools 1–8’s particularly large value-added for high-achieving students can be due to multiple factors.
+First, human capital returns to education by these schools could be higher for high-achieving students (due to demanding curriculum, for example).
 Gathering high-achieving students in these schools may also produce higher peer effects.
@@
 While our data and variation do not allow for a clear decomposition into these different factors, our analysis suggests that the effect of meritocratic centralization remains similar even after controlling for the size of same-cohort colleagues from the same schools or focusing on those with the highest educational qualification (Appendix Section A.4.4).
-Much of Schools 1-8’s effects therefore likely come from the first human capital factor.
-One potential threat to our identification is that the number of available top-ranking positions might increase during the Capp periods.
+Much of Schools 1–8’s effects therefore likely come from the first human capital factor.
+One potential threat to our identification is that the number of available top-ranking positions might increase during the centralization periods.
 To investigate this concern, recall that we define our cohort by the year of entering a higher school or its equivalent (not by the year of becoming top-ranking officials).
@@
 This limits the concern that a potential correlation between the number of top-ranking positions and the lagged periods of centralized admissions may drive our results.
-Another potential threat is that the number of top-ranking officials may be differentially impacted by wars between Dapp and Capp cohorts.
-In particular, Japan was occupied by the US after WWII.
-During the postwar US Occupation, a substantial number of top officials were purged from public service, and a few were sentenced to death for war crimes.
-However, we show that the number of top-ranking officials who died in wars, or were purged or executed after WWII, changes little between Dapp and Capp cohorts (Appendix Table A.16).
-For career bureaucrats, the meritocratic centralization turns out to be more productively efficient than the decentralized system, as the government envisioned.
-Whether or not this was true for all career elites is an open question.
-For example, it is possible that the quality of career bureaucrats improved at the expense of other types of elites.
-To examine this possibility, we use the 1939 PIR data to compute the share of top bureaucrats among all “socially distinguished individuals” listed in the PIR.
-[^30] Top bureaucrats are defined as high-ranking central government officials listed in the PIR as in Section As Appendix Figure A.10 shows, we observe no positive association between the share of government officials and the cohort’s exposure to the centralized admissions (although the share of government officials is rising over time reflecting a growing public sector).
-Our result is unlikely to be driven by students’ brain drain from other elite occupations into top bureaucrats under the meritocratic centralization.
-Finally, we obtain suggestive evidence that the meritocratic centralization might produce a greater number of elites in general.
-As Appendix Figure A.11 and Table A.17 show, higher exposure to the centralized system is associated with a greater total number of individuals listed in the PIR.
-A potential concern with this analysis is that this may be driven by the age effect.
-For example, people of certain ages may have had a higher chance of holding important positions and being selected into the PIR, while such ages may happen to be correlated with exposure to centralization.
-To distinguish the cohort effect from the age effect, we use both 1934 and 1939 editions of the PIR while controlling for the age of each cohort.
-The resulting estimates show that the number of individuals listed in the PIR is 4–5% higher among Capp cohorts.
-The centralized system may therefore be more productively efficient not only for producing top bureaucrats but also for producing other types of career elites.
+Another potential threat is that the number of top-ranking officials may be differentially impacted by wars between decentralized and centralized cohorts.
+During the US occupation of Japan right after the WWII, a substantial number of top officials were purged from public service, and a few were sentenced to death for war crimes.
+However, we show that the number of top-ranking officials who died in wars, or were purged or executed after WWII, changes little between decentralized and centralized cohorts (Appendix Table A.17).
+The last potential concern is that the number of top bureaucrats might increase at the expense of other types of elites.
+To examine this possibility, we use the 1939 PIR data to compute the share of top bureaucrats among all socially distinguished individuals listed in the PIR.
+As Appendix Figure A.10 shows, we observe no positive association between the share of government officials and the cohort’s exposure to the centralized admissions (although the share of government officials is rising over time reflecting a growing public sector).
+Therefore, our result is unlikely to be driven by students’ brain drain from other elite occupations into top bureaucrats under the centralized admissions.
+Effect on All Occupational Elites
+Moreover, we find that meritocratic centralization produced more occupational elites across fields, not just government officials.
+As Appendix Figure A.11 shows, higher exposure to the centralized system is associated with a greater total number of all occupational elites (as defined in Section 4.1) listed in the PIR (1939).
+A potential concern is that this analysis may conflate age effects and cohort effects because the centralized cohorts, who were in their fifties in the PIR (1939), may have a higher likelihood of holding prominent positions simply due to their age rather than the centralized system’s impact.
+To distinguish the cohort effect from the age effect, we additionally use the 1934 edition of the PIR, in which we observe the centralized cohorts in their late forties.
+The total number of all occupational elites listed in the PIR (1934) also increased for the centralized cohorts, as Appendix Figure A.11 shows.
+Table 5 shows this result in a regression analysis using both 1934 and 1939 editions of the PIR.
+We count the number of individuals who are aged 40–69 by birth cohort in each PIR edition, and obtain the cohort-edition level data.
+We regress this count on each cohort’s exposure to centralization, while controlling for non-linear age trends defined by either quadratic age trends, quartic age trends, or edition-specific quadratic age trends.
+We also control for the edition fixed effect.
+Panel A shows that the total number of all occupational elites listed in the PIR is 5–13% higher among centralized cohorts.
+Looking at some specific groups of elites, the number of top 0.05% income earners and the number of medal recipients increased by 8–10% (Panel B) and 12-24% (Panel C), respectively, among centralized cohorts.
+The centralized system may therefore be more productively efficient not only for producing top bureaucrats but also for producing other types of individuals in leadership positions.
 5 Conclusion
-The design of school admissions persistently impacts the production and geography of career elites.
-We reveal this fact by looking at the world’s first recorded use of nationally centralized admissions and its subsequent abolitions in early twentieth-century.
-While centralization was designed to make the school seat allocations more meritocratic, there turns out to be a tradeoff between meritocracy and equal regional access to elite higher education and career advancement.
-The meritocratic centralization led students to apply to more selective schools and make more inter-regional applications.
-As high-achieving students were located disproportionately in urban areas, however, centralization caused urban applicants to crowd out rural applicants from advancing to higher education.
-Most importantly, these impacts persist: Several decades later, the centralized system produced more top-ranking government officials, indicating that the meritocratic centralization may be more productively efficient.
-The distributional effect also persists: The meritocratic centralization increased the number of high income earners, medal recipients, and other elite professionals born in urban areas relative to those born in rural areas.
+The design of selective school admissions persistently impacts the production and distribution of society’s leaders.
+We reveal this fact by looking at the world’s first recorded use of nationally centralized admissions and its subsequent abolitions in the early twentieth-century.
+While centralization was designed to make the school seat allocations more meritocratic, there turns out to be a tradeoff between meritocracy and equal access to selective higher education and career advancement.
+Meritocratic centralization led students to apply to more selective schools and make more inter-regional applications.
+As high-achieving students were located disproportionately in urban areas, however, centralization caused urban high achievers to crowd out rural applicants from advancing to higher education.
+Most importantly, these impacts persist: Several decades later, the centralized system produced more individuals in leadership positions, especially top-ranking government officials.
+The distributional effect also persists: Meritocratic centralization increased the number of high income earners, medal recipients, and other occupational elites coming from urban areas relative to those from rural areas.
 Even though our study uses the admission reforms unique to our historical setting, the implications of our study may have broader relevance.
 For instance, distributional consequences of centralized meritocratic admissions may be a reason why many countries continue to use seemingly inefficient decentralized college admissions.
-While this paper focuses on the decentralized system with a single application, the takeaway from our results would be relevant for other types of decentralized admissions as long as costly applications make applicants self-select into a limited number of schools.
-For any such decentralized systems, the meritocracy-equity tradeoff would be an important potential concern.
+While this paper focuses on the particular decentralized system, the takeaway from our results would be relevant for other types of decentralized admissions as long as costly applications make applicants self-select into a limited number of schools.
+For any such decentralized systems, the meritocracy-equity tradeoff is an important concern.
 Methodologically, the use of natural experiments in history may be also valuable for studying the long-run effects of market designs in other areas, such as housing, labor, and health markets.
-It is the multiple bidirectional policy changes in history that allow us to measure the long-run effects.
 The disadvantage of using historical events is the limited availability of data.
 The ideal way to alleviate the data concerns would be to use modern administrative data.
-For example, one may imagine linking administrative tax return data and school district data to measure the long-run effects of school choice reforms around the world in the past few decades.
+One may imagine linking administrative tax return data and school district data to measure the long-run effects of school choice reforms around the world in the past few decades.
 Such an effort would be a fruitful complement to our historical study.
-A. Baseline Specification
-Age 17 under centralization 3.18 0.61 1.68 2.82 1.84 0.84 0.41
-  $\times$ Tokyo area ($<$100 km) (0.027)** (0.016)** (0.060)* (0.010)*** (0.088)* (0.002)*** (0.025)**
-[0.000]*** [0.020]** [0.008]*** [0.001]*** [0.045]** [0.016]** [0.182]
-B. With Control Variables
-Age 17 under centralization 1.93 0.52 1.59 2.44 1.57 0.63 0.39
-  $\times$ Tokyo area ($<$100 km) (0.002)*** (0.025)** (0.026)** (0.002)*** (0.071)* (0.001)*** (0.041)**
-[0.014]** [0.017]** [0.007]*** [0.000]*** [0.040]** [0.041]** [0.143]
-C. Bidirectional Specification with Control Variables
-[…]
-  $\times$ Tokyo area ($<$100 km) (0.165) (0.078)* (0.006)*** (0.004)*** (0.057)* (0.010)*** (0.149)
-[0.034]** [0.049]** [0.006]*** [0.003]*** [0.023]** [0.170] [0.359]
-[…]
-  $\times$ Tokyo area ($<$100 km) (0.001)*** (0.022)** (0.183) (0.004)*** (0.338) (0.006)*** (0.025)**
-[0.005]*** [0.131] [0.021]** [0.008]*** [0.488] [0.032]** [0.079]*
-D. Centralization Exposure with Control Variables
-Cohort’s exposure to centralization 2.03 0.64 1.73 2.51 1.73 0.60 0.43
-  $\times$ Tokyo area ($<$100 km) (0.001)*** (0.008)*** (0.020)** (0.002)*** (0.024)** (0.006)*** (0.074)*
-[0.005]*** [0.001]*** [0.007]*** [0.004]*** [0.055]* [0.017]** [0.188]
-Observations 705 705 705 705 705 705 705
-Birth cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes
-Mean dep var 8.77 1.24 5.19 6.21 7.99 1.34 0.82
-Mean dep var (Tokyo area 10.62 1.51 6.75 6.86 11.21 1.25 1.02
-   under decentralization)
-[…]
-Table 3: Long-run Impacts of Centralization: Destinations of Career Elites
-(1) (2) (3) (4) (5) (6) (7)
-Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial
-Univ. income income recipients executives politicians & Univ.
-grads earners earners bureaucrats professors
-A. Baseline Specification
-Age 17 under centralization 4.20 0.48 2.03 3.11 3.47 0.61 0.71
-  $\times$ Tokyo area ($<$100 km) (0.060)* (0.151) (0.320) (0.056)* (0.238) (0.000)*** (0.120)
-[0.000]*** [0.154] [0.060]* [0.005]*** [0.003]*** [0.077]* [0.092]*
-B. Adding Control Variables
-Age 17 under centralization 3.16 0.74 2.44 2.90 3.64 0.41 0.54
-  $\times$ Tokyo area ($<$100 km) (0.021)** (0.221) (0.263) (0.038)** (0.223) (0.017)** (0.043)**
-[0.008]*** [0.010]*** [0.012]** [0.010]*** [0.002]*** [0.122] [0.134]
-Observations 705 705 705 705 705 705 705
-Birth cohort FE, Pref. FE Yes Yes Yes Yes Yes Yes Yes
-Mean dep var 8.30 1.21 5.08 5.88 7.74 1.27 0.78
-Mean dep var (Tokyo area 14.73 2.46 9.60 9.64 15.54 2.14 1.32
-   under decentralization)
-Notes: This table shows difference-in-differences estimates of the long-run effects of the centralized admission system on the geographical destinations of career elites.
-By birth cohort (born in 1880–1894), we count the number of elites who reside in each prefecture in 1939.
-Unlike the previous tables, all outcome variables are measured at the prefecture of residence.
-In Panel B, we control for time- and cohort-varying prefecture characteristics, i.e., the number of primary schools in the prefecture in the year when the cohort turned age 17, the number of middle-school graduates in the prefecture in the year when the cohort turned age 17, log GDP of the prefecture when the cohort turned age 20, and birth population of the cohort in the prefecture.
-Parentheses contain p-values based on standard errors clustered at the prefecture level.
-Square brackets contain wild cluster bootstrap p-values based on standard errors clustered at the cohort level.
-***, **, and * mean significance at the 1%, 5%, and 10% levels, respectively.
-See Section 4.1 for discussions about this table.
-Table 4: Long-run Impacts of Centralization: National Production of Top Government Officials
-(1) (2) (3) (4)
-Exam passers Exam passers Exam passers Exam passers
-graduated graduated not
-from from graduated
-School 1 Schools 2–8 from
-(Tokyo) Schools 1–8
-Centralized 5.96 -2.81 14.93 -12.11
-(0.802) (0.221) (0.014)** (0.044)**
-Observations 33 33 33 33
-Control time trend Yes Yes Yes Yes
-Control exam passers No Yes Yes Yes
-Mean dep var 189.55 37.29 68.47 83.80
-Mean dep var (decentralization) 194.22 37.96 64.09 92.17
-(1) (2) (3) (4)
-Top-ranking Top-ranking Top-ranking Top-ranking
-officials officials officials officials
-graduated graduated not
-from from graduated
-School 1 Schools 2–8 from
-(Tokyo) Schools 1–8
-Centralized 4.19 -0.48 5.22 -0.55
-(0.006)*** (0.648) (0.000)*** (0.596)
-Observations 33 33 33 33
-Control time trend & exam passers Yes Yes Yes Yes
-Mean dep var 29.77 8.27 12.97 8.53
-Mean dep var (decentralization) 28.66 8.30 11.22 9.14
-Notes: Panel (a) shows OLS estimates of the effects of the centralized admissions on the number of individuals who passed the administrative division of the Higher Civil Service Exams (administrative HCSE).
-Panel (b) shows OLS estimates of the effects of the centralized admissions on the number of top-ranking higher civil officials.
-The estimates are based on the cohort level data (1898–1930), where cohort is defined by the year of entering a higher school or its equivalent.
-The data is compiled from the complete list of individuals who passed the administrative HCSE in 1894–1941 and their biographical information.
-“Exam passers” is the number of individuals in cohort $t$ who passed the administrative HCSE.
-“Top-ranking officials” is the number of top-ranking officials in cohort $t$ (i.e., the number of individuals who entered a higher school or its equivalent in year $t$, passed the administrative HCSE, and were internally promoted to the top three ranks of higher civil service in their lifetime).
-“Centralized” is the indicator variable that takes 1 if cohort $t$ entered a higher school or its equivalent under the centralized admissions in 1902–07, 1917–18, and 1926–27.
-“Mean dep var (decentralization)” is the mean of the dependent variable for the cohorts who entered a higher school or its equivalent under the decentralized admissions.
-In all regressions, we control for quadratic time trends.
-Parentheses contain P values based on Newey-West standard errors with the maximum lag order of 3.
-***, **, and * mean significance at the 1%, 5%, and 10% levels, respectively.
-See Section 4.2 for discussions about this table.
 Appendix A Online Empirical Appendix
 A.1 The Evolution of the Admission System
-Table A.1 shows changes in the number and the admission system of National Higher Schools 1-8 (Kanritsu Koutou Gakkou in Japanese) from 1900 to 1930.313131National Higher Schools were initially called the University Preparatory Division of National Higher School.
-See Moriguchi (2021) for historical details and the sources of information in Table A.1.
-Despite the population increase and the growing demand for higher education, the number of Schools 1-8 increased only slightly from 6 to 8 in 1900–1918 due to tight fiscal constraints.
-With the economic boom of WWI, the government expanded the higher education system and increased the number of Schools from 8 in 1918 to 25 in 1925.323232In the late 1920s, in addition to 25 national higher schools, 3 local public higher schools, 4 private higher schools, and one colonial higher school were established.
+Table A.1 shows changes in the number and the admission system of Schools 1–8 (Kanritsu Koutou Gakkou in Japanese) from 1900 to 1930.
+Despite the population increase and the growing demand for higher education, the number of National Higher Schools increased only slightly from 6 to 8 in 1900–1918 due to tight fiscal constraints.
+With the economic boom of WWI, the government expanded the higher education system and increased the number of Schools from 8 in 1918 to 25 in 1925.333333In the late 1920s, in addition to 25 national higher schools, 3 local public higher schools, 4 private higher schools, and one colonial higher school were established.
 Such expansions notwithstanding, Schools 1–8 remained the most distinguished among all higher schools throughout the pre-WWII period.
@@
 […]
+[…]
 A.3 Data for the Short-run Analysis
-To analyze short-run effects of the centralized admission system (Capp), we collect data on application and enrollment outcomes for Schools 1–8 by newly digitizing several administrative and non-administrative sources.
+To analyze short-run effects of the centralized admission system, we collect data on application and enrollment outcomes for Schools 1–8 by newly digitizing several administrative and non-administrative sources.
 A.3.1 Data on Application Behavior
-First, we collect data on the number of applicants by school of their first choice (hereafter first-choice school) for 1900–1930 from the following sources: Ministry of Education Yearbook (Monbushou Nenpou in Japanese) for 1900, 1901, 1908–1916, 1919–1925, and 1928–1930; [^33] Digital images are available online at the National Diet Library Digital Collections.
+We collect data on the number of applicants by school of their first choice (hereafter first-choice school) for 1900–1930 from the following sources: the Ministry of Education Yearbook (Monbushou Nenpou in Japanese) for 1900, 1901, 1908–1916, 1919–1925, and 1928–1930; [^34] Digital images are available online at the National Diet Library Digital Collections.
 Correspondences from the Ministry of Education to the Tokyo Imperial University for 1903 and 1904; [^35] Digital images are available online at the University of Tokyo Digital Archives.
@@
 no data is available for 1902, 1905, 1906, and 1926.
-In addition, we collect data on the number of entrants by school for 1900–1930 from Ministry of Education Yearbook.
-First-choice school is defined as the school to which an applicant apply under the decentralized admission system and the school which an applicant rank as his first choice under the centralized admission system.
+In addition, we collect data on the number of entrants by school for 1900–1930 from the Ministry of Education Yearbook.
+First-choice school is defined as the school to which an applicant applies under the decentralized admission system and the school which an applicant ranks as his first choice under the centralized admission system.
 There is one issue in this data that requires data cleaning.
@@
 Using this data, we compute the share of applicants who choose School 1 as their first choice (in Figure 1(a)) and the ratio of first-choice applicants to entrants for each school (in Appendix Figure A.4).
-Taking advantage of more detailed data in a supplementary volume of Higher School Entrance Examination Investigation Report of 1917, we collect data on the number of applicants by their first-choice school and by their middle-school prefecture for 1916 and 1917.
+Taking advantage of more detailed data in a supplementary volume of the Higher School Entrance Examination Investigation Report of 1917, we collect data on the number of applicants by their first-choice school and by their middle-school prefecture for 1916 and 1917.
 Middle-school prefecture is defined as the prefecture in which an applicant’s middle school is located.
@@
 A.3.2 Data on Enrollment Outcomes
-To analyze enrollment outcomes, we use Higher School Student Registers (Gakkou Ichiran in Japanese) published annually by Schools 1–8.363636Digital images are available online at the National Diet Library Digital Collections.
+To analyze enrollment outcomes, we use Higher School Student Registers (Gakkou Ichiran in Japanese) published annually by Schools 1–8.373737Digital images are available online at the National Diet Library Digital Collections.
 For each school, as a proxy for the number of entrants, we collect data on the number of first-year students in the university preparatory course (daigaku yoka in Japanese).
@@
 Birth prefecture is the prefecture of a student’s legal domicile (honsekichi in Japanese) recorded in the official family registry system (koseki in Japanese).
-We include students born in any prefectures, but exclude foreign-born students and students born in colonies.
+We include students born in any prefecture, but exclude foreign-born students and students born in colonies.
 To measure the geographical mobility of entrants in Figure 1(b), we define enrollment distance (i.e., the distance between an entrant’s birth prefecture and the school he entered) by the direct distance between the capital of the birth prefecture and the capital of the prefecture in which the school he entered is located.
@@
 As Schools 7 and 8 are allowed to conduct their entrance exams on different dates from other schools in 1901 (School 7), 1908 (Schools 7 and 8), and 1909–1910 (School 7), applicants can apply for these schools in addition to one of the rest of the schools in these years even under the decentralized system.
-We exclude these schools in these years from the sample, since the feature of multiple-school application contaminates the essential aspect of the decentralized system.
+We exclude these schools in these years from the sample.
 To analyze the impact of centralized admissions on the geographical composition of entrants (Figure 2(a), Figure 2(b), and Table 1), we control for the size of potential applicants and the number of competing schools in each prefecture, using the following variables.
-From Ministry of Education Yearbook, we collect data on the number of middle-school graduates (including both public and private schools) by prefecture (defined by school location)  [^38] The data on middle-school graduates is missing for 1920 and is linearly interpolated.
+From the Ministry of Education Yearbook, we collect data on the number of middle-school graduates (including both public and private schools) by prefecture (defined by school location) [^39] The data on middle-school graduates is missing for 1920 and is linearly interpolated.
 as well as the number of higher schools (including national, public, and private schools) by prefecture (defined by school location) for 1900–1930.
 The total number of higher schools increased from 8 (consisting of 8 national schools) in 1918 to 32 (consisting of 25 national, 3 public, and 4 private schools) in 1930.
-The Tokyo area is defined by seven prefectures located within 100 km from Tokyo (i.e., Chiba, Gumma, Ibaraki, Kanagawa, Saitama, Tochigi, and Tokyo prefectures) measured by the distance between two prefectural capitals (Appendix Table A.3).
 […]
-Finally, to test the exogeneity of the timing of admission reforms, in addition to the data collected above, we collect data on the mean age of entrants for 1905–1930 from Ministry of Education Yearbook.
+Finally, to test the exogeneity of the timing of admission reforms, in addition to the data collected above, we collect data on the mean age of entrants for 1905–1930 from the Ministry of Education Yearbook.
 [^40] No data on the mean age of entrants is available for 1900–1904.
-We also collect data on government expenditures on national higher education (the sum of ordinary and extraordinary expenditures spent on National Higher Schools and Imperial Universities) for 1900–1930 from Ministry of Education Yearbook.
+We also collect data on government expenditures on national higher education (the sum of ordinary and extraordinary expenditures spent on National Higher Schools and Imperial Universities) for 1900–1930 from the Ministry of Education Yearbook.
 Our analysis is presented in Appendix Table A.7.
 A.4 Data for the Long-run Analysis
-To assess long-run effects of the centralized admission system (Capp), we provide two sets of empirical analyses.
-In the first analysis, we compile prefecture-cohort level data of career elites and compare career outcomes of urban- and rural-born individuals across birth cohorts that differed in their exposure to Capp.
-In the second analysis, we construct cohort level data of central government officials using the complete list of individuals who have passed the Higher Civil Service Examinations and compare promotion outcomes of these individuals across cohorts that differed in their exposure to Capp.
-A.4.1 Data on the Regional Distribution of Career Elites
+A.4.1 Data on the Regional Distribution of Elites
 Personnel Inquiry Records (PIR) data
-To measure students’ career outcomes, the first long-run analysis uses Personnel Inquiry Records (Jinji Koushin-roku in Japanese), which compiles a highly selective list of “socially distinguished individuals” (shakai-teki meishi in Japanese) and provides their biographical information.
+To measure students’ career outcomes, the first long-run analysis uses Personnel Inquiry Records (Jinji Koushin-roku in Japanese), which compiles a selective list of “socially distinguished individuals” (shakai-teki meishi in Japanese) and provides their biographical information.
 We use the 1939 edition of the PIR in our main analysis and the 1934 edition for a supplementary analysis.
-[^40] Digital images of the PIR are available online at the National Diet Library Digital Collections.
-We thank Hidehiko Ichimura and Yasuyuki Sawada for sharing their digitized PIR data.
-In selecting the socially distinguished individuals, the PIR refers to multiple sources such as the government personnel directory (Shokuin-roku in Japanese), the directories of banks and companies (Teikoku Ginkou Kaisha Youroku and Zenkoku Ginkou Kaisha Youroku in Japanese), the directory of the national chamber of commerce and industry members (Zenkoku Shoukou Kaigisho Giin Meibo in Japanese), and the directory of Japanese notables (Nihon Shinshi-roku in Japanese) (see PIR 1934, p.2, for details).
-[^41] Digital images of these directories are available online at the National Diet Library Digital Collections.
-The government personnel directory provides a complete list of public servants (with their job titles and ranks) in national and local governments, including the national diet members, civil and military officials, and Imperial University professors.
-One of the directories of banks and companies provides a complete list of directors of banks and companies whose capital is 300,000 JPY or above (Teikoku Ginkou Kaisha Youroku 1938, p.2).
-The directory of Japanese notables includes “wealthy persons” (shisanka in Japanese, defined by individuals who paid income tax of 80 JPY or greater or business tax of 70 JPY or greater) living in urban areas in 24 prefectures (out of 47 prefectures) in Japan (Nihon Shinshi-roku, 1938, p.i).
+[^41] We thank Hidehiko Ichimura and Yasuyuki Sawada for sharing their digitized the 1934 and 1939 PIR data.
+In compiling a list of socially distinguished individuals, the PIR explicitly states that it used multiple sources, including the government personnel directory (Shokuin-roku in Japanese), the directories of banks and companies (Teikoku Ginkou Kaisha Youroku and Zenkoku Ginkou Kaisha Youroku in Japanese), the directory of the national chamber of commerce and industry members (Zenkoku Shoukou Kaigisho Giin Meibo in Japanese), and the directory of Japanese notables (Nihon Shinshi-roku in Japanese) (PIR (1934), p.2).
+[^42] Digital images of these materials are available online at the National Diet Library Digital Collections.
+The government personnel directory provides a complete list of public servants (with their job titles and ranks) in national and local governments, including the Imperial Diet members, civil and military officials, and Imperial University professors.
+The directory of banks and companies provides a complete list of directors of banks and companies whose capital is 300,000 JPY or above (Teikoku Ginkou Kaisha Youroku (1938), p.2).
+The directory of Japanese notables includes “wealthy persons,” defined by individuals who paid income tax of 80 JPY or greater or business tax of 70 JPY or greater, living in urban areas in 24 prefectures (out of 47 prefectures) in Japan (Nihon Shinshi-roku (1938), p.i).
 We use the following information from the PIR for each individual: full name, birth year, birth prefecture, the prefecture of residence, final education, occupation titles and organization names, corporation type (if the organization is incorporated), the medal for merit and the court rank awarded (if any), and the amounts of national income tax and business tax paid in the previous year (if any).
-To check for accuracy, the PIR verifies the information of birth date and birth prefecture for each individual by obtaining a transcript of the official family register (PIR 1934, p.2).
-Defining the Groups of Career Elites
-As described in the main text, we define (mutually non-exclusive) groups of elites among the individuals listed in the PIR.
-We also define landlords as individuals whose occupations include landlord, but excluding the top 0.05% income earners, prestigious medal recipients, corporate executives, top politicians and bureaucrats, and Imperial University professors.
-To determine individual’s occupations, we check if one’s occupational titles and organization names contain specific Japanese characters that signify a given occupation.
+To check for accuracy, the PIR verifies the information of birth date and birth prefecture for each individual by obtaining a transcript of the official family register (PIR (1934), p.2).
+Defining the Groups of Elites
+As described in the main text, we define (mutually non-exclusive) subgroups of elites among the individuals listed in the PIR.
+To determine an individual’s occupations, we search if one’s occupational titles and organization names contain specific Japanese characters that signify a given occupation.
 The precise definitions of corporate executives, top politicians and bureaucrats, Imperial University professors, landlords, scholars, engineers, physicians, and lawyers are provided in Appendix Table A.9.
@@
 […]
+Finally, we define all occupational elites as all individuals listed in the PIR except for those individuals whose only occupation is peerage membership or landlord.
+We exclude these elites because they became elites typically by inheritance.
 Control Variables
@@
 […]
-Third, to control for local demographic changes, we estimate the prefecture-level male birth population of the cohorts born in 1874–1894 using the following data sources: (a) Japanese Population Census (Nihon Zenkoku Kokou-hyo in Japanese) in 1880–1892 that provides the number of male births by prefecture in each year and (b) Japanese Imperial Population Census (Nihon Teikoku Minseki Kokou-hyou in Japanese) in 1886 that provides the age-specific population by prefecture (from which we use the population of males who were born in 1874–1885 and were 1–12 years old in 1886).
+Third, to control for local demographic changes, we estimate the prefecture-level male birth population of the cohorts born in 1874–1894 using the following data sources: (a) Japanese Population Census (Nihon Zenkoku Kokou-hyo in Japanese) in 1880–1892, which provides the number of male births by prefecture in each year, and (b) Japanese Imperial Population Census (Nihon Teikoku Minseki Kokou-hyou in Japanese) in 1886, which provides the age-specific population by prefecture (from which we use the population of males who were born in 1874–1885 and were 1–12 years old in 1886).
 [^44] Digital images of population censuses are available online at the National Diet Library Digital Collections.
@@
 Therefore, we impute them by the average birth population in the prefecture in 1891 and 1892.
-Estimating the Intensity of Exposure to Capp
+Estimating the Intensity of Exposure to Centralization
 Before 1918, the qualifications for the higher school entrance exam were male middle-school graduates who were at least 17 years old.
 Accordingly, applicants typically took the exam for the first time at age 17.
-Given this, in the regression analyses in Table 2 Panels A–C, Table A.14, and Table 3, we simply assume that, among cohorts who were born in 1880–1894 and turned age 17 in 1897–1911, only the cohorts who turned age 17 during the first period of centralized admissions (Capp) in 1902–1907 were fully exposed to Capp, while the rest of the cohorts were fully exposed to Dapp.
+Given this, in the regression analyses in Table 2 Panels A–C, Table A.15, and Table 3, we simply assume that, among cohorts who were born in 1880–1894 and turned age 17 in 1897–1911, only the cohorts who turned age 17 during the first period of centralized admissions in 1902–1907 were fully exposed to centralization, while the rest of the cohorts were fully exposed to decentralization.
 Precisely speaking, however, a nontrivial number of unsuccessful applicants retook the exam multiple times in the subsequent years at age greater than 17.
-For this reason, we also estimate each cohort’s intensity of exposure to Capp that takes into account the presence of exam re-takers as follows.
-We use a special table in Government Gazette No.5838 (published in December 17, 1902) that reports the number of applicants and the number of admitted applicants in 1902 by the year of middle-school graduation.
+For this reason, we also estimate each cohort’s intensity of exposure to centralization by taking into account the presence of exam re-takers as follows.
+We use a special table in the Government Gazette No.5838 (published in December 17, 1902) that reports the number of applicants and the number of admitted applicants in 1902 by the year of middle-school graduation.
 According to the table, 63.9% of the applicants in 1902 graduated from middle schools in 1902 (the same year), 27.3% in 1901 (the previous year), 7.2% in 1900 (two years before), 1.3% in 1899 (three years before), 0.2% in 1898 (four years before), and 0.1% in 1897 (five years before).
-We use the 1902 data to compute the total number of exams taken and the number of exams taken under Capp for each cohort, where cohort is defined by the year of middle-school graduation.
+We use the 1902 data to compute the total number of exams taken and the number of exams taken under centralized admissions for each cohort, where cohort is defined by the year of middle-school graduation.
 This exercise requires us to estimate two sets of parameters.
@@
 The probability of retaking the exam is estimated to be $r_{0}=0.660$, $r_{1}=0.423$, $r_{2}=0.255$, $r_{3}=0.175$, $r_{4}=0.011$, $r_{5}=0.000$.
-Using these parameters, we simulate the total number of exams taken by each cohort and compute the share of the exams taken under Capp, which we define as the intensity of exposure to Capp.
+Using these parameters, we simulate the total number of exams taken by each cohort and compute the share of the exams taken under centralized admissions, which we define as the intensity of exposure to centralization.
 Here we assume that all cohorts graduated from middle-schools at age 17 (so that birth cohorts and cohorts defined by the year of middle-school graduation coincide).
-The estimated intensity of exposure to Capp is 0.00, 0.02, 0.09, 0.36, 1.00, 0.98, 0.91, 0.64, and 0.00 for the cohort who turned age 17 in 1896–98, 1899, 1900, 1901, 1902–04, 1905, 1906, 1907, and 1908–11, respectively.
-We color cohorts according to their intensity of exposure to Capp in Figure 3 and Appendix Figures A.7, A.10, and A.11.
-We also use the estimated intensity of exposure in the regression analyses in Table 2 Panel D and Appendix Table A.17.
-A.4.2 Data on the National Production of Career Elites
-In the second set of long-run analysis, we focus on a specific group of career elites, i.e., higher civil officials for whom we have complete count data.
+The estimated intensity of exposure to centralization is 0.00, 0.02, 0.09, 0.36, 1.00, 0.98, 0.91, 0.64, and 0.00 for the cohort who turned age 17 in 1896–98, 1899, 1900, 1901, 1902–04, 1905, 1906, 1907, and 1908–11, respectively.
+We color cohorts according to their intensity of exposure to centralization in Figure 3 and Appendix Figures A.7, A.10, and A.11.
+We also use the estimated intensity of exposure in the regression analyses in Table 2 Panel D and Table 5.
+A.4.2 Data on the National Production of Top-Ranking Higher Civil Officials
+In the second set of long-run analysis, we focus on a specific group of elites, i.e., higher civil officials for whom we have complete count data.
 Higher Civil Service Examinations (HCSE) data
 Our main data source is Hata (1981), which provides not only the list of all individuals who passed the Higher Civil Service Examinations (HCSE, Bunkan Koutou Shiken in Japanese), but also their biographical information available as of 1981 (Hata, 1981, Section 3).
-The HCSE were highly selective national qualification exams held annually from 1894 to 1947.
+The HCSE were selective national qualification exams held annually from 1894 to 1947.
 The administrative division of the HCSE consisted of two parts: the preliminary exams and the main exams.
@@
 For a robustness check, we alternatively assume that the year of entering a higher school or its equivalent is “the year of university graduation minus 7” and show that the empirical results are qualitatively the same (see Appendix Table A.16).
-To proceed to an empirical analysis, we redefine cohort by “the year of entering a higher school or its equivalent” estimated above (so that we can see which cohorts are exposed to Capp as opposed to Dapp).
+To proceed to an empirical analysis, we redefine cohort by “the year of entering a higher school or its equivalent” estimated above (so that we can see which cohorts are exposed to centralization as opposed to decentralization).
 That is, to create cohort-level data, we count the number of exam passers (i.e., those who passed the administrative division of the HCSE) and the number of top-ranking officials by the year in which they entered a higher school or its equivalent.
 Our dataset consists of 6,255 individuals who entered a higher school or its equivalent in 1898–1930 and passed the HCSE in 1901–1941.
-Out of 6,255 exam passers, 3,490 individuals or 55.8% are Schools 1–8 graduates, 4,767 individuals or 76.2% are Imperial University graduates, and 982 individuals or 15.7% are top-ranking officials.
+Out of 6,255 exam passers, 3,490 individuals (55.8%) are Schools 1–8 graduates, 4,767 individuals (76.2%) are Imperial University graduates, and 982 individuals (15.7%) are top-ranking officials.
 The share of Imperial University graduates is greater than that of Schools 1–8 graduates because the number of higher schools increased from 8 to 25 in 1919–1930.
-Among 982 top-ranking officials, 701 officials or 71.4% are Schools 1–8 graduates, and 891 officials or 90.7% are Imperial University graduates.
+Among 982 top-ranking officials, 701 officials (71.4%) are Schools 1–8 graduates, and 891 officials (90.7%) are Imperial University graduates.
 The HCSE consisted of preliminary and main exams, and because Imperial University law graduates were the only group exempted from the administrative preliminary exams from 1894 to 1922, they had substantial advantages in passing the HCSE (Hata, 1981, pp.663-666; Spaulding, 1967, Chapter 12).
@@
 A.4.3 Variation in Years Taken to be Promoted to Top-ranking Positions
-One potential threat to our identification in Table 4 Panel (b) is a possibility that the number of available top-ranking positions happened to have increased during the periods of Capp.
+One potential threat to our identification in Table 4 Panel (b) is a possibility that the number of available top-ranking positions happened to have increased during the periods of centralization.
 However, we argue that even if this was the case, it is not likely to affect our results, since our cohort is defined by the year of entering a higher school or its equivalent, not by the year of becoming top-ranking officials.
 Namely, as long as individuals in a given cohort were not promoted to a top-ranking position in the same year, a potential correlation between the number of top-ranking positions and the lagged periods of centralized admissions does not bias our results.
-To examine this possibility, we first randomly selected two cohorts exposed to Capp (1903 and 1916) and two cohorts exposed to Dapp (1913 and 1922).
+To examine this possibility, we first randomly selected two cohorts exposed to centralization (1903 and 1916) and two cohorts exposed to decentralization (1913 and 1922).
 For all top-ranking officials who graduated from Schools 1–8 in these cohorts, we searched the years in which these officials were appointed to their first top-ranking positions.
@@
 This within-cohort variation is a sum of two variations, the first variation in the number of years taken from entering a higher school to passing the exams and the second variation in the number of years taken from passing the exams to becoming a top-ranking official.
-Therefore, it is unlikely that the result in Table 4 Panel (b) is driven by a greater number of available top-ranking positions that coincided with the periods of Capp.
+Therefore, it is unlikely that the result in Table 4 Panel (b) is driven by a greater number of available top-ranking positions that coincided with the periods of centralization.
 A.4.4 Connections and Signaling Hypotheses
-We investigate whether connections or signaling was an important mechanism through which the meritocratic centralization increased the total number of top-ranking officials.
+We investigate whether connections or signaling was an important mechanism through which meritocratic centralization increased the total number of top-ranking officials.
 First, to explore the connection channel, we reexamine the result of Table 4 Panel (b) Column (3) by controlling for the number of exam passers from the same schools (Schools 2–8) in the same cohort, instead of the total number of exam passers.
 This variable takes into account the size of same-cohort colleagues from the same schools who are likely to have network values.
-The coefficient of Capp remains sizable (2.72) and statistically significant at the 1% level, implying 24% increase in the number of top-ranking officials from these schools relative to the mean during Dapp (11.22).
+The coefficient of centralization remains sizable (2.72) and statistically significant at the 1% level, implying a 24% increase in the number of top-ranking officials from these schools relative to the mean during decentralization (11.22).
 Second, to investigate the signaling channel, we further restrict our sample to top-ranking officials who graduated from both Schools 2–8 and Tokyo Imperial University, the most prestigious institution with the strongest signaling value.
-We regress the number of this subgroup of top-ranking officials on Capp, controlling for the number of exam passers from the same schools in the same cohort.
-If signaling is a major channel, we expect the effect of Capp among this subgroup to be larger.
-The coefficient of Capp decreases to 1.76, implying 18% increase in the number of top-ranking officials from these schools relative to the mean during Dapp (10.00).
+We regress the number of this subgroup of top-ranking officials on centralization, controlling for the number of exam passers from the same schools in the same cohort.
+If signaling is a major channel, we expect the effect of centralization among this subgroup to be larger.
+The coefficient of centralization decreases to 1.76, implying a 18% increase in the number of top-ranking officials from these schools relative to the mean during decentralization (10.00).
 This effect size is smaller than the effect size of 24% above.
@@
 Light orange color (e.g.
-US): Decentralized college admissions where each college defines its own admission standards and rules.
+U.S.): Decentralized college admissions where each college defines its own admission standards and rules.
 Yellow with diagonal lines (e.g.
@@
 Notes: This figure shows the locations of Schools 1–8 and the location of the Tokyo area (in dark gray color) defined as a set of 7 prefectures that are within 100 km from Tokyo (i.e., Tokyo, Chiba, Kanagawa, Saitama, Ibaraki, Tochigi, and Gunma prefectures).
-See Sections 2 and 3.4 for discussions about this figure.
+See Sections 2 and 3.5 for discussions about this figure.
 Table A.1: The Evolution of the Admission System
-Notes: This table shows changes in the number and the admission system of Schools 1-8 from 1900 to 1930.
+Notes: This table shows changes in the admission system of National Higher Schools (including Schools 1–8) from 1900 to 1930.
 In 1901, all schools held their exams on the same date, with the exception of newly established School 7.
 In 1908, all schools held their exams on the same date, with the exception of Schools 7 and 8.
+See Moriguchi (2021) for historical details and the sources of information.
 See Section 2.1 and Appendix Section A.1 for discussions about this table.
@@
 Notes: This figure is a reprint of the assignment algorithm of the centralized admission system stated in the Ordinance of the Ministry of Education No.4 published in Government Gazette No.1419, pp.580-581, on April 27, 1917.
-See Sections 2 and 3.3 for an English translation and discussions.
+See Sections 2 and 3.4 for an English translation and discussions.
 Table A.2: Summary Statistics
@@
 No. of entrants to School 8 5.67 12.45 2 1093
-No. of public middle-school graduates 415.38 350.01 299 1410
-No. of private middle-school graduates 118.49 435.12 0 1410
+No. of middle-school graduates 533.35 617.82 359 1457
 No. of national higher schools other than Schools 1–8 0.14 0.43 0 1469
@@
 No. of civilians receiving medal of the Order of Fifth Class and above 6.21 4.94 5 705
-No. of corporate executives with a positive amount of tax payment 7.99 8.91 5 705
+No. of corporate executives with a positive amount of tax payment 8.44 9.57 6 705
 No. of top politicians and high-ranking bureaucrats 1.34 1.38 1 705
 No. of Imperial University professors 0.82 1.18 0 705
+No. of all occupational elites 38.00 35.29 28 705
 Cohort level data on long-run outcomes, government officials entering higher school or equivalent in 1898–1930
@@
 We control for year fixed effects, prefecture fixed effects, and the number of higher schools other than Schools 1–8 in prefecture $p$ in year $t$.
-We also control for “Born in school’s prefecture,” “Born near school’s prefecture (1-100 km),” and “Born near school’s prefecture (100-300 km)” as in Table 1.
+We also control for “Born in school’s prefecture,” “Born near school’s prefecture (1–100 km),” and “Born near school’s prefecture (100-300 km)” as in Table 1.
 Parentheses contain p-values based on standard errors clustered at the prefecture level.
@@
 education Tokyo area
-Centralized 1.357 -0.101 0.00556 -0.460 0.0224 0.0523 0.760 0.164 57.02 0.0394
+Centralized 1.357 -0.101 0.00556 -0.460 0.0224 0.0523 0.760 0.164 58.54 0.0392
 (0.418) (0.109) (0.104) (0.486) (0.110) (0.518) (0.676) (0.001)*** (0.000)*** (0.000)***
-No. of middle-school graduates 0.00493 -0.000942 -0.150 0.00134 0.0656 -0.429 0.00194 1.345 0.00147
+No. of middle-school graduates 0.00493 -0.000942 -0.150 0.00134 0.0656 -0.429 0.00194 1.448 0.00150
 (0.291) (0.022)** (0.046)** (0.165) (0.000)*** (0.062)* (0.468) (0.104) (0.033)**
 Observations 31 31 31 27 27 26 31 27 31 31
-Mean dep var 23.41 2.007 0.180 10.78 0.217 19.03 13.56 0.291 228.9 0.174
+Mean dep var 23.41 2.007 0.180 10.78 0.217 19.03 13.56 0.291 228.4 0.174
 Notes: Columns 1–7 test if important institutional variables are correlated with the timing of centralization using year-level data.
-Columns 8–10 examine the robustness of our main short-run outcomes using year-level data.
+Columns 8–10 do the same our main short-run outcomes using year-level data.
 All numbers are at the national-level from 1900 to 1930.
@@
 See Section 3.6 for discussions about this table.
-Table A.8: Definitions of Career Elites in the PIR
-Notes: This table provides the precise definitions of career elites, defined as a subset of individuals listed in PIR (1939). See Section 4.1 for the analysis using these variables.
+Table A.8: Occupations of Individuals Listed in the PIR (1939)
+(1) All individuals listed (2) Sampled cohorts (3) Top 0.05% income earners (4) Medal recipients
+Corporate executives and managers 0.660 0.672 0.803 0.320
+Politicians and bureaucrats 0.075 0.070 0.030 0.213
+Scholars, lawyers, and artists 0.119 0.142 0.048 0.473
+Engineers and physicians 0.090 0.110 0.048 0.259
+Military personnels 0.045 0.047 0.019 0.000
+Landlords 0.048 0.033 0.060 0.001
+Imperial and peerage family members 0.006 0.004 0.004 0.016
+None of the above 0.130 0.109 0.098 0.087
+No. observations 55,742 28,423 3,864 4,739
+Notes: This table shows occupational distributions of individuals listed in the PIR (1939).
+Column (1) shows the share of each specified occupational category among all individuals listed in the PIR (1939).
+Column (2) restricts the sample to cohorts who were born in 1880–1894.
+Columns (3) and (4) further restrict the sample to the top 0.05% income earners and civilian medal recipients (as defined in Section 4.1), respectively.
+Occupational categories are not mutually exclusive, except for “None of the above” which is defined as individuals not included in the specified categories.
+See Section 4.1 for discussions about this table.
+Table A.9: Definitions of Occupational Categories of Elites
+Notes: This table provides the precise definitions of occupational categories of elites.
+Individuals in each category are defined as a subset of individuals listed in the PIR.
+See Section 4.1 for the analysis using these categories.
 Figure A.5: Sampling Rates of High Income Earners in the PIR Data
 Notes: This figure plots the sampling rate of the high income earners listed in PIR (1939) by the income level expressed as a top percentile of the national income distribution.
-The sampling rates and the top income percentiles are computed from the complete count data in National Tax Bureau Yearbook.
+The sampling rates and the top income percentiles are computed from the complete count data in the National Tax Bureau Yearbook.
 The vertical lines indicate the top 0.05% and 0.01% thresholds used in our analysis.
@@
 Figure A.6: High Income Earners in the PIR vs Income Tax Statistics across Prefectures
-Notes: These figures compare the number of high income earners in each prefecture listed in PIR (1939) and the complete count of high income earners in each prefecture reported in National Tax Bureau Yearbook (1936).
+Notes: These figures compare the number of high income earners in each prefecture listed in the PIR (1939) and the complete count of high income earners in each prefecture reported in the National Tax Bureau Yearbook (1936).
 The vertical axis is the log of the number of individuals in the PIR who earned more than 50,000 JPY taxable income (corresponding to the top 0.01% income group) or 16,950 JPY taxable income (corresponding to the top 0.05% income group) in 1938.
@@
 earners earners earners earners
-Entrants to Schools 1–8 -0.000035 0.000056
+Entrants to Schools 1–8 -0.000078 0.000018
 (0.000060) (0.000036)
@@
 Observations 47 47 47 47
-R-squared 0.003527 0.033977 0.006313 0.008153
+R-squared 0.015 0.003 0.006 0.008
 Mean dep var 0.44 0.24 0.44 0.24
@@
 See Section 4.1 for discussions about this table.
-Figure A.7: Long-run Impacts of Centralization: Geographical Origins of Other Career Elites
+Figure A.7: Long-run Impacts of Centralization: Geographical Origins of Other Elites
 Notes: This figure shows additional difference-in-differences plots that compare the average number of elites born in prefectures inside and outside the Tokyo area by cohort (see Figure 3).
 The plots are based on the data from PIR (1939), which covers cohorts who were born in 1879–1894 and turned age 17 (main application age) in 1896–1911.
-The vertical axis shows the number of indicated elites who were born in the indicated area in the indicated cohort.
+The vertical axis shows the average number of indicated elites (per prefecture) born in the indicated area in the indicated cohort.
 The cohorts are colored according to their intensity of exposure to the centralized admissions in 1902–07, where the darker color indicates the higher intensity.
 See Section 4.1 for discussions about this figure.
-Table A.10: Long-run Impacts: Other Elite Professionals
+Table A.11: Long-run Impacts: Other Elites
 (1) (2) (3) (4)
 Scholars Engineers Physicians Lawyers
-Age 17 under centralization 0.87 0.66 0.46 0.18
-$\times$ Tokyo area ($<$100 km) (0.026)** (0.048)** (0.038)** (0.420)
-[0.094]* [0.125] [0.091]* [0.544]
+Age 17 under centralization 1.36 0.66 0.46 0.18
+$\times$ Tokyo area ($<$100 km) (0.004)*** (0.048)** (0.038)** (0.420)
+[0.003]*** [0.125] [0.091]* [0.544]
 Observations 705 705 705 705
 Birth cohort FE, Birth pref. FE Yes Yes Yes Yes
-Mean dep var 3.61 2.46 1.92 1.11
-Mean dep var (Tokyo area 4.76 3.13 2.52 1.16
+Mean dep var 4.29 2.46 1.92 1.11
+Mean dep var (Tokyo area 5.41 3.13 2.52 1.16
 under decentralization)
@@
 Table A.12: Long-run Impacts: Excluding Cohorts who Turned Age 17 in 1901 or 1907
-Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial
-Univ. income income recipients executives politicians & Univ.
-grads earners earners bureaucrats professors
-A. Baseline Specification
-Age 17 under centralization 3.05 0.68 1.65 2.61 1.89 0.68 0.45
-  $\times$ Tokyo area ($<$100 km) (0.003)*** (0.007)*** (0.048)** (0.007)*** (0.030)** (0.019)** (0.067)*
-[0.000]*** [0.026]** [0.011]** [0.000]*** [0.066]* [0.023]** [0.189]
-B. Adding Control Variables
-Age 17 under centralization 1.70 0.58 1.53 2.18 1.60 0.45 0.41
-   $\times$ Tokyo area ($<$100 km) (0.021)** (0.013)** (0.007)*** (0.000)*** (0.014)** (0.028)** (0.095)*
-[0.017]** [0.008]*** [0.020]** [0.006]*** [0.069]* [0.086]* [0.186]
-Observations 611 611 611 611 611 611 611
-Birth cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes
-Mean dep var 8.68 1.23 5.16 6.08 7.96 1.34 0.83
-Mean dep var (Tokyo area 10.38 1.41 6.57 6.61 10.91 1.21 1.02
+Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial All
+Univ. income income recipients executives politicians & Univ. occupational
+grads earners earners bureaucrats professors elites
+Panel A: Baseline Specification
+Age 17 under centralization 3.05 0.68 1.65 2.61 1.94 0.68 0.45 9.27
+    $\times$ Tokyo area ($<$100 km) (0.003)*** (0.007)*** (0.048)** (0.008)*** (0.020)** (0.019)** (0.067)* (0.068)*
+[0.000]*** [0.038]** [0.019]** [0.000]*** [0.081]* [0.022]** [0.193] [0.001]***
+Panel B: Adding Control Variables
+Age 17 under centralization 1.70 0.58 1.53 2.18 1.61 0.45 0.41 6.48
+    $\times$ Tokyo area ($<$100 km) (0.021)** (0.013)** (0.007)*** (0.001)*** (0.007)*** (0.028)** (0.095)* (0.007)***
+[0.019]** [0.010]** [0.013]** [0.001]*** [0.081]* [0.095]* [0.207] [0.009]***
+Observations 611 611 611 611 611 611 611 611
+Cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes Yes
+Mean dep var 8.68 1.23 5.16 6.08 8.41 1.32 0.83 37.72
+Mean dep var (Tokyo area 10.38 1.41 6.57 6.61 11.39 1.18 1.02 49.43
 under decentralization)
@@
 Table A.13: Long-run Impacts: Pre-event Trends Are Parallel
-Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial
-Univ. income income recipients executives politicians & Univ.
-grads earners earners bureaucrats professors
+Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial All
+Univ. income income recipients executives politicians & Univ. occupational
+grads earners earners bureaucrats professors elites
 A. Baseline Specification
-Tokyo area ($<$ 100 km) 0.19 -0.02 0.07 0.17 0.11 -0.01 0.02
-  $\times$ Time trend (0.673) (0.802) (0.822) (0.476) (0.760) (0.879) (0.835)
+Tokyo area ($<$ 100 km) 0.19 -0.02 0.07 0.17 0.11 -0.01 0.02 0.92
+     $\times$ Time trend (0.673) (0.802) (0.822) (0.476) (0.768) (0.879) (0.835) (0.511)
 B. Adding Control Variables
-Tokyo area ($<$ 100 km) 0.21 -0.01 0.10 0.18 0.14 -0.01 0.02
-  $\times$ Time trend (0.630) (0.836) (0.751) (0.448) (0.701) (0.885) (0.841)
+Tokyo area ($<$ 100 km) 0.21 -0.01 0.10 0.18 0.13 -0.01 0.02 1.01
+    $\times$ Time trend (0.630) (0.836) (0.751) (0.448) (0.703) (0.885) (0.841) (0.463)
 C. Squared Trend Term
-Tokyo area ($<$ 100km) $\times$ Time trend -0.93 -0.08 -0.48 0.02 0.09 -0.13 0.22
-(0.415) (0.817) (0.275) (0.981) (0.903) (0.629) (0.412)
-[…]
-(0.465) (0.824) (0.288) (0.878) (0.927) (0.695) (0.333)
-Observations 470 470 470 470 470 470 470
-Birth cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes
-Mean dep var 5.02 1.21 4.50 4.34 6.34 0.66 0.39
-Mean dep var (Tokyo area 6.47 1.29 5.96 5.23 8.80 0.61 0.54
+Tokyo area ($<$ 100km) -0.93 -0.08 -0.48 0.01 0.06 -0.13 0.22 -0.11
+    $\times$ Time trend (0.415) (0.817) (0.275) (0.990) (0.927) (0.629) (0.412) (0.923)
+Tokyo area ($<$ 100km) 0.07 0.00 0.03 0.01 0.00 0.01 -0.01 0.07
+    $\times$ Time trend2 (0.465) (0.824) (0.288) (0.870) (0.893) (0.695) (0.333) (0.586)
+Observations 470 470 470 470 470 470 470 470
+Cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes Yes
+Mean dep var 5.02 1.21 4.50 4.34 6.71 0.66 0.39 28.62
+Mean dep var (Tokyo area 6.47 1.29 5.96 5.23 9.21 0.61 0.54 38.97
      under decentralization)
@@
 For cohorts who were born in 1874–1883 and thus turned age 17 in 1891–1900, we run the following regression in Panels A and B:
+$Y_{pt}=\beta\times Trend_{t}\times Tokyo\_area_{p}+\alpha_{p}+\alpha_{t}+\epsilon_{pt},$
 where $Trend_{t}$ is defined as the cohort’s birth year minus 1870 (the linear time trend). In Panel C, for the same cohorts, we run the following regression:
+$Y_{pt}=\beta_{1}\times Trend_{t}\times Tokyo\_area_{p}+\beta_{2}\times(Trend_{t})^{2}\times Tokyo\_area_{p}+\alpha_{p}+\alpha_{t}+\epsilon_{pt},$
 In Panels B and C, we control for the number of primary schools in the prefecture in the year when the cohort turned eligible age and the birth population of the cohort in the prefecture.
@@
 Table A.14: Long-run Impacts: Difference-in-Differences Estimates Using the PIR in 1934
-(1) (2) (3) (4) (5) (6) (7)
-Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial
-Univ. income income recipients executives politicians & Univ.
-grads earners earners bureaucrats professors
-Age 17 under centralization 1.71 0.50 1.81 1.44 1.11 0.39 0.08
-  $\times$ Tokyo area ($<$100 km) (0.016)** (0.044)** (0.029)** (0.070)* (0.013)** (0.018)** (0.717)
-[0.004]*** [0.003]*** [0.004]*** [0.008]*** [0.008]*** [0.282] [0.496]
-Observations 705 705 705 705 705 705 705
-Birth cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes
-Mean dep var 4.79 0.81 3.64 3.86 2.80 1.18 0.70
-Mean dep var (Tokyo area 5.68 1.11 6.18 4.46 4.46 1.14 0.70
+(1) (2) (3) (4) (5) (6) (7) (8)
+Imperial Top 0.01% Top 0.05% Medal Corporate Top Imperial All
+Univ. income income recipients executives politicians & Univ. occupational
+grads earners earners bureaucrats professors elites
+Age 17 under centralization 1.71 0.50 1.81 1.44 1.20 0.39 0.08 4.86
+(0.016)** (0.044)** (0.029)** (0.070)* (0.026)** (0.018)** (0.717) (0.072)*
+[0.011]** [0.004]*** [0.005]*** [0.002]*** [0.001]*** [0.261] [0.690] [0.005]***
+Observations 705 705 705 705 705 705 705 705
+Cohort FE, Birth pref. FE Yes Yes Yes Yes Yes Yes Yes Yes
+Mean dep var 4.79 0.81 3.64 3.86 2.95 1.15 0.52 14.91
+Mean dep var (Tokyo area 5.68 1.11 6.17 4.46 4.63 1.13 0.70 22.35
      under decentralization)
@@
 A. Baseline Specification
-Age 17 under centralization 0.34 -0.10 -0.01 -18.41
+Age 17 under centralization 0.34 -0.12 -0.01 -18.41
      $\times$ Tokyo area ($<$100 km) (0.270) (0.474) (0.449) (0.173)
-[0.245] [0.818] [0.422] [ 0.307]
+[0.245] [0.767] [0.422] [0.307]
 B. Adding Control Variables
-Age 17 under centralization -0.06 0.01 -0.02 -23.44
+Age 17 under centralization -0.06 -0.01 -0.02 -23.44
     $\times$ Tokyo area ($<$100 km) (0.872) (0.955) (0.360) (0.173)
-[0.790] [0.969] [0.361] [0.259]
+[0.790] [0.978] [0.361] [0.259]
 Observations 705 705 705 705
 Birth cohort FE, Birth pref. FE Yes Yes Yes Yes
-Mean dep var 11.67 0.94 0.29 297.92
-Mean dep var (Tokyo area 13.18 2.89 0.37 463.52
+Mean dep var 11.67 0.93 0.29 297.92
+Mean dep var (Tokyo area 13.18 2.88 0.37 463.52
      under decentralization)
@@
 Notes: This figure plots the average number of years taken from entering a higher school to graduating from an Imperial University by cohort (where cohort is defined by the year of university graduation) for two groups of individuals who passed the administrative division of the Higher Civil Service Exams.
-The first group is top-ranking officials who graduated from Schools 1–8 (shown in black round markers).
+The first group is top-ranking officials who graduated from Schools 1–8 (shown in black circles).
 The second group is non-top-ranking officials who graduated from Schools 1–8 (shown in red rectangular markers for two representative years, 1914 and 1922, only).
@@
 See Section 4.2 for discussions about this table.
-Figure A.10: Share of Government Officials in PIR-listed Individuals
-Notes: This figure plots the share of central government officials in all individuals listed in PIR (1939) by birth cohort.
+Figure A.10: Share of Government Officials in Elites Listed in the PIR
+Notes: This figure plots the share of central government officials in all individuals listed in the PIR (1939) by birth cohort.
 The cohorts are colored according to their exposure to the centralized admissions in 1902–1907, where darker color indicates higher intensity of exposure.
-We observe no significant increase in the share of central government officials for the cohorts who are exposed to the centralized admissions.
+We observe no significant increase in the share of central government officials for the cohorts exposed to the centralized admissions.
 See Section 4.2 for discussions about this table.
-Figure A.11: Total Number of PIR-listed Individuals by Cohort
-Notes: This figure plots the total number of individuals listed in PIR (1939) and PIR (1934) by birth cohort.
+Figure A.11: Total Number of All Occupational Elites by Cohort
+Notes: This figure plots the total number of all occupational elites listed in the PIR (1939) and PIR (1934) by birth cohort.
 The cohorts are colored according to their exposure to the centralized admissions in 1902–1907, where the darker color indicates the higher intensity of exposure.
-In both PIR (1939) and PIR (1934), cohorts exposed to the centralization tend to have a greater number of individuals selected into the PIR.
-See Appendix Table A.17 for a statistical analysis and Section 4.2 for discussions about this figure.
-Table A.17: Long-run Impacts: Total Number of PIR-listed Individuals
-(1) (2)
-No. of Individuals in PIR
-(No. of individuals aged 50 in PIR = 100)
-Age 17 under centralization 3.77
-(0.086)*
-Cohort’s exposure to centralization 4.78
-(0.060)*
-Age 40.32 39.69
-(0.000)*** (0.000)***
-Age squared -0.39 -0.38
-(0.000)*** (0.000)***
-PIR (1939) dummy 0.37 0.37
-(0.808) (0.805)
-Observations 46 46
-Notes: This table examines the long-run effects of the centralized admissions on the total number of career elites in the PIR.
-To distinguish the cohort effect from the age effect, we use two editions of the PIR, i.e., PIR (1934) and PIR (1939), count the number of individuals aged 40–62 by birth cohort in each edition (see Appendix Figure A.11), and pool these data.
-We standardize the dependent variable by setting the number of individuals at age 50 in each edition to be 100.
-“Age 17 under centralization” takes 1 if the cohort turned 17 years old under the centralized system in 1902–1907, and takes 0 otherwise.
-“Cohort’s exposure to centralization” is defined as in Table 2 Panel D.
-We control for quadratic age trends and the edition fixed effect.
-Parentheses contain p-values based on robust standard errors.
-***, **, and * mean significance at the 1%, 5%, and 10% levels, respectively.
-See Section 4.2 for discussions about this table.
+In both PIR (1939) and PIR (1934), cohorts exposed to centralization tend to have a greater number of individuals selected into the PIR.
+See Table 5 for a statistical analysis and Section 4.2 for discussions about this figure.
 Appendix B Online Theoretical Appendix
 To guide our empirical investigation, we develop a simple model to predict the impacts of centralization on application behavior and assignment.
-We first confirm that centralized admissions (Capp) was indeed designed to make the school seat allocation more meritocratic compared to decentralized admissions (Dapp).
+We first confirm that the centralized admissions (referred to as “Capp” in this section) was indeed designed to make the school seat allocation more meritocratic compared to the decentralized admissions (referred to as “Dapp” in this section).
 Our model also predicts changes in application behavior.